### v1

# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

Abstract
Evolutionary multitasking (EMT) has been attracting much attention over the past years. It aims to handle multiple optimization tasks simultaneously within limited computing resources assisted by inter-task knowledge transfer techniques. Numerous multitask evolutionary algorithms (MTEAs) for solving multitask optimization (MTO) problems have been proposed in the EMT field, but there lacks a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications. To address this issue, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT. It incorporates more than 30 MTEAs, more than 150 MTO problem cases with real-world applications, and more than 10 performance metrics. Moreover, for comparing MTEAs with traditional evolutionary algorithms, we modified more than 30 popular single-task evolutionary algorithms to be able to solve MTO problems in MToP. MToP is a user-friendly tool with a graphical user interface that makes it easy to analyze results, export data, and plot schematics. More importantly, MToP is extensible, allowing users to develop new algorithms and define new problems. The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
Index Terms:
Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
I Introduction
Evolutionary computation (EC) is inspired by natural evolution and has grown rapidly due to effectiveness and efficiency. With its robustness and ease of use, EC algorithms, also known as evolutionary algorithms (EAs), have achieved great success in solving black-box optimization problems. Designing specific EAs for different complex black-box optimization problems, such as constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6], has also been studied extensively. In recent years, with the increase in computational demands and the development of cloud computing, the use of EAs to handle a large number of optimization tasks simultaneously has gained much focus. This requirement is called evolutionary multitasking (EMT) [7]. A multitask optimization (MTO) problem of EMT with $K$ optimization tasks aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks. It can be formulated as follows:
$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t. }\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task. For multi-objective optimization task, its $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$ task. EMT has been successfully applied to engineering scheduling [5], feature selection in machine learning [8], anomaly detection [9], point cloud registration [10], and reinforcement learning [11].
In order to solve multiple optimization tasks faster and better concurrently, researchers have attempted to exploit the similarity among tasks to empower EAs with inter-task knowledge extraction and transfer [12]. Through knowledge transfer, EAs are then able to fully utilize implicit parallelism to obtain better solutions for multiple tasks with fewer computational resources [13]. The first attempt for EMT is the multifactorial EA [14], which proposes an implicit knowledge representation through random mating among optimization tasks. After that, a large number of EAs for solving MTO problems have been proposed. These developed multitask EAs (MTEAs) employ either a multifactorial framework [15, 16] that utilizes a single population for multiple tasks or a multi-population framework [17, 18] that assign separate populations for multiple tasks. In addition, for decision space mapping among different tasks, various techniques in MTEAs such as unified search space [14, 16], autoencoding [19, 20], affine transformation [21, 22], and adversarial generative model [23] have been successively proposed. Due to the significant influence of the amount of knowledge transfer for solving MTO problems, MTEAs with knowledge transfer adaptive control strategies, such as similarity judgment [15, 24], knowledge selection [18, 25], and historical feedback [26, 27], have also been studied. There is a growing number of MTEAs proposed by researchers, yet there is no standardized programming language, code pattern, or software platform for the EMT source codes. This presents difficulties for beginners to enter the field of EMT and for researchers to make convenient experimental comparisons of the algorithms.
As EMT is booming, it has become urgent to provide a convenient and user-friendly comparison software platform for benchmarking the performance of MTEAs. Moreover, the available source code and platform are essential for exploring real-world applications of EMT. Open-source and user-friendly software platforms can greatly facilitate the development of a research field. As for EC, some popular and successful software platforms helped researchers advance the development of EA and evolutionary optimization. PlatEMO [28] is developed for evolutionary multi-objective optimization and benchmarking multi-objective EAs. It facilitates researchers to design multi-objective EAs to deal with various types of multi-objective optimization problems. EvoX [29] is a distributed GPU-accelerated library that helps researchers design parallel EAs to accelerate the solving of complex optimization problems and training for reinforcement learning. However, different from the traditional single-task EC field, solving MTO problems with MTEAs requires simultaneous evolution of multiple optimization tasks with inter-task solution space mapping and knowledge transfer. Additionally, the performance metrics of EMT are diverse, with both single-task metrics and multitask metrics. These special requirements are hard to implement in existing tools and platforms. Moreover, existing research work on EMT presents various code implementations of algorithms and problems. Particularly, EMT has been extended to the subfields of many-task optimization [30, 31], multi-objective multitask optimization [32, 33], competitive multitask optimization [26, 34], and constrained multi-task optimization [35]. The researchers cannot easily benchmark MTEAs on various types of MTO problems.
In order to address these issues, we introduce an open-source MATLAB platform, named MTO-Platform (MToP), for the development of the EMT field. The main contributions of MToP are summarized as follows:
  1. 1.
    MToP has an easy-to-use graphical user interface (GUI) including test, experiment, and data-process modules. These modules can assist researchers in realizing problem characteristics, performing comparative experiments, parallel execution, statistically analyzing results, plotting result figures, and processing experimental data. Moreover, MToP provides a modular implementation of algorithms, problems and metrics. On top of that, extensive public application programming interface (API) functions are provided, such as population initialization, function evaluation, evolutionary operator, and environmental selection.
  2. 2.
    An abundance of algorithms, problems, and metrics are implemented with public API and included in MToP. More than 30 MTEAs (see Table I) cover single-objective, multi-objective, constrained, many-task, and competitive EMT types. Additionally, to facilitate the comparison of MTEAs with popular traditional EAs, more than 30 single-task EAs (see Table II) of various types, such as genetic algorithm, differential evolution, swarm intelligence, and evolution strategy, are also modified and included in MToP. For benchmark test problems, more than 150 benchmark MTO problems and several real-world applications (see Table III) are incorporated. Finally, various types of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, are also implemented in MToP.
  3. 3.
    MToP is extensible to easily adding new code of algorithms, problems, and metrics. After following the coding patterns and implementing functionality based on the public API, new code can be recognized and performed. Since MToP is completely open-source, researchers can utilize existing algorithms, problems, and metrics as the basis for new ideas. Through the open-source website GitHub, MToP will be continuously updated and receive new code [^1] MToP has been continuously updated since Sep. 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
TABLE I: Multitask evolutionary algorithms included in MToP (current version).
Algorithm Year Publication Objective Task Framework Constraint Description
MFEA [14] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
LDA-MFEA [36] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
MFDE&MFPSO [37] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
IMEA [38] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
G-MFEA [39] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
EMEA [19] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
MFEA-GHS [40] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
MFEA-DV [41] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
SBO [31] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
MFEA-II [15] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
SREMTO [42] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
MFMP [17] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
TLTLA [43] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
MaTDE [30] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
MFEA-AKT [16] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
MTEA-AD [18] 2022 TEVC Single Multi Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
DEORA [26] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
AT-MFEA [21] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
MFEA-DGD [44] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
MTES [45] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
BoKT [24] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
EMaTO-MKT [25] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
MTES-KG [46] 2023 TEVC Single Multi Multi-population Yes Multitask ES with knowledge-guided external sampling
BLKT-DE [47] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
MTEA-SaO [48] 2023 INS Single/Multi Multi Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
MTSRA [34] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
MO-MFEA [32] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominanted sort and crowding distance
MM-DE [49] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
AMT-NSGA-II [50] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
MO-MFEA-II [51] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
EMT-ET [27] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
EMT-PD [33] 2021 SMCA Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
MOMFEA-SADE [52] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
EMT-GS [23] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
MTEA-D-DN [53] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
TABLE II: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
Algorithms Year Publication Objective Category Constraint Description
GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
PSO [54] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
CMA-ES [55] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
jDE [56] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
MTV-DE [57] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
JADE [58] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
ECHT [59] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
xNES-as [60] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
rank-DE [61] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
SHADE [62] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
R1-NES [63] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
xNES&SNES [64] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
L-SHADE [65] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
CSO [66] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
FROFI [67] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
jSO [68] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
LSHADE44 [69] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
CAL-SHADE [70] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
OpenAI-ES [71] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
C2oDE [72] 2019 SMCA Single Constraint Handling Yes Composite DE for constrained optimization
DES [73] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
MPA [74] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
EO [75] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
NL-SHADE-RSP [76] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
AO [77] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
DeCODE [78] 2021 SMCA Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
VMCH [79] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
CEDE-DRL [80] 2023 SWEVO Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
SPEA2 [81] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
NSGA-II [82] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
MO-CMA-ES [83] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
MOEA/D-DE [84] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
LMOCSO [85] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
CMOCSO [86] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
TABLE III: Multitask optimization problems included in MToP (current version).
Problem Case Task Dimension Objective Special Description
CEC17-MTSO [87] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO [88] 6 2-50 50 Single - Single-objective many-task optimization problems
WCCI20-MTSO [89] 10 2 50 Single - Complex single-objective multitask optimization problems
WCCI20-MaTSO [89] 10 2-50 50 Single - Complex single-objective many-task optimization problems
CMT [35] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP [26] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
C-CPLX [34] 20 2 50 Single Competitive Complex competitive multitask optimization problems
CEC17-MTMO [90] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO [88] 10 2 50 Multi - Complex multi-objective multitask optimization problems
CEC19-MaTMO [88] 10 2-50 50 Multi - Multi-objective many-task optimization problems
WCCI20-MaTMO [89] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
CEC21-MTMO [91] 10 2 50 Multi - More complex multi-objective multi optimization problems
PKACP [24] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
SCP [26] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
PEPVM [48] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
OPF [46] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
SOPM 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
MO-OPF 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
MO-SCP 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
TABLE IV: Metrics included in MToP (current version).
Metric Objective Description
Obj Single Objective value for each task
Obj (MTS) Single Multitask score of Obj for all tasks
Obj (AV) Single Average Obj for all tasks
Obj (CMT) Single Competitive multitask Obj for all tasks
Obj (NBR) Single Number of best Obj result for all tasks
CV Single Constraint violation for each task
FR Single Feasible rate for each task
HV Multi Hypervolume for each task
HV (MTS) Multi Multitask score of HV for all tasks
HV (CMT) Multi Competitive multitask HV for all tasks
IGD Multi Inverted generational distance for each task
IGD+ Multi Improved IGD plus for each task
IGD+ (MTS) Multi Multitask score of IGD+ for all tasks
IGD+ (CMT) Multi Competitive multitask IGD+ for all tasks
Run Time All Algorithm running time for all tasks
The rest of this paper is organized as follows. Section II elaborates on the architecture of MToP. Section III describes approaches for performing experiments with MToP. Section IV introduces how to extend contents of MToP. Finally, the conclusion and future work are given in Section V.
II Architecture of MToP
II-A Functional Modules
II-A1 Test Module
(a) Function landscapes
(b) Feasible regions
(c) Pareto front
(d) Metric convergence
Figure 1: Examples of graphic display in the test module of MToP.
The test module is designed to facilitate researchers in analyzing the characteristics of MTO problems and algorithms. The test module of MToP provides functions to run the selected algorithm on the selected problem. It provides visualization functions for both single-objective and multi-objective MTO problems. For single-objective MTO problems, the landscapes of the objective function with different tasks in the one-dimensional unified search space can be displayed as in Fig. 1 (a). For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be shown as in Fig. 1 (b). For multi-objective MTO problems, the Pareto front figures can be displayed for multiple tasks as in Fig. 1 (c). The metric convergence behavior can be drawn after running algorithms on problems as in Fig. 1 (d).
Figure 2: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
The workflow of API calls for the execution of an algorithm on a problem is displayed in Fig. 2. The workflow starts with the Start_Button event in MTO_GUI.m, which calls the parameter setting API Set_Parameter in the selected algorithm MFEA.m. Following this, it enters the Run function of the algorithm, which at the beginning calls the Initialization API of the algorithm base class Algorithm.m and later loops through the function evaluation API Evaluation. The Initialization gets the parameter settings of the problem CEC17_MTP1.m to be solved via the problem base class Problem.m. Similarly, the Evaluation gets the specific optimization task object through the problem base class. After the algorithm has run, GUI then displays the resultant performance metrics after calling the metrics function Obj.m via Call_Metric API. All functions, including the subsequent experiment module, in MToP that execute algorithms are based on this procedure. Both algorithm and problem parameters can be modified directly through the GUI of MToP.
II-A2 Experiment Module
TABLE V: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
GA MFEA MTEA-AD
CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
+ / - / = Base 8 / 5 / 5 9 / 6 / 3
(a) Convergence
(b) Pareto front
Figure 3: schematic plotting by the experiment module of MToP.
The experiment module provides functions for performing comprehensive experiments and analyzing results. Users can execute multiple independent runs of multiple algorithms to solve multiple MTO problems. Data such as experimental settings, objective values, and decision variables of the execution are recorded and stored after running is completed. Based on these data, metric results, significance tests, and schematic plotting would be able to be continued. Metrics functions running on the data are divided into single-task and multitask metrics. Selecting single-task metrics displays the results of each task of the MTO problem on the table in MToP, and selecting multitask metrics displays the results of each MTO problem. The results of the significance tests are also shown in the table along with the metric results. All table data can be exported in tex, csv, and xlsx types as shown in Table V. Convergence plots are displayed based on the selected table data to analyze the comparison of the convergence behavior of different algorithms as in Fig. 3 (a). Additionally, for the multi-objective MTO algorithms and problems, the Pareto front can plotted as in Fig. 3 (b).
II-A3 Data Process Module
TABLE VI: Properties of experimental data.
Property Description
Reps Repetitions number of independent runs
Algorithms Algorithms data contains names and parameter settings
Problems Problems data contains names and parameter settings
Results Total results data contains Obj, CV, and Dec
RunTimes Running time data of algorithms on problems
Metrics Calculated metric results data
Each unique execution in the experiment module is treated as a data object and saved to a data file MTOData.mat. The settings and intermediate results of this experiment are stored in data properties as shown in Table VI. The saved data can subsequently be reloaded by the experiment module of MToP. With the data process module, the data can be merged or split which allows data to be reused and customized. The processing of data is grouped into three types, which are independent runs, algorithms and problems. Having several sets of data with fewer independent runs, the data can then be merged to become new data with a larger number of independent runs. In cases where existing data needs to be augmented with new algorithms for comparison, with the data process module there is no need to re-execute all the comparative experiments. It is only necessary to run the new algorithms and merge the resulting data with the original data. Similarly, if not all algorithms in the data are needed, it can be split and re-merged. As with the data processing for algorithms, the data processing for problems is applied in cases where new problems need to be added or deleted.
II-B File Structure
Figure 4: File structure of MToP.
In the root directory MTO/ of MToP, there is a running script file mto.m and four folders, Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is shown in Fig. 4. The algorithms folder Algorithms/ holds various categories folder, utility functions, and the algorithm base class Algorithm.m. The specific algorithm files, e.g., MFEA.m, MO_MFEA.m, GA.m, are stored under their corresponding classification folder. The utility functions folder Algorithms/Utils/ provides various public utility functions of generation operators, environmental selection approaches, constraint handling techniques, etc. The problems folder Problems/ contains various categories folder, base function folder, and the problem base class Problem.m. The specific problem files, e.g., WCCI20_MTSO1.m, CMT3.m, CEC19_MaTMO6.m, are stored under their corresponding classification folder. In particular, files for real-world application problems are placed in Real-World Application/ folder. The Problems/Base/ folder includes the public base functions for problems. The third folder Metrics/ in MToP contains all the result evaluation metrics and is with two subfolders, Metrics/Single-objective/ and Metrics/Multi-objective for single-objective and multi-objective optimization respectively. Specific metric files are Obj.m, Obj_MTS.m, IGD.m, etc. The last folder is the GUI/ which contains all the files used by the GUI of MToP. Among them, MTO_GUI.m is the main file of the GUI, and MTO_CMD.m provides the function of executing experiments by the command line.
II-C Code Pattern
II-C1 Algorithm
TABLE VII: Properties and methods of algorithm base class.
Property or method Description
FE Number of fitness function evaluations
Gen Number of evolutionary generations
Best Best individual found for single-objective optimization
Result Result data contains Obj, CV, and Dec
getParameter() Get customized parameters for algorithm object
setParameter() Set customized parameters for algorithm object
notTerminated() Determine whether to terminate and update result data
Evaluation() Fitness function evaluation
run() Executing algorithm
1    label={list:algo},
2    caption={Algorithm implementation example.},
3    captionpos=b,
4    float=htbp
5]
6classdef Algo_Example < Algorithm
7% <Multi-task> <Single-objective> <None/Constrained>
8
9properties % set parameter
10  para_example = 0.5;
11end
12
13methods
14  function run(Algo, Prob)
15    % initialize populations for all tasks
16    pops = Initialization(Algo, Prob, Individual);
17
18    while Algo.notTerminated(Prob, pops) % main loop
19      for t = 1:Prob.T % for each task
20        % generate new offspring
21        off = Generation(pops{t});
22        % evaluate the fitness of offspring
23        off = Algo.Evaluation(off, Prob, t);
24        % environmental selection
25        pops{t} = Selection(pops{t}, off);
26      end
27    end
28  end
29end
All algorithms in MToP inherit from the algorithm base class Algorithm.m. The properties and methods of the algorithm base class are shown in Table VII. The base class provides all the functions that interact with the GUI, so the implemented algorithms only need to consider the evolutionary workflow itself. A simple example of algorithm implementation is shown in Listing LABEL:list:algo. The second line here is labels for the algorithms, inspired by PlatEMO [28], for the algorithm classification displayed in the GUI. The function run() of the class is for the execution of the algorithm, with its first argument Algo denoting the object itself and the second argument Prob denoting the problem object to be solved. At the beginning of the run() function, the population initialization is performed with the Initialization() function. MToP provides two population initialization functions, the multi-population method and the multifactorial method, which produce either multiple populations or single populations with skill factors respectively. Here the pops produced in line 11 are of cell data type containing multiple populations. Each individual in the pops is Individual object and contains properties including objective value Obj, constraint violation CV, and decision variables Dec. Then, the main loop is entered with the function notTerminated() implemented in the algorithm base class. The function notTerminated() provides data updates and generation counting methods. In the main loop, the functions of offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection() are performed separately for each task. For algorithms with the multifactorial framework, these functions are executed simultaneously for all tasks. The function Generation() acts on the decision variables Dec of offspring individuals and is implemented by specific algorithms. The evaluation function Evaluation() is implemented in the algorithm base class, which automatically calls the problem object Prob for offspring fitness evaluation, i.e., calculating the objective value Obj and constraint violation CV for all individuals. It also updates the best solution automatically for single-objective optimization. At the end of each loop, through the environmental selection function Selection(), the new populations pops are updated. MToP provides some universal Selection() functions, and this function can also be reimplemented by specific algorithms. After that, the code goes to the next loop and calls notTerminated() to record changes. The algorithm code pattern in MToP is adapted to all EAs for solving MTO problems.
II-C2 Problem
TABLE VIII: Properties and methods of problem base class.
Property or method Description
T Number of optimization tasks
N Population size for each task
M Number of objectives for all tasks
D Number of decision variable dimensions for all tasks
Fnc Fitness function for all tasks
Lb Lower bound of decision variables for all tasks
Ub Upper bound of decision variables for all tasks
maxFE Maximum fitness function evaluations
getRunParameter() Get public parameters for problem object
getParameter() Get customized parameters for problem object
setParameter() Set customized parameters for problem object
setTasks() Set optimization tasks
getOptimum() Get optimum solutions for multi-objective optimization
1    label={list:prob},
2    caption={Problem implementation example.},
3    captionpos=b,
4    float=htbp
5]
6classdef Prob_Example < Problem
7% <Multi-task> <Multi-objective> <None>
8
9methods
10function Prob = Prob_Example(varargin)
11    Prob = Prob@Problem(varargin);
12    % set default maximum function evaluations
13    Prob.maxFE = 1000 * 100;
14  end
15
16  function setTasks(Prob)
17    Prob.T = 2;
18    % task 1
19    Prob.D(1) = 10; % variable dimensions
20    Prob.M(1) = 2; % objective number
21    Prob.Fnc{1} = @func1 % fitness function
22    Prob.Lb{1} = zeros(1, 10); % lower bound
23    Prob.Ub{1} = ones(1, 10); % upper bound
24    % task 2
25    Prob.D(2) = 20; % variable dimensions
26    Prob.M(2) = 3; % objective dimensions
27    Prob.Fnc{2} = @func2 % objective function
28    Prob.Lb{2} = [0,-ones(1, 19)]; % lower bound
29    Prob.Ub{2} = [1,ones(1, 19)]; % upper bound
30  end
31
32  function optimum = getOptimum(Prob)
33    % return optimum points for each task
34    optimum{1} = optimum_matrix1;
35    optimum{2} = optimum_matrix2;
36  end
37end
All problems in MToP inherit from the problem base class Problem.m. The properties and methods of the Problem.m are shown in Table VIII. A simple example of problem implementation is presented in Listing LABEL:list:prob. The maximum number of function evaluations maxFE for the problem can be defined in the constructor of the class. The function setTasks() is for setting the properties and evaluation function for each optimization task. In this function, the number of tasks T of the problem is first defined followed by a detailed per-task setup. After that, decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub need to be set separately for each task. M and D for multiple tasks of the problem are denoted as vector type. Fnc, Lb and Ub are denoted as cell type. Note that the function Fnc for each task takes inputs of the decision variable Dec and returns objective values Obj and the constraint violation value CV. Additionally, for multi-objective optimization, the external function getOptimum(), which obtains optimal solutions (i.e., reference points), needs to be defined. With these problem code patterns, all problems that are completely defined can be solved using the corresponding type of algorithm in MToP.
II-C3 Metric
1    label={list:metric},
2    caption={Metric implementation example.},
3    captionpos=b,
4    float=htbp
5]
6function result = Metric_Example(MTOData)
7% <Metric> <Multi-objective>
8
9result.Metric = ’Min’;
10
11% Data for shown in the GUI table
12result.RowName = {MTOData.Problems.Name};
13result.ColumnName = {MTOData.Algorithms.Name};
14result.TableData = CalculateTableData(MTOData);
15
16% Data for shown in the GUI convergence plot
17result.ConvergeData = CalculateConvergeData(MTOData);
18
19% Data for shown in the GUI Pareto plot
20result.ParetoData = CalculateParetoData(MTOData);
21end
In contrast to algorithm and problem class implementation, metric codes in MToP are recognized as the function type. A simple example of metric implementation is shown in Listing LABEL:list:metric. The input parameter of the function is MTOData generated by the experimental execution. The function returns result, which contains properties of Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The property Metric can be either ‘Min’ or ‘Max’, indicating that a smaller or larger metric value is better. RowName, ColumnName, and TableData are used to return the metric results for the GUI table. ConvergeData and ParetoData are used to return metric convergence results for convergence plots and non-dominated solutions for Pareto plots respectively. Note that specific metric calculation functions are not covered here for the simplicity of this example.
III Performing Experiments with MToP
(a) Test module
(b) Experiment module
Figure 5: Graphical user interface of MToP.
Running MToP requires first executing the script file mto.m in the root directory. After that, the MToP GUI is shown as presented in Fig. 5.
III-A Testing Algorithms and Problems
As shown in Fig. 5 (a), the test module is divided into left, center, and right parts. The left part is for the selection and setup of algorithms and problems. After selecting the corresponding labels of Task, Objective, and Special drop-down buttons, the Algorithm and Problem drop-down buttons update their contents. Selecting specific algorithms and problems displays the parameters they contain in the edit box, which can be also clicked on and set. After parameters are updated, new parameter values are fed back into the algorithm and problem objects through the setParameter() method. The parameters of the algorithms generally contain evolutionary operator parameters and their customized parameters. The problem parameters generally contain population size N, maximum function evaluations maxFE, and their customized parameters. The center part of the MToP GUI can execute the test experiment and display the graphic results. After executing the procedure of the algorithm on the problem, the metric results of the problem are displayed in the right part of the test module.
III-B Executing Comparative Experiments
The GUI of the experiment module is shown in Fig. 5 (b). It is divided into a two-column experimental setup part on the left and a results display part on the right. The first column contains the basic settings to perform experiments. The text box No. of Runs indicates the number of independent repeat runs for each algorithm on each problem. The text box No. of Results denotes the number of results saved by the evolutionary process of each algorithm, which is uniformly sampled according to the number of generations. The Parallel button controls whether experiments are executed in parallel processing. Parallel execution invokes multiple threads to run simultaneously multiple independent runs. The drop-down buttons Task, Objective, and Special serve the same purpose as the test module, filtering algorithms and problems based on their pre-defined labels. After selecting the algorithm and problem in the two boxes on the bottom side of the first column, it can be added to the preparation area of the second column by clicking the Add button. The names and parameters of the added algorithms and problems can be edited by clicking on them. After these are ready, click on the green Start button and MToP starts executing the experiment. It is possible to click on the buttons Pause and Stop to pause or stop the experiment during the execution.
When the experiment has been completed, the content of the table is available in the results area on the right part of the experiment module. The original data of the experiment can be read and saved by Load Data and Save Data buttons. In the right part, The buttons in the first function row operate on the experimental data. The first button Save Table in the line can export the results of the table to xlsx, csv, and tex files. The second and third Converge buttons control convergence y-lim types and draw convergence plots based on the results selected in the table. The fourth button Pareto draws Pareto front and non-dominated solutions based on the selected results. The fifth edit box controls the type of result formatting of the table. The drop-down box right of the edit box is for selecting performance metrics. The drop-down boxes following the metrics are the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After the statistical test method is selected, the calculated statistical results, denoted by “+/-/=” followed by each metric result, can be also displayed in the table. MToP provides two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test. The highlight button highlights the best and worst algorithm’s metrics results for each problem or task.
III-C Executing via Alternative Command Line
In addition to GUI execution, MToP provides alternative options for running experiments via the command line. After inputting the algorithm list, problem list and other parameters into the script file mto.m, the experiments can be run directly in the command line and the data MTOData.mat are returned after running completed. Similarly, inputting the data into the metrics function yields the calculated metric results.
IV Extending MToP
New algorithms, problems and performance metrics can be easily added in MToP. Users can create and implement code files following the code pattern described in Section II-C.
IV-A Adding New Algorithms
The first step in adding an algorithm to MToP is to create a MATLAB class file that inherits the base class Algorithm.m in the Algorithms/ folder. The algorithmic framework can then be implemented following the code pattern as in Listing LABEL:list:algo. Users can use the population initialization function Initialization(), offspring generation function Generation(), and environmental selection function Selection() available in MToP, or of course reimplement these functions according to their requirements. In addition, users who want to create customized parameters that can be clicked and modified in the MToP GUI require an additional implementation of getParameter() and setParameter() API in the new algorithm class file.
IV-B Adding New Problems
Similar to adding algorithms, adding problems in MToP requires the creation of a specific problem class file that inherits the base class Problem.m in the Problems/ folder. The suggested maximum function evaluations maxFE need to be predefined in the constructor of the specific problem. After that, refer to Listing LABEL:list:prob to implement the configuration of each optimization task in setTasks() function. When optimization tasks include multi-objective optimization problems, it is necessary to additionally implement the getOptimum() function to return the optimum solution set or reference points. Finally, for setting customized parameters that can be modified in the GUI, it is also required to implement getParameter() and setParameter() functions.
IV-C Adding New Performance Metrics
Users can add new performance metrics by creating a new metrics function file in the Metrics/ folder. Follow the pattern as in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. The metric results displayed in the GUI table need to return the RowName, ColumnName, and TableData fields. The RowName for combined multitask metrics is typically the problem name, while metrics on individual tasks need to return the name of each task. The ColumnName is typically the algorithm name. The calculation function of TableData needs to be implemented for the specific performance metric. To implement the function to plot the convergence of the metric, the ConvergeData field is also required. Additionally, for multi-objective optimization, the metric function should also return non-dominated solutions in the ParetoData field.
V Conclusion and Future Work
This paper introduced an open-source MATLAB platform MToP for EMT. MToP provides a user-friendly GUI, enriched algorithms and problems, and convenient code patterns. The current version of MToP includes more than 30 MTEAs, more than 30 single-task EAs (that can handle MTO problems), more than 150 MTO benchmark problems, and several real-world applications of EMT. The functional modules, file structure and code pattern of MToP are described in detail. Moreover, the procedures and extensions for conducting experiments using MToP are also presented.
The current MToP, while carefully reimplementing, modifying, and testing the included algorithms and problems, might still have some reproducing errors and bugs. We have been actively looking for and fixing the issues and continue to update MToP. Fortunately, MToP is uploaded on the open-source website GitHub, where followers and researchers can test and give feedback on issues. After the upload, we also received and updated many codes from researchers.
In the future, we will continue to develop and update MToP with inspiration and contributions from the academic community. We hope that MToP can become a convenient tool to help researchers advance the field of EMT and consequently advance the evolutionary computation field.
Acknowledgement
The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes. Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [28], for which the authors are also thankful.

### v2

@@
 Abstract
-Evolutionary multitasking (EMT) has been attracting much attention over the past years.
-It aims to handle multiple optimization tasks simultaneously within limited computing resources assisted by inter-task knowledge transfer techniques.
-Numerous multitask evolutionary algorithms (MTEAs) for solving multitask optimization (MTO) problems have been proposed in the EMT field, but there lacks a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications.
-To address this issue, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT.
-It incorporates more than 30 MTEAs, more than 150 MTO problem cases with real-world applications, and more than 10 performance metrics.
-Moreover, for comparing MTEAs with traditional evolutionary algorithms, we modified more than 30 popular single-task evolutionary algorithms to be able to solve MTO problems in MToP.
-MToP is a user-friendly tool with a graphical user interface that makes it easy to analyze results, export data, and plot schematics.
-More importantly, MToP is extensible, allowing users to develop new algorithms and define new problems.
+Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years.
+It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques.
+Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications.
+To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT.
+MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 10 performance metrics.
+Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems.
+MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization.
+More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains.
 The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
@@
 I Introduction
-Evolutionary computation (EC) is inspired by natural evolution and has grown rapidly due to effectiveness and efficiency.
-With its robustness and ease of use, EC algorithms, also known as evolutionary algorithms (EAs), have achieved great success in solving black-box optimization problems.
-Designing specific EAs for different complex black-box optimization problems, such as constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6], has also been studied extensively.
-In recent years, with the increase in computational demands and the development of cloud computing, the use of EAs to handle a large number of optimization tasks simultaneously has gained much focus.
-This requirement is called evolutionary multitasking (EMT) [7].
-A multitask optimization (MTO) problem of EMT with $K$ optimization tasks aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks.
-It can be formulated as follows:
+Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency.
+Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature.
+Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6].
+In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7].
+A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
 $\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t.
@@
 where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task.
-For multi-objective optimization task, its $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$ task.
-EMT has been successfully applied to engineering scheduling [5], feature selection in machine learning [8], anomaly detection [9], point cloud registration [10], and reinforcement learning [11].
-In order to solve multiple optimization tasks faster and better concurrently, researchers have attempted to exploit the similarity among tasks to empower EAs with inter-task knowledge extraction and transfer [12].
-Through knowledge transfer, EAs are then able to fully utilize implicit parallelism to obtain better solutions for multiple tasks with fewer computational resources [13].
-The first attempt for EMT is the multifactorial EA [14], which proposes an implicit knowledge representation through random mating among optimization tasks.
-After that, a large number of EAs for solving MTO problems have been proposed.
-These developed multitask EAs (MTEAs) employ either a multifactorial framework [15, 16] that utilizes a single population for multiple tasks or a multi-population framework [17, 18] that assign separate populations for multiple tasks.
-In addition, for decision space mapping among different tasks, various techniques in MTEAs such as unified search space [14, 16], autoencoding [19, 20], affine transformation [21, 22], and adversarial generative model [23] have been successively proposed.
-Due to the significant influence of the amount of knowledge transfer for solving MTO problems, MTEAs with knowledge transfer adaptive control strategies, such as similarity judgment [15, 24], knowledge selection [18, 25], and historical feedback [26, 27], have also been studied.
-There is a growing number of MTEAs proposed by researchers, yet there is no standardized programming language, code pattern, or software platform for the EMT source codes.
-This presents difficulties for beginners to enter the field of EMT and for researchers to make convenient experimental comparisons of the algorithms.
-As EMT is booming, it has become urgent to provide a convenient and user-friendly comparison software platform for benchmarking the performance of MTEAs.
-Moreover, the available source code and platform are essential for exploring real-world applications of EMT.
-Open-source and user-friendly software platforms can greatly facilitate the development of a research field.
-As for EC, some popular and successful software platforms helped researchers advance the development of EA and evolutionary optimization.
-PlatEMO [28] is developed for evolutionary multi-objective optimization and benchmarking multi-objective EAs.
-It facilitates researchers to design multi-objective EAs to deal with various types of multi-objective optimization problems.
-EvoX [29] is a distributed GPU-accelerated library that helps researchers design parallel EAs to accelerate the solving of complex optimization problems and training for reinforcement learning.
-However, different from the traditional single-task EC field, solving MTO problems with MTEAs requires simultaneous evolution of multiple optimization tasks with inter-task solution space mapping and knowledge transfer.
-Additionally, the performance metrics of EMT are diverse, with both single-task metrics and multitask metrics.
-These special requirements are hard to implement in existing tools and platforms.
-Moreover, existing research work on EMT presents various code implementations of algorithms and problems.
-Particularly, EMT has been extended to the subfields of many-task optimization [30, 31], multi-objective multitask optimization [32, 33], competitive multitask optimization [26, 34], and constrained multi-task optimization [35].
-The researchers cannot easily benchmark MTEAs on various types of MTO problems.
-In order to address these issues, we introduce an open-source MATLAB platform, named MTO-Platform (MToP), for the development of the EMT field.
-The main contributions of MToP are summarized as follows:
+Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task.
+EMT has found successful applications in various domains, including engineering scheduling [5, 8], nonlinear equation systems [9], feature selection in machine learning [10], anomaly detection [11], point cloud registration [12, 13], and reinforcement learning [14].
+To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [15, 16].
+Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [17, 18].
+The first attempt of EMT can be traced back to the multifactorial EA [19], which introduced an implicit knowledge representation via random mating among optimization tasks.
+Subsequently, numerous EAs tailored for MTO have emerged.
+These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23].
+Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [19, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs.
+Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated.
+Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes.
+This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
+As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs.
+Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field.
+Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field.
+In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization.
+For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems.
+EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training.
+Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions.
+EDOLAB [39] is an optimization tool specifically for dynamic optimization.
+However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks.
+Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques.
+Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics.
+These distinctive requirements pose unsolvable challenges for MTEA implementation in existing tools and platforms.
+Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45].
+Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
+To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. The main contributions of MToP are summarized as follows:
   1. 1.
-    MToP has an easy-to-use graphical user interface (GUI) including test, experiment, and data-process modules.
-These modules can assist researchers in realizing problem characteristics, performing comparative experiments, parallel execution, statistically analyzing results, plotting result figures, and processing experimental data.
-Moreover, MToP provides a modular implementation of algorithms, problems and metrics.
-On top of that, extensive public application programming interface (API) functions are provided, such as population initialization, function evaluation, evolutionary operator, and environmental selection.
+    MToP features a user-friendly graphical user interface (GUI) comprising test, experiment, and data-process modules.
+These modules facilitate researchers in understanding problem characteristics, conducting comparative experiments, solving problems in parallel, statistically analyzing results, plotting result figures, and managing experimental data.
+Moreover, MToP offers a modular implementation of algorithms, problems and performance metrics.
+On top of these, it provides an extensive public application programming interface (API) with template functions for population initialization, function evaluation, evolutionary operators, and environmental selection.
   2. 2.
-    An abundance of algorithms, problems, and metrics are implemented with public API and included in MToP.
-More than 30 MTEAs (see Table I) cover single-objective, multi-objective, constrained, many-task, and competitive EMT types.
-Additionally, to facilitate the comparison of MTEAs with popular traditional EAs, more than 30 single-task EAs (see Table II) of various types, such as genetic algorithm, differential evolution, swarm intelligence, and evolution strategy, are also modified and included in MToP.
-For benchmark test problems, more than 150 benchmark MTO problems and several real-world applications (see Table III) are incorporated.
-Finally, various types of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, are also implemented in MToP.
+    MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API.
+Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types (see Table IV).
+Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types (see Table V).
+In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications (see Table VII).
+Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities (see Table VIII).
   3. 3.
-    MToP is extensible to easily adding new code of algorithms, problems, and metrics.
-After following the coding patterns and implementing functionality based on the public API, new code can be recognized and performed.
-Since MToP is completely open-source, researchers can utilize existing algorithms, problems, and metrics as the basis for new ideas.
-Through the open-source website GitHub, MToP will be continuously updated and receive new code [^1] MToP has been continuously updated since Sep.
+    MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics.
+By adhering to established coding patterns and implementing functionality based on the public API, new code can be seamlessly integrated and utilized within MToP.
+Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas.
+Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep.
 21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
-TABLE I: Multitask evolutionary algorithms included in MToP (current version).
-Algorithm Year Publication Objective Task Framework Constraint Description
-MFEA [14] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
-LDA-MFEA [36] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
-MFDE&MFPSO [37] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
-IMEA [38] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
-G-MFEA [39] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
-EMEA [19] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
-MFEA-GHS [40] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
-MFEA-DV [41] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
-SBO [31] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
-MFEA-II [15] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
-SREMTO [42] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
-MFMP [17] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
-TLTLA [43] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
-MaTDE [30] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
-MFEA-AKT [16] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
-MTEA-AD [18] 2022 TEVC Single Multi Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
-DEORA [26] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
-AT-MFEA [21] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
-MFEA-DGD [44] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
-MTES [45] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
-BoKT [24] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
-EMaTO-MKT [25] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
-MTES-KG [46] 2023 TEVC Single Multi Multi-population Yes Multitask ES with knowledge-guided external sampling
-BLKT-DE [47] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
-MTEA-SaO [48] 2023 INS Single/Multi Multi Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
-MTSRA [34] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
-MO-MFEA [32] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominanted sort and crowding distance
-MM-DE [49] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
-AMT-NSGA-II [50] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
-MO-MFEA-II [51] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
-EMT-ET [27] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
-EMT-PD [33] 2021 SMCA Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
-MOMFEA-SADE [52] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
-EMT-GS [23] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
-MTEA-D-DN [53] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
-TABLE II: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
-Algorithms Year Publication Objective Category Constraint Description
-GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
-DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
-PSO [54] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
-CMA-ES [55] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
-jDE [56] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
-MTV-DE [57] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
-JADE [58] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
-ECHT [59] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
-xNES-as [60] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
-rank-DE [61] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
-SHADE [62] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
-R1-NES [63] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
-xNES&SNES [64] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
-L-SHADE [65] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
-CSO [66] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
-FROFI [67] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
-jSO [68] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
-LSHADE44 [69] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
-CAL-SHADE [70] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
-OpenAI-ES [71] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
-C2oDE [72] 2019 SMCA Single Constraint Handling Yes Composite DE for constrained optimization
-DES [73] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
-CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
-MPA [74] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
-EO [75] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
-NL-SHADE-RSP [76] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
-AO [77] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
-DeCODE [78] 2021 SMCA Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
-VMCH [79] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
-CEDE-DRL [80] 2023 SWEVO Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
-SPEA2 [81] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
-NSGA-II [82] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
-MO-CMA-ES [83] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
-MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
-MOEA/D-DE [84] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
-NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
-LMOCSO [85] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
-CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
-CMOCSO [86] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
-TABLE III: Multitask optimization problems included in MToP (current version).
-Problem Case Task Dimension Objective Special Description
-CEC17-MTSO [87] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MaTSO [88] 6 2-50 50 Single - Single-objective many-task optimization problems
-WCCI20-MTSO [89] 10 2 50 Single - Complex single-objective multitask optimization problems
-WCCI20-MaTSO [89] 10 2-50 50 Single - Complex single-objective many-task optimization problems
-CMT [35] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
-C2TOP&C4TOP [26] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
-C-CPLX [34] 20 2 50 Single Competitive Complex competitive multitask optimization problems
-CEC17-MTMO [90] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MTMO [88] 10 2 50 Multi - Complex multi-objective multitask optimization problems
-CEC19-MaTMO [88] 10 2-50 50 Multi - Multi-objective many-task optimization problems
-WCCI20-MaTMO [89] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
-CEC21-MTMO [91] 10 2 50 Multi - More complex multi-objective multi optimization problems
-PKACP [24] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
-SCP [26] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
-PEPVM [48] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
-OPF [46] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
-SOPM 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
-MO-OPF 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
-MO-SCP 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
-TABLE IV: Metrics included in MToP (current version).
-Metric Objective Description
-Obj Single Objective value for each task
-Obj (MTS) Single Multitask score of Obj for all tasks
-Obj (AV) Single Average Obj for all tasks
-Obj (CMT) Single Competitive multitask Obj for all tasks
-Obj (NBR) Single Number of best Obj result for all tasks
-CV Single Constraint violation for each task
-FR Single Feasible rate for each task
-HV Multi Hypervolume for each task
-HV (MTS) Multi Multitask score of HV for all tasks
-HV (CMT) Multi Competitive multitask HV for all tasks
-IGD Multi Inverted generational distance for each task
-IGD+ Multi Improved IGD plus for each task
-IGD+ (MTS) Multi Multitask score of IGD+ for all tasks
-IGD+ (CMT) Multi Competitive multitask IGD+ for all tasks
-Run Time All Algorithm running time for all tasks
 The rest of this paper is organized as follows.
@@
 Section III describes approaches for performing experiments with MToP.
-Section IV introduces how to extend contents of MToP.
+Section IV introduces how to extend the contents of MToP.
 Finally, the conclusion and future work are given in Section V.
@@
 II-A1 Test Module
-(a) Function landscapes
-(b) Feasible regions
-(c) Pareto front
-(d) Metric convergence
+(a) Function landscapes (1D)
+(b) Function landscapes (2D)
+(c) Feasible regions
+(d) Pareto front
+(e) Metric convergence
 Figure 1: Examples of graphic display in the test module of MToP.
-The test module is designed to facilitate researchers in analyzing the characteristics of MTO problems and algorithms.
-The test module of MToP provides functions to run the selected algorithm on the selected problem.
-It provides visualization functions for both single-objective and multi-objective MTO problems.
-For single-objective MTO problems, the landscapes of the objective function with different tasks in the one-dimensional unified search space can be displayed as in Fig.
-1 (a).
-For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be shown as in Fig.
-1 (b).
-For multi-objective MTO problems, the Pareto front figures can be displayed for multiple tasks as in Fig.
+The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms.
+It facilitates running the selected algorithm on the chosen problem within MToP.
+The module includes visualization functions for both single-objective and multi-objective MTO problems.
+In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig.
+1 (a)-(b).
+For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig.
 1 (c).
-The metric convergence behavior can be drawn after running algorithms on problems as in Fig.
+Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig.
 1 (d).
+Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig.
+1 (e).
+(a) Gen=1
+(b) Gen=100
+(c) Gen=200
+(d) Gen=500
+(e) Gen=700
+(f) Gen=1000
+Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space with the number of generations.
+(a) Gen=1
+(b) Gen=100
+(c) Gen=200
+(d) Gen=500
+(e) Gen=700
+(f) Gen=1000
+Figure 3: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the decision space with the number of generations.
+In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior.
+Figures 2 and 3 illustrate an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space.
+Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
+In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space.
+As evolution progresses through the first and middle phases (Gen=100,200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity.
+Subsequently, in the middle and late stages (Gen=500,700), the population gradually emphasizes diversity.
+By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task.
+The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task.
+However, these two tasks do not precisely overlap in the decision space.
+This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy.
+In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
 Figure 4: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
-The workflow of API calls for the execution of an algorithm on a problem is displayed in Fig.
-2.
-The workflow starts with the Start_Button event in MTO_GUI.m, which calls the parameter setting API Set_Parameter in the selected algorithm MFEA.m.
-Following this, it enters the Run function of the algorithm, which at the beginning calls the Initialization API of the algorithm base class Algorithm.m and later loops through the function evaluation API Evaluation.
-The Initialization gets the parameter settings of the problem CEC17_MTP1.m to be solved via the problem base class Problem.m.
-Similarly, the Evaluation gets the specific optimization task object through the problem base class.
-After the algorithm has run, GUI then displays the resultant performance metrics after calling the metrics function Obj.m via Call_Metric API.
-All functions, including the subsequent experiment module, in MToP that execute algorithms are based on this procedure.
-Both algorithm and problem parameters can be modified directly through the GUI of MToP.
+The workflow of API calls for executing an algorithm on a problem is depicted in Fig.
+4.
+The process initiates with the Start_Button event in MTO_GUI.m, triggering the parameter setting API Set_Parameter within the selected algorithm MFEA.m.
+Subsequently, it proceeds to the Run function of the algorithm, which initially invokes the Initialization API of the algorithm base class Algorithm.m, and then iterates through the function evaluation API Evaluation.
+During the initialization phase, the parameter settings of the problem CEC17_MTP1.m to be solved are obtained via the problem base class Problem.m.
+Similarly, the Evaluation function retrieves the specific optimization task object through the problem base class.
+Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function Obj.m via the Call_Metric API.
+All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure.
+Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
 II-A2 Experiment Module
-TABLE V: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
+TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
 GA MFEA MTEA-AD
@@
 Figure 5: schematic plotting by the experiment module of MToP.
-The experiment module provides functions for performing comprehensive experiments and analyzing results.
-Users can execute multiple independent runs of multiple algorithms to solve multiple MTO problems.
-Data such as experimental settings, objective values, and decision variables of the execution are recorded and stored after running is completed.
-Based on these data, metric results, significance tests, and schematic plotting would be able to be continued.
-Metrics functions running on the data are divided into single-task and multitask metrics.
-Selecting single-task metrics displays the results of each task of the MTO problem on the table in MToP, and selecting multitask metrics displays the results of each MTO problem.
-The results of the significance tests are also shown in the table along with the metric results.
-All table data can be exported in tex, csv, and xlsx types as shown in Table V.
-Convergence plots are displayed based on the selected table data to analyze the comparison of the convergence behavior of different algorithms as in Fig.
+The experiment module offers functionalities for conducting comprehensive experiments and analyzing results.
+Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems.
+Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored.
+These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
+Metrics functions operating on the data are categorized into single-task and multitask metrics.
+The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem.
+Significance test results are also included in the table alongside the metric results.
+All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
+Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig.
 5 (a).
-Additionally, for the multi-objective MTO algorithms and problems, the Pareto front can plotted as in Fig.
+The convergence plot function can be applied to both single- and multi-objective problems.
+Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig.
 5 (b).
 II-A3 Data Process Module
-TABLE VI: Properties of experimental data.
+TABLE II: Properties of experimental data.
 Property Description
@@
 Metrics Calculated metric results data
-Each unique execution in the experiment module is treated as a data object and saved to a data file MTOData.mat.
-The settings and intermediate results of this experiment are stored in data properties as shown in Table VI.
-The saved data can subsequently be reloaded by the experiment module of MToP.
-With the data process module, the data can be merged or split which allows data to be reused and customized.
-The processing of data is grouped into three types, which are independent runs, algorithms and problems.
-Having several sets of data with fewer independent runs, the data can then be merged to become new data with a larger number of independent runs.
-In cases where existing data needs to be augmented with new algorithms for comparison, with the data process module there is no need to re-execute all the comparative experiments.
-It is only necessary to run the new algorithms and merge the resulting data with the original data.
-Similarly, if not all algorithms in the data are needed, it can be split and re-merged.
-As with the data processing for algorithms, the data processing for problems is applied in cases where new problems need to be added or deleted.
+Each unique execution within the experiment module is considered a data object and is saved to a data file named MTOData.mat.
+The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II.
+This saved data can subsequently be reloaded by the experiment module of MToP.
+With the data process module, the data can be merged or split, enabling data reuse and customization.
+The processing of data is categorized into three types: independent runs, algorithms, and problems.
+If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs.
+Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments.
+Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data.
+Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly.
+The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
 II-B File Structure
 Figure 6: File structure of MToP.
-In the root directory MTO/ of MToP, there is a running script file mto.m and four folders, Algorithms/, Problems/, Metrics/, and GUI/.
-The structure of these folders is shown in Fig.
-4.
-The algorithms folder Algorithms/ holds various categories folder, utility functions, and the algorithm base class Algorithm.m.
-The specific algorithm files, e.g., MFEA.m, MO_MFEA.m, GA.m, are stored under their corresponding classification folder.
-The utility functions folder Algorithms/Utils/ provides various public utility functions of generation operators, environmental selection approaches, constraint handling techniques, etc.
-The problems folder Problems/ contains various categories folder, base function folder, and the problem base class Problem.m.
-The specific problem files, e.g., WCCI20_MTSO1.m, CMT3.m, CEC19_MaTMO6.m, are stored under their corresponding classification folder.
-In particular, files for real-world application problems are placed in Real-World Application/ folder.
-The Problems/Base/ folder includes the public base functions for problems.
-The third folder Metrics/ in MToP contains all the result evaluation metrics and is with two subfolders, Metrics/Single-objective/ and Metrics/Multi-objective for single-objective and multi-objective optimization respectively.
-Specific metric files are Obj.m, Obj_MTS.m, IGD.m, etc.
-The last folder is the GUI/ which contains all the files used by the GUI of MToP.
-Among them, MTO_GUI.m is the main file of the GUI, and MTO_CMD.m provides the function of executing experiments by the command line.
+In the root directory MTO/ of MToP, you’ll find a script file named mto.m and four folders: Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is depicted in Fig. 6.
+The Algorithms/ folder contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m.
+Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders.
+The Algorithms/Utils/ folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
+The Problems/ folder houses various categories of problems, the base function folder, and the problem base class Problem.m.
+Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders.
+Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
+The Metrics/ folder contains all result evaluation metrics, organized into subfolders for single-objective and multi-objective optimization (Metrics/Single-objective/ and Metrics/Multi-objective).
+Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
+Lastly, the GUI/ folder houses all files used by the GUI of MToP.
+Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
 II-C Code Pattern
 II-C1 Algorithm
-TABLE VII: Properties and methods of algorithm base class.
+TABLE III: Properties and methods of algorithm base class.
 Property or method Description
@@
 1    label={list:algo},
-2    caption={Algorithm implementation example.},
+2    caption={Algorithm implementation example for multi-population EA.},
 3    captionpos=b,
@@
 5]
-6classdef Algo_Example < Algorithm
+6classdef Algo_Example1_MP < Algorithm
 7% <Multi-task> <Single-objective> <None/Constrained>
@@
 13methods
-14  function run(Algo, Prob)
-15    % initialize populations for all tasks
-16    pops = Initialization(Algo, Prob, Individual);
+14function run(Algo, Prob)
+15% initialize multiple populations for all tasks
+16pops = Initialization(Algo, Prob, Individual);
 17
-18    while Algo.notTerminated(Prob, pops) % main loop
+18while Algo.notTerminated(Prob, pops) % main loop
 19  for t = 1:Prob.T % for each task
 20    % generate new offspring
-21        off = Generation(pops{t});
+21    off(index1) = Generation(pops{t});
+22    % knowledge transfer
+23    off(index2) = KnowledgeTransfer(pops)
 24    % evaluate the fitness of offspring
@@
 28  end
+29end
+30end
+1    label={list:algo2},
+2    caption={Algorithm implementation example for multifactorial EA.},
+3    captionpos=b,
+4    float=htbp
+5]
+6classdef Algo_Example2_MF < Algorithm
+7% <Multi-task> <Multi-objective> <None/Constrained>
+8
+9properties % set parameter
+10  para_example = 0.2;
+11end
+12
+13methods
+14function run(Algo, Prob)
+15% initialize populations for multifactorial evolution
+16population=Initialization_MF(Algo, Prob, Individual);
+17
+18while Algo.notTerminated(Prob, pops) % main loop
+19  % generate new offspring with random mating
+20  off = Generation(population);
+21  % evaluate based on offspring skill factor
+22  for t = 1:Prob.T
+23    idx = [off.MFFactor] == t;
+24    off(idx) = Algo.Evaluation(off(idx), Prob, t);
 25  end
-28  end
+26  % environmental selection
+27  population = Selection(population, off);
+28end
 29end
-All algorithms in MToP inherit from the algorithm base class Algorithm.m.
-The properties and methods of the algorithm base class are shown in Table VII.
-The base class provides all the functions that interact with the GUI, so the implemented algorithms only need to consider the evolutionary workflow itself.
-A simple example of algorithm implementation is shown in Listing LABEL:list:algo.
-The second line here is labels for the algorithms, inspired by PlatEMO [28], for the algorithm classification displayed in the GUI.
-The function run() of the class is for the execution of the algorithm, with its first argument Algo denoting the object itself and the second argument Prob denoting the problem object to be solved.
-At the beginning of the run() function, the population initialization is performed with the Initialization() function.
-MToP provides two population initialization functions, the multi-population method and the multifactorial method, which produce either multiple populations or single populations with skill factors respectively.
-Here the pops produced in line 11 are of cell data type containing multiple populations.
-Each individual in the pops is Individual object and contains properties including objective value Obj, constraint violation CV, and decision variables Dec.
-Then, the main loop is entered with the function notTerminated() implemented in the algorithm base class.
-The function notTerminated() provides data updates and generation counting methods.
-In the main loop, the functions of offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection() are performed separately for each task.
-For algorithms with the multifactorial framework, these functions are executed simultaneously for all tasks.
-The function Generation() acts on the decision variables Dec of offspring individuals and is implemented by specific algorithms.
-The evaluation function Evaluation() is implemented in the algorithm base class, which automatically calls the problem object Prob for offspring fitness evaluation, i.e., calculating the objective value Obj and constraint violation CV for all individuals.
-It also updates the best solution automatically for single-objective optimization.
-At the end of each loop, through the environmental selection function Selection(), the new populations pops are updated.
-MToP provides some universal Selection() functions, and this function can also be reimplemented by specific algorithms.
-After that, the code goes to the next loop and calls notTerminated() to record changes.
-The algorithm code pattern in MToP is adapted to all EAs for solving MTO problems.
+All algorithms within MToP inherit from the algorithm base class Algorithm.m.
+The properties and methods of this base class are detailed in Table III.
+This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself.
+The MTEAs implemented in MToP are outlined in Table IV.
+Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems.
+Specific algorithms belonging to this category are listed in Table V.
+Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively.
+The algorithm labels provided in the second line serve as identifiers for classification within the GUI.
+The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved.
+Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
+In Listing LABEL:list:algo, the variable pops is of cell data type, housing multiple populations as popspopulation1, population2, population3….
+Each individual within pops is an Individual object, containing properties such as objective value (Obj), constraint violation (CV), and decision variables (Dec).
+Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the population variable, directly containing individuals as population(individual1:factor1, individual2:factor1, individual3:factor2…), along with an additional property (MFFactor).
+Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class.
+This function oversees data updates and generation counting within the loop.
+During the main loop, distinct operations are carried out for offspring generation Generation(), offspring evaluation Evaluation(), and environmental selection Selection().
+In the context of knowledge transfer, the multi-population algorithm requires the implementation of the KnowledgeTransfer() function to acquire knowledge from other tasks.
+Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function.
+The Generation() function operates on the decision variables (Dec) of offspring individuals and is tailored to specific algorithms.
+Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object (Prob) for offspring fitness evaluation.
+This entails calculating the objective value (Obj) and constraint violation (CV) for all individuals.
+Following each loop iteration, the environmental selection function Selection() is invoked to update the new population.
+While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function.
+Subsequently, the code progresses to the next loop and invokes notTerminated() to document changes.
+The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
+TABLE IV: Multitask evolutionary algorithms included in MToP (current version).
+Algorithm Year Publication Objective Task Framework Constraint Description
+MFEA [19] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
+LDA-MFEA [46] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
+MFDE&MFPSO [47] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
+G-MFEA [48] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
+MFEA-GHS [49] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
+MFEA-DV [50] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
+MFEA-II [20] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
+SREMTO [51] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
+TLTLA [52] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
+MFEA-AKT [21] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
+ASCMFDE [53] 2021 TEVC Single Multi Multifactorial Yes Intertask transfer in aligned subspace for MFDE
+AT-MFEA [26] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
+MFEA-DGD [54] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
+IMEA [55] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
+EMEA [24] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
+MFMP [22] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
+DEORA [33] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
+MKTDE [56] 2022 TEVC Single Multi Multi-population Yes Meta-knowledge transfer-based multitask DE
+MTES [57] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
+BLKT-DE [58] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
+MTSRA [44] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
+MTEA-AD [23] 2022 TEVC Single Multi/Many Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
+MTES-KG [59] 2023 TEVC Single Multi/Many Multi-population Yes Multitask ES with knowledge-guided external sampling
+BoKT [30] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
+TRADE [60] 2023 TCYB Single Many Multi-population Yes Transferable adaptive parameter DE for many-task optimization
+MTEA-SaO [61] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
+KR-MTEA [62] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive multi-task EA based on knowledge reuse
+SBO [41] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
+MaTDE [40] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
+EMaTO-MKT [32] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
+MO-MFEA [42] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominated sort and crowding distance
+MO-MFEA-II [63] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
+EMT-PD [64] 2021 TSMC-S Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
+EMT-GS [29] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
+MM-DE [65] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
+AMT-NSGA-II [66] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
+EMT-ET [34] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
+MOMFEA-SADE [67] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
+MTEA-D-DN [68] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
+MTDE-MKTA [69] 2024 TEVC Multi Multi Multi-population Yes Multitask DE with multiple knowledge types and transfer adaptation
+TABLE V: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
+Algorithms Year Publication Objective Category Constraint Description
+GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
+DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
+jDE [70] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
+JADE [71] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
+rank-DE [72] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
+SHADE [73] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
+jSO [74] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
+LSHADE44 [75] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
+L-SHADE [76] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
+CAL-SHADE [77] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
+NL-SHADE-RSP [78] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
+PSO [79] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
+CSO [80] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
+MPA [81] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
+EO [82] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
+AO [83] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
+CMA-ES [84] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
+IPOP-CMA-ES [85] 2005 CEC Single Evolution Strategy Yes Restart CMA-ES with increasing population size
+sep-CMA-ES [86] 2008 PPSN Single Evolution Strategy Yes Separable CMA-ES for large-scale optimization
+xNES-as [87] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
+R1-NES [88] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
+xNES&SNES [89] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
+OpenAI-ES [90] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
+DES [91] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
+MTV-DE [92] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
+ECHT [93] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
+FROFI [94] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
+C2oDE [95] 2019 TSMC-S Single Constraint Handling Yes Composite DE for constrained optimization
+CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
+DeCODE [96] 2021 TSMC-S Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
+VMCH [97] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
+CEDE-DRL [98] 2023 SWEC Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
+CCEF-ECHT [99] 2024 TSMC-S Single Constraint Handling Yes Competitive and cooperative ensemble of constraint handling techniques
+SPEA2 [100] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
+NSGA-II [101] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
+NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
+MO-CMA-ES [102] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
+MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
+MOEA/D-DE [103] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
+CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
+LMOCSO [104] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
+CMOCSO [105] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
 II-C2 Problem
-TABLE VIII: Properties and methods of problem base class.
+TABLE VI: Properties and methods of problem base class.
 Property or method Description
@@
 13  Prob.maxFE = 1000 * 100;
-14  end
-15
-16  function setTasks(Prob)
+14end
+15function setTasks(Prob)
 16  Prob.T = 2;
@@
 28  Prob.Ub{2} = [1,ones(1, 19)]; % upper bound
-30  end
-31
-32  function optimum = getOptimum(Prob)
+29end
+30function optimum = getOptimum(Prob)
 31  % return optimum points for each task
@@
 33  optimum{2} = optimum_matrix2;
-36  end
-37end
-All problems in MToP inherit from the problem base class Problem.m.
-The properties and methods of the Problem.m are shown in Table VIII.
-A simple example of problem implementation is presented in Listing LABEL:list:prob.
-The maximum number of function evaluations maxFE for the problem can be defined in the constructor of the class.
-The function setTasks() is for setting the properties and evaluation function for each optimization task.
-In this function, the number of tasks T of the problem is first defined followed by a detailed per-task setup.
-After that, decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub need to be set separately for each task.
-M and D for multiple tasks of the problem are denoted as vector type.
-Fnc, Lb and Ub are denoted as cell type.
-Note that the function Fnc for each task takes inputs of the decision variable Dec and returns objective values Obj and the constraint violation value CV.
-Additionally, for multi-objective optimization, the external function getOptimum(), which obtains optimal solutions (i.e., reference points), needs to be defined.
-With these problem code patterns, all problems that are completely defined can be solved using the corresponding type of algorithm in MToP.
-II-C3 Metric
+34end
+35end
+All problems within MToP inherit from the problem base class Problem.m.
+The properties and methods of Problem.m are detailed in Table VI.
+As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [19].
+For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$, respectively, and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is linearly mapped as follows:
+$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
+The dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$.
+During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
+It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
+A simple problem implementation example is illustrated in Listing LABEL:list:prob.
+The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor.
+The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task.
+Within this function, the number of tasks T is first defined, followed by a detailed setup for each task.
+Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task.
+For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type.
+Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint violation value CV.
+Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
+With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP.
+The existing problems available in MToP are enumerated in Table VII.
+In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
+TABLE VII: Multitask benchmark and real-world optimization problems included in MToP (current version).
+Problem Case Task Dimension Objective Special Description
+CEC17-MTSO [106] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
+CEC19-MaTSO [107] 6 2-50 50 Single - Single-objective many-task optimization problems
+WCCI20-MTSO [108] 10 2 50 Single - Complex single-objective multitask optimization problems
+WCCI20-MaTSO [108] 10 2-50 50 Single - Complex single-objective many-task optimization problems
+CMT [45] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
+C2TOP&C4TOP [33] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
+C-CPLX [44] 20 2 50 Single Competitive Complex competitive multitask optimization problems
+CEC17-MTMO [109] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
+CEC19-MTMO [107] 10 2 50 Multi - Complex multi-objective multitask optimization problems
+CEC19-MaTMO [107] 10 2-50 50 Multi - Multi-objective many-task optimization problems
+WCCI20-MaTMO [108] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
+CEC21-MTMO [110] 10 2 50 Multi - More complex multi-objective multi optimization problems
+PKACP [30] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
+SCP [33] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
+PEPVM [61] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
+OPF [59] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
+SOPM [69] 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
+MO-OPF [69] 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
+MO-SCP [61] 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
+II-C3 Performance Metric
 1    label={list:metric},
@@
 21end
-In contrast to algorithm and problem class implementation, metric codes in MToP are recognized as the function type.
-A simple example of metric implementation is shown in Listing LABEL:list:metric.
-The input parameter of the function is MTOData generated by the experimental execution.
-The function returns result, which contains properties of Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData.
-The property Metric can be either ‘Min’ or ‘Max’, indicating that a smaller or larger metric value is better.
-RowName, ColumnName, and TableData are used to return the metric results for the GUI table.
-ConvergeData and ParetoData are used to return metric convergence results for convergence plots and non-dominated solutions for Pareto plots respectively.
-Note that specific metric calculation functions are not covered here for the simplicity of this example.
+Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions.
+An illustrative example of metric implementation is presented in Listing LABEL:list:metric.
+The input parameter of the function is MTOData, which is generated during experimental execution.
+The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData.
+The Metric property can take values of either Min’ or Max’, indicating whether a smaller or larger metric value is preferable.
+RowName, ColumnName, and TableData are utilized to present metric results in the GUI table.
+On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively.
+Note that specific metric calculation functions are not elaborated here for the sake of simplicity.
+The metrics currently integrated into MToP are enumerated in Table VIII.
+TABLE VIII: Metrics included in MToP (current version).
+Metric Objective Task Description
+Obj Single All Objective value for each task
+Obj (MTS) Single Multi Multitask score of Obj for all tasks
+Obj (AV) Single Multi Average Obj for all tasks
+Obj (CMT) Single Multi Competitive multitask Obj for all tasks
+Obj (NBR) Single Multi Number of best Obj result for all tasks
+CV Single All Constraint violation for each task
+FR Single All Feasible rate for each task
+HV Multi All Hypervolume for each task
+HV (MTS) Multi Multi Multitask score of HV for all tasks
+HV (CMT) Multi Multi Competitive multitask HV for all tasks
+IGD Multi All Inverted generational distance for each task
+IGD+ Multi All Improved IGD plus for each task
+IGD+ (MTS) Multi Multi Multitask score of IGD+ for all tasks
+IGD+ (CMT) Multi Multi Competitive multitask IGD+ for all tasks
+Run Time All All Algorithm running time for all tasks
 III Performing Experiments with MToP
@@
 Figure 7: Graphical user interface of MToP.
-Running MToP requires first executing the script file mto.m in the root directory. After that, the MToP GUI is shown as presented in Fig. 5.
+To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 7.
 III-A Testing Algorithms and Problems
-As shown in Fig.
-5 (a), the test module is divided into left, center, and right parts.
-The left part is for the selection and setup of algorithms and problems.
-After selecting the corresponding labels of Task, Objective, and Special drop-down buttons, the Algorithm and Problem drop-down buttons update their contents.
-Selecting specific algorithms and problems displays the parameters they contain in the edit box, which can be also clicked on and set.
-After parameters are updated, new parameter values are fed back into the algorithm and problem objects through the setParameter() method.
-The parameters of the algorithms generally contain evolutionary operator parameters and their customized parameters.
-The problem parameters generally contain population size N, maximum function evaluations maxFE, and their customized parameters.
-The center part of the MToP GUI can execute the test experiment and display the graphic results.
-After executing the procedure of the algorithm on the problem, the metric results of the problem are displayed in the right part of the test module.
+As depicted in Fig.
+7 (a), the test module interface is structured into left, center, and right sections.
+The left section facilitates the selection and configuration of algorithms and problems.
+Upon choosing the desired labels such as Task, Objective, and Special from the drop-down buttons, the options available for Algorithm and Problem are updated accordingly.
+Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them.
+Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method.
+Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size N, maximum function evaluations maxFE, and other specific parameters.
+In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically.
+Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
 III-B Executing Comparative Experiments
-The GUI of the experiment module is shown in Fig.
-5 (b).
-It is divided into a two-column experimental setup part on the left and a results display part on the right.
-The first column contains the basic settings to perform experiments.
-The text box No.
-of Runs indicates the number of independent repeat runs for each algorithm on each problem.
-The text box No.
-of Results denotes the number of results saved by the evolutionary process of each algorithm, which is uniformly sampled according to the number of generations.
-The Parallel button controls whether experiments are executed in parallel processing.
-Parallel execution invokes multiple threads to run simultaneously multiple independent runs.
-The drop-down buttons Task, Objective, and Special serve the same purpose as the test module, filtering algorithms and problems based on their pre-defined labels.
-After selecting the algorithm and problem in the two boxes on the bottom side of the first column, it can be added to the preparation area of the second column by clicking the Add button.
-The names and parameters of the added algorithms and problems can be edited by clicking on them.
-After these are ready, click on the green Start button and MToP starts executing the experiment.
-It is possible to click on the buttons Pause and Stop to pause or stop the experiment during the execution.
-When the experiment has been completed, the content of the table is available in the results area on the right part of the experiment module.
-The original data of the experiment can be read and saved by Load Data and Save Data buttons.
-In the right part, The buttons in the first function row operate on the experimental data.
-The first button Save Table in the line can export the results of the table to xlsx, csv, and tex files.
-The second and third Converge buttons control convergence y-lim types and draw convergence plots based on the results selected in the table.
-The fourth button Pareto draws Pareto front and non-dominated solutions based on the selected results.
-The fifth edit box controls the type of result formatting of the table.
-The drop-down box right of the edit box is for selecting performance metrics.
-The drop-down boxes following the metrics are the result display type, statistical test method, base algorithm, and result highlighting type, respectively.
-After the statistical test method is selected, the calculated statistical results, denoted by “+/-/=” followed by each metric result, can be also displayed in the table.
-MToP provides two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests.
-The base algorithm drop-down button controls the base algorithm in the statistical test.
-The highlight button highlights the best and worst algorithm’s metrics results for each problem or task.
+The GUI of the experiment module is displayed in Fig. 7 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
+In the first column, users can configure the basic experiment settings.
+The No.
+of Runs textbox specifies the number of independent repeat runs for each algorithm on each problem.
+The No.
+of Results textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations.
+The Parallel button toggles parallel processing, enabling multiple threads to run independent experiments simultaneously.
+The drop-down buttons Task, Objective, and Special filter algorithms and problems based on predefined labels, similar to the test module.
+Users can add selected algorithms and problems by clicking the Add button after choosing them from the respective drop-down boxes at the bottom.
+In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them.
+Once all configurations are set, clicking the green Start button initiates the experiment execution.
+Users can also pause or stop the experiment using the Pause and Stop buttons if needed.
+Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module.
+Users can read and save the original data of the experiment using the Load Data and Save Data buttons.
+In the results area, the first function row contains buttons for operating on the experimental data.
+The first button, labeled Save Table, exports the table results to xlsx, csv, and tex files.
+The subsequent buttons labeled Converge control the convergence y-lim types and generate convergence plots based on the selected results from the table.
+The Pareto button draws the Pareto front and non-dominated solutions based on the selected results.
+The edit box controls the formatting type of the table result.
+Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively.
+After selecting the statistical test method, calculated statistical results denoted by “+/-/=” followed by each metric result can also be displayed in the table.
+MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests.
+The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
 III-C Executing via Alternative Command Line
-In addition to GUI execution, MToP provides alternative options for running experiments via the command line.
-After inputting the algorithm list, problem list and other parameters into the script file mto.m, the experiments can be run directly in the command line and the data MTOData.mat are returned after running completed.
-Similarly, inputting the data into the metrics function yields the calculated metric results.
+In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line.
+By inputting the algorithm list, problem list, and other parameters into the script file mto.m, experiments can be executed directly from the command line.
+Upon completion, the experiment data MTOData.mat is returned.
+Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
 IV Extending MToP
-New algorithms, problems and performance metrics can be easily added in MToP. Users can create and implement code files following the code pattern described in Section II-C.
+Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP.
+By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
 IV-A Adding New Algorithms
-The first step in adding an algorithm to MToP is to create a MATLAB class file that inherits the base class Algorithm.m in the Algorithms/ folder.
-The algorithmic framework can then be implemented following the code pattern as in Listing LABEL:list:algo.
-Users can use the population initialization function Initialization(), offspring generation function Generation(), and environmental selection function Selection() available in MToP, or of course reimplement these functions according to their requirements.
-In addition, users who want to create customized parameters that can be clicked and modified in the MToP GUI require an additional implementation of getParameter() and setParameter() API in the new algorithm class file.
+To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the Algorithms/ folder.
+This class file should inherit from the base class Algorithm.m.
+Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo.
+Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended.
+However, users can also choose to reimplement these functions to suit their specific requirements.
+Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
 IV-B Adding New Problems
-Similar to adding algorithms, adding problems in MToP requires the creation of a specific problem class file that inherits the base class Problem.m in the Problems/ folder.
-The suggested maximum function evaluations maxFE need to be predefined in the constructor of the specific problem.
-After that, refer to Listing LABEL:list:prob to implement the configuration of each optimization task in setTasks() function.
-When optimization tasks include multi-objective optimization problems, it is necessary to additionally implement the getOptimum() function to return the optimum solution set or reference points.
-Finally, for setting customized parameters that can be modified in the GUI, it is also required to implement getParameter() and setParameter() functions.
+To integrate new problems into MToP, users should create a dedicated problem class file within the Problems/ folder.
+This class file should inherit from the base class Problem.m.
+It’s advisable to define the suggested maximum function evaluations (maxFE) within the constructor of the specific problem.
+Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function.
+In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points.
+Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
 IV-C Adding New Performance Metrics
-Users can add new performance metrics by creating a new metrics function file in the Metrics/ folder.
-Follow the pattern as in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object.
-The metric results displayed in the GUI table need to return the RowName, ColumnName, and TableData fields.
-The RowName for combined multitask metrics is typically the problem name, while metrics on individual tasks need to return the name of each task.
-The ColumnName is typically the algorithm name.
-The calculation function of TableData needs to be implemented for the specific performance metric.
-To implement the function to plot the convergence of the metric, the ConvergeData field is also required.
-Additionally, for multi-objective optimization, the metric function should also return non-dominated solutions in the ParetoData field.
+To incorporate new performance metrics into MToP, users should create a new metrics function file within the Metrics/ folder.
+Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object.
+For the metrics displayed in the GUI table, ensure that the RowName, ColumnName, and TableData fields are properly populated.
+Typically, the RowName for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task.
+The ColumnName typically represents the algorithm name.
+Implement the calculation function for the specific performance metric in the TableData field.
+For plotting metric convergence, ensure the ConvergeData field is populated.
+In the case of multi-objective optimization, include non-dominated solutions in the ParetoData field.
 V Conclusion and Future Work
-This paper introduced an open-source MATLAB platform MToP for EMT.
-MToP provides a user-friendly GUI, enriched algorithms and problems, and convenient code patterns.
-The current version of MToP includes more than 30 MTEAs, more than 30 single-task EAs (that can handle MTO problems), more than 150 MTO benchmark problems, and several real-world applications of EMT.
-The functional modules, file structure and code pattern of MToP are described in detail.
-Moreover, the procedures and extensions for conducting experiments using MToP are also presented.
-The current MToP, while carefully reimplementing, modifying, and testing the included algorithms and problems, might still have some reproducing errors and bugs.
-We have been actively looking for and fixing the issues and continue to update MToP.
-Fortunately, MToP is uploaded on the open-source website GitHub, where followers and researchers can test and give feedback on issues.
-After the upload, we also received and updated many codes from researchers.
-In the future, we will continue to develop and update MToP with inspiration and contributions from the academic community.
-We hope that MToP can become a convenient tool to help researchers advance the field of EMT and consequently advance the evolutionary computation field.
+This paper introduces MToP, an open-source MATLAB platform designed for EMT.
+MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures.
+The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT.
+This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
+While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs.
+Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub.
+The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
+Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community.
+It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field.
+Specifically, MToP is considering the incorporation of multitask combinatorial optimization problems [5, 6] and more real-world applications [8, 111] in the future.
 Acknowledgement
+All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles.
+The code provided by the original authors undergoes modifications based on the source code during its integration into MToP.
 The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes.

### v3

@@
 To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT.
-MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 10 performance metrics.
+MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 20 performance metrics.
 Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems.
@@
 Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task.
+For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space [8] as follows:
+$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
+In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$.
+During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
 EMT has found successful applications in various domains, including engineering scheduling [5, 9], nonlinear equation systems [10], feature selection in machine learning [11], anomaly detection [12], point cloud registration [13, 14], and reinforcement learning [15].
@@
 Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics.
-These distinctive requirements pose unsolvable challenges for MTEA implementation in existing tools and platforms.
+Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem.
+These bring uncertainty to the pattern of results display and analysis.
+These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms.
 Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45].
 Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
-To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field. The main contributions of MToP are summarized as follows:
+To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field.
+MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications.
+The reason for using MATLAB is that it is a widely used programming language in the field of EC.
+Moreover, MATLAB provides a user-friendly graphical interface and a powerful numerical computing environment, making it an ideal choice for developing MToP.
+More importantly, the existing MTEAs and MTO problems are mainly implemented in MATLAB, which can be directly integrated into MToP.
+The main contributions of MToP are summarized as follows:
   1. 1.
@@
     MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API.
-Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types (see Table IV).
-Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types (see Table V).
-In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications (see Table VII).
-Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities (see Table VIII).
+Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types.
+Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types.
+In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications.
+Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
   3. 3.
@@
 II Architecture of MToP
+MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis.
+These components allow for flexible experimentation through both GUI and command line interfaces.
+Following standardized code patterns, MToP integrates various algorithms and performance metrics.
 II-A Functional Modules
@@
 (e) Metric convergence
-Figure 1: Examples of graphic display in the test module of MToP.
+Figure 1: Examples of graphical display in the test module of MToP.
+(a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively.
+(c) shows the feasible regions of a single-objective problem with different tasks in the two-dimensional unified search space.
+(d) depicts the Pareto front of a multi-objective problem with multiple tasks.
+(e) displays the convergence behavior of metrics after executing algorithms on problems.
 The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms.
@@
 The module includes visualization functions for both single-objective and multi-objective MTO problems.
+Users can select different options to display depending on the problem type, resulting in different visual outputs, as demonstrated below.
 In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig.
@@
 1 (e).
+In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior.
+Figure 2 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space.
+CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives.
+The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$.
+The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.4$.
+Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
 (a) Gen=1
@@
 (f) Gen=1000
-Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space with the number of generations.
-(a) Gen=1
-(b) Gen=100
-(c) Gen=200
-(d) Gen=500
-(e) Gen=700
-(f) Gen=1000
-Figure 3: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the decision space with the number of generations.
-In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior.
-Figures 2 and 3 illustrate an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space.
-Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
+(g) Gen=1
+(h) Gen=100
+(i) Gen=200
+(j) Gen=500
+(k) Gen=700
+(l) Gen=1000
+Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations.
+The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
 In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space.
-As evolution progresses through the first and middle phases (Gen=100,200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity.
-Subsequently, in the middle and late stages (Gen=500,700), the population gradually emphasizes diversity.
+As evolution progresses through the first and middle phases (Gen=100 to 200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity.
+Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity.
 By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task.
@@
 The workflow of API calls for executing an algorithm on a problem is depicted in Fig.
-4.
-The process initiates with the Start_Button event in MTO_GUI.m, triggering the parameter setting API Set_Parameter within the selected algorithm MFEA.m.
-Subsequently, it proceeds to the Run function of the algorithm, which initially invokes the Initialization API of the algorithm base class Algorithm.m, and then iterates through the function evaluation API Evaluation.
-During the initialization phase, the parameter settings of the problem CEC17_MTP1.m to be solved are obtained via the problem base class Problem.m.
-Similarly, the Evaluation function retrieves the specific optimization task object through the problem base class.
-Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function Obj.m via the Call_Metric API.
+3.
+The process initiates with the <Start_Button> event in <MTO_GUI.m>, triggering the parameter setting API SetParameter() within the selected algorithm <MFEA.m>.
+Subsequently, it proceeds to the Run() function of the algorithm, which initially invokes the Initialization() API of the algorithm base class <Algorithm.m>, and then iterates through the function evaluation API Evaluation().
+During the initialization phase, the parameter settings of the problem <CEC17_MTP1.m> to be solved are obtained via the problem base class <Problem.m>.
+Similarly, the Evaluation() function retrieves the specific optimization task object through the problem base class.
+Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function <Obj.m> via the CallMetric() API.
 All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure.
@@
 (b) Pareto front
-Figure 5: schematic plotting by the experiment module of MToP.
+Figure 4: schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
 The experiment module offers functionalities for conducting comprehensive experiments and analyzing results.
@@
 Metrics Calculated metric results data
-Each unique execution within the experiment module is considered a data object and is saved to a data file named MTOData.mat.
+Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>.
 The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II.
@@
 The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
+In practical use, the specific operations for splitting and merging data are as follows:
+  • Data Splitting: If an experiment involves running both the MFEA and EMEA algorithms on same problems, the data can be split by selecting it in the <Data Tree> and clicking the <Algorithms Split> button to generate separate datasets for each algorithm.
+Similarly, data can also be split by independent runs or problems using the respective options.
+  • Data Merging: If the MFEA and EMEA algorithms were run on same problems in separate experiments, the data can be merged by selecting both datasets and clicking the <Algorithms Merge> button, provided the settings are identical except for the algorithm.
+This process can also be used to merge datasets by independent runs or problems.
 II-B File Structure
-Figure 6: File structure of MToP.
-In the root directory MTO/ of MToP, you’ll find a script file named mto.m and four folders: Algorithms/, Problems/, Metrics/, and GUI/. The structure of these folders is depicted in Fig. 6.
-The Algorithms/ folder contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m.
-Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders.
-The Algorithms/Utils/ folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
-The Problems/ folder houses various categories of problems, the base function folder, and the problem base class Problem.m.
-Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders.
-Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
-The Metrics/ folder contains all result evaluation metrics, organized into subfolders for single-objective and multi-objective optimization (Metrics/Single-objective/ and Metrics/Multi-objective).
-Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
-Lastly, the GUI/ folder houses all files used by the GUI of MToP.
-Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
+Figure 5: File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
+In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
+The structure of these folders is depicted in Fig.
+5.
+The <Algorithms/> folder contains various categories of algorithms, utility functions, and the algorithm base class <Algorithm.m>.
+Specific algorithm files such as <MFEA.m>, <MO_MFEA.m>, and <GA.m> are stored within their respective classification folders.
+The <Algorithms/Utils/> folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
+The <Problems/> folder houses various categories of problems, the base function folder, and the problem base class <Problem.m>.
+Specific problem files like <WCCI20_MTSO1.m>, <CMT3.m>, and <CEC19_MaTMO6.m> are stored under their corresponding classification folders.
+Real-world application problem files are located in the <Real-World Application/> folder, while the <Problems/Base/> folder contains public base functions for problems.
+The <Metrics/> folder contains all result evaluation metrics, organized into subfolders for single-objective <Metrics/Single-objective/> and multi-objective <Metrics/Multi-objective> optimization.
+Specific metric files include <Obj.m>, <Obj_MTS.m>, and <IGD.m>, among others.
+Lastly, the <GUI/> folder houses all files used by the GUI of MToP.
+Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
 II-C Code Pattern
@@
 29end
-All algorithms within MToP inherit from the algorithm base class Algorithm.m.
+All algorithms within MToP inherit from the algorithm base class <Algorithm.m>.
 The properties and methods of this base class are detailed in Table III.
 This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself.
-The MTEAs implemented in MToP are outlined in Table IV.
+The MTEAs implemented in MToP are outlined in Table LABEL:supp:tab:algo-mt of the supplementary files.
 Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems.
-Specific algorithms belonging to this category are listed in Table V.
+Specific algorithms belonging to this category are listed in Table LABEL:supp:tab:algo-st.
 Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively.
 The algorithm labels provided in the second line serve as identifiers for classification within the GUI.
-The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved.
+The run() function within each class orchestrates algorithm execution, with <Algo> representing the object itself and <Prob> denoting the problem object to be solved.
 Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
-In Listing LABEL:list:algo, the variable pops is of cell data type, housing multiple populations as popspopulation1, population2, population3….
-Each individual within pops is an Individual object, containing properties such as objective value (Obj), constraint violation (CV), and decision variables (Dec).
-Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the population variable, directly containing individuals as population(individual1:factor1, individual2:factor1, individual3:factor2…), along with an additional property (MFFactor).
+In Listing LABEL:list:algo, the variable <pops> is of <cell> data type, housing multiple populations as <population1, population2, population3...>.
+Each individual within <pops> is an <Individual> object, containing properties such as objective value <Obj>, constraint violation <CV>, and decision variables <Dec>.
+Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the <population> variable, directly containing individuals as <individual1:factor1, individual2:factor1, individual3:factor2...>, along with an additional property <MFFactor>.
 Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class.
@@
 Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function.
-The Generation() function operates on the decision variables (Dec) of offspring individuals and is tailored to specific algorithms.
-Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object (Prob) for offspring fitness evaluation.
-This entails calculating the objective value (Obj) and constraint violation (CV) for all individuals.
+The Generation() function operates on the decision variables <Dec> of offspring individuals and is tailored to specific algorithms.
+Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object <Prob> for offspring fitness evaluation.
+This entails calculating the objective value <Obj> and constraint violation <CV> for all individuals.
 Following each loop iteration, the environmental selection function Selection() is invoked to update the new population.
@@
 The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
-TABLE IV: Multitask evolutionary algorithms included in MToP (current version).
-Algorithm Year Publication Objective Task Framework Constraint Description
-MFEA [19] 2016 TEVC Single Multi Multifactorial Yes Multifactorial evolution-based GA for multitask optimization
-LDA-MFEA [46] 2017 CEC Single Multi Multifactorial - Linearized domain adaptation in MFEA
-MFDE&MFPSO [47] 2017 CEC Single Multi Multifactorial Yes Multifactorial evolution-based DE and PSO
-G-MFEA [48] 2019 TEVC Single Multi Multifactorial - Generalized MFEA with decision variable translation and shuffling
-MFEA-GHS [49] 2019 ESWA Single Multi Multifactorial Yes Genetic Transform and hyper-rectangle search in MFEA
-MFEA-DV [50] 2019 CEC Single Multi Multifactorial Yes Enhanced MFEA with cross-task search direction
-MFEA-II [20] 2020 TEVC Single Multi Multifactorial Yes MFEA with online transfer parameter estimation
-SREMTO [51] 2020 TEVC Single Multi Multifactorial Yes Self-regulated multitask framework for varying relatedness among tasks
-TLTLA [52] 2020 FNS Single Multi Multifactorial Yes MFEA with two-level inter- and intra-task transfer learning
-MFEA-AKT [21] 2021 TCYB Single Multi Multifactorial - MFEA with adaptive knowledge transfer via multiple crossover operator
-ASCMFDE [53] 2021 TEVC Single Multi Multifactorial Yes Intertask transfer in aligned subspace for MFDE
-AT-MFEA [26] 2022 TCYB Single Multi Multifactorial Yes Affine transformation-enhanced domain adaptation for MFEA
-MFEA-DGD [54] 2022 TCYB Single Multi Multifactorial - MFEA based on diffusion gradient descent
-IMEA [55] 2018 GECCO Single Multi Multi-population Yes Multi-population framework for multitask optimization through island model
-EMEA [24] 2019 TCYB Single Multi Multi-population Yes Explicit knowledge transfer for multitask optimization via autoencoding
-MFMP [22] 2020 INS Single Multi Multi-population Yes Multi-population-based adaptive DE for multitask optimization
-DEORA [33] 2022 TEVC Single Multi Multi-population - Adaptive task selection for competitive multitask optimization
-MKTDE [56] 2022 TEVC Single Multi Multi-population Yes Meta-knowledge transfer-based multitask DE
-MTES [57] 2022 TCYB Single Multi Multi-population Yes Multitask OpenAI-ES via gradient-free evolution multitasking
-BLKT-DE [58] 2023 TCYB Single Multi Multi-population Yes Multitask DE with block-level knowledge transfer
-MTSRA [44] 2023 ESWA Single Multi Multi-population - Improved adaptive DE with competitive task selection
-MTEA-AD [23] 2022 TEVC Single Multi/Many Multi-population Yes MTEA with adaptive knowledge transfer via anomaly detection
-MTES-KG [59] 2023 TEVC Single Multi/Many Multi-population Yes Multitask ES with knowledge-guided external sampling
-BoKT [30] 2022 TEVC Single Many Multi-population - Bi-objective knowledge transfer framework for many-task optimization
-TRADE [60] 2023 TCYB Single Many Multi-population Yes Transferable adaptive parameter DE for many-task optimization
-MTEA-SaO [61] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive solver multitask framework with implicit knowledge transfer
-KR-MTEA [62] 2023 INS Single/Multi Multi/Many Multi-population Yes Adaptive multi-task EA based on knowledge reuse
-SBO [41] 2019 AAAI Single/Multi Many Multi-population Yes Symbiosis in biocoenosis framework for many-task optimization
-MaTDE [40] 2020 TETCI Single/Multi Many Multi-population Yes Many-task DE with adaptive archive-based knowledge transfer
-EMaTO-MKT [32] 2022 TEVC Single/Multi Many Multi-population - Multi-source knowledge transfer via local distribution estimation
-MO-MFEA [42] 2017 TCYB Multi Multi Multifactorial Yes Multiobjective MFEA with non-dominated sort and crowding distance
-MO-MFEA-II [63] 2021 TCYB Multi Multi Multifactorial Yes Cognizant Multitasking for parameter estimation in MO-MFEA
-EMT-PD [64] 2021 TSMC-S Multi Multi Multifactorial Yes Two-stage adaptive knowledge transfer based on population distribution
-EMT-GS [29] 2023 TEVC Multi Multi Multifactorial Yes Generative adversarial networks for knowledge transfer
-MM-DE [65] 2018 CEC Multi Multi Multi-population Yes Fast memetic DE for multiobjective multitask optimization
-AMT-NSGA-II [66] 2019 TCYB Multi Multi Multi-population Yes Curbing negative influences online for seamless transfer
-EMT-ET [34] 2021 TCYB Multi Multi Multi-population Yes Effective knowledge transfer approach via non-dominated sort
-MOMFEA-SADE [67] 2022 TCYB Multi Multi Multi-population Yes Subspace alignment and adaptive Differential Evolution
-MTEA-D-DN [68] 2023 TEVC Multi Multi Multi-population - Neighborhood as a bridge for decomposition-based knowledge transfer
-MTDE-MKTA [69] 2024 TEVC Multi Multi Multi-population Yes Multitask DE with multiple knowledge types and transfer adaptation
-TABLE V: Single-task evolutionary algorithms for multitask optimization included in MToP (current version).
-Algorithms Year Publication Objective Category Constraint Description
-GA - - Single Genetic Algorithm Yes Genetic algorithm with SBX crossover and polynomial mutation
-DE - - Single Differential Evolution Yes Differential evolution algorithm with DE/rand/1/bin operator
-jDE [70] 2006 TEVC Single Differential Evolution Yes Self-adapting control parameters in DE
-JADE [71] 2009 TEVC Single Differential Evolution Yes Adaptive DE with fast convergence performance
-rank-DE [72] 2013 TCYB Single Differential Evolution Yes Ranking-based mutation operators for DE
-SHADE [73] 2013 CEC Single Differential Evolution Yes Success-history based parameter adaptation for DE
-jSO [74] 2017 CEC Single Differential Evolution Yes Champion of CEC 2017 numerical optimization competition via improved DE
-LSHADE44 [75] 2017 CEC Single Differential Evolution Yes L-SHADE with competing strategies applied to constrained optimization
-L-SHADE [76] 2014 CEC Single Differential Evolution Yes Linear population size reduction for SHADE
-CAL-SHADE [77] 2017 CEC Single Differential Evolution Yes Adaptive constraint handling technique for L-SHADE
-NL-SHADE-RSP [78] 2021 CEC Single Differential Evolution - Champion of CEC 2021 numerical optimization competition via improved DE
-PSO [79] 1995 ICNN Single Swarm Intelligence Yes Particle swarm optimization with global and particle best update
-CSO [80] 2015 TCYB Single Swarm Intelligence Yes Competitive swarm optimizer for large-scale optimization
-MPA [81] 2020 ESWA Single Swarm Intelligence Yes Marine predators algorithm for numerical optimization
-EO [82] 2020 KBS Single Swarm Intelligence Yes Equilibrium optimizer for numerical optimization
-AO [83] 2021 CIE Single Swarm Intelligence Yes Aquila optimizer for numerical optimization
-CMA-ES [84] 2001 ECJ Single Evolution Strategy Yes Evolution strategy with derandomized covariance matrix adaptation
-IPOP-CMA-ES [85] 2005 CEC Single Evolution Strategy Yes Restart CMA-ES with increasing population size
-sep-CMA-ES [86] 2008 PPSN Single Evolution Strategy Yes Separable CMA-ES for large-scale optimization
-xNES-as [87] 2012 GECCO Single Evolution Strategy Yes Natural ES with adaptation sampling
-R1-NES [88] 2013 GECCO Single Evolution Strategy Yes A Linear Time natural ES for non-separable functions
-xNES&SNES [89] 2014 JMLR Single Evolution Strategy Yes ES with adaptive natural gradients to update distribution
-OpenAI-ES [90] 2017 ArXiv Single Evolution Strategy Yes Parallelized ES with standard normal distribution gradients
-DES [91] 2020 TEVC Single Evolution Strategy Yes Matrix-free covariance matrix adaptation ES
-MTV-DE [92] 2007 EngOpt Single Constraint Handling Yes Multiple trial vectors in DE for handling constraints
-ECHT [93] 2010 TEVC Single Constraint Handling Yes Ensemble of constraint handling techniques
-FROFI [94] 2016 TCYB Single Constraint Handling Yes Incorporating objective function information into the feasibility rule
-C2oDE [95] 2019 TSMC-S Single Constraint Handling Yes Composite DE for constrained optimization
-CORCO [1] 2020 TEVC Single Constraint Handling Yes Utilizing the correlation between constraints and objective function
-DeCODE [96] 2021 TSMC-S Single Constraint Handling Yes Decomposition-based multiobjective approach for constrained optimization
-VMCH [97] 2022 TEVC Single Constraint Handling Yes Voting-mechanism-based ensemble of constraint handling techniques
-CEDE-DRL [98] 2023 SWEC Single Constraint Handling Yes Deep reinforcement learning assisted co-evolutionary DE
-CCEF-ECHT [99] 2024 TSMC-S Single Constraint Handling Yes Competitive and cooperative ensemble of constraint handling techniques
-SPEA2 [100] 2001 EUROGEN Multi Non-dominated Sort Yes Improving the strength Pareto approach for multiobjective optimization
-NSGA-II [101] 2002 TEVC Multi Non-dominated Sort Yes Multiobjective GA with non-dominated sort and crowding distance
-NSGA-III [4] 2014 TEVC Multi Non-dominated Sort Yes Reference-point-based nondominated sorting for multiobjective optimization
-MO-CMA-ES [102] 2007 ECJ Multi Evolution Strategy - Covariance matrix adaptation for multiobjective optimization
-MOEA/D [3] 2007 TEVC Multi Decomposition - Decomposition-based multiobjective evolutionary algorithm
-MOEA/D-DE [103] 2009 TEVC Multi Decomposition - MOEA/D for complex Pareto sets in multiobjective optimization
-CCMO [2] 2021 TEVC Multi Constraint Handling Yes Coevolutionary framework for constrained multiobjective optimization
-LMOCSO [104] 2020 TCYB Multi Swarm Intelligence Yes Efficient large-scale multiobjective optimization based on CSO
-CMOCSO [105] 2022 TEVC Multi Swarm Intelligence Yes Competitive and cooperative swarm optimizer for multiobjective optimization
 II-C2 Problem
-TABLE VI: Properties and methods of problem base class.
+TABLE IV: Properties and methods of problem base class.
 Property or method Description
@@
 setTasks() Set optimization tasks
-getOptimum() Get optimum solutions for multi-objective optimization
+getOptimum() Get optimal solutions for multi-objective optimization
 1 label={list:prob},
@@
 35end
-All problems within MToP inherit from the problem base class Problem.m.
-The properties and methods of Problem.m are detailed in Table VI.
+All problems within MToP inherit from the problem base class <Problem.m>.
+The properties and methods of <Problem.m> are detailed in Table IV.
 As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [8].
-For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$, respectively, and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is linearly mapped as follows:
-$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
-The dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$.
-During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
 It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
 A simple problem implementation example is illustrated in Listing LABEL:list:prob.
-The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor.
+The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor.
 The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task.
-Within this function, the number of tasks T is first defined, followed by a detailed setup for each task.
-Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task.
-For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type.
-Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint violation value CV.
+Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task.
+Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task.
+For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type.
+Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>.
 Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
 With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP.
-The existing problems available in MToP are enumerated in Table VII.
+The existing problems available in MToP are enumerated in Table LABEL:supp:tab:prob-mt of the supplementary files.
 In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
-TABLE VII: Multitask benchmark and real-world optimization problems included in MToP (current version).
-Problem Case Task Dimension Objective Special Description
-CEC17-MTSO [106] 9 2 50 Single - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MaTSO [107] 6 2-50 50 Single - Single-objective many-task optimization problems
-WCCI20-MTSO [108] 10 2 50 Single - Complex single-objective multitask optimization problems
-WCCI20-MaTSO [108] 10 2-50 50 Single - Complex single-objective many-task optimization problems
-CMT [45] 9 2 Any Single Constrained Large/partial/no feasible regions interaction and different inter-task similarity
-C2TOP&C4TOP [33] 51 2/4 50 Single Competitive Different optimal objective values among competitive tasks
-C-CPLX [44] 20 2 50 Single Competitive Complex competitive multitask optimization problems
-CEC17-MTMO [109] 9 2 50 Multi - Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MTMO [107] 10 2 50 Multi - Complex multi-objective multitask optimization problems
-CEC19-MaTMO [107] 10 2-50 50 Multi - Multi-objective many-task optimization problems
-WCCI20-MaTMO [108] 10 2-50 50 Multi - Complex multi-objective many-task optimization problems
-CEC21-MTMO [110] 10 2 50 Multi - More complex multi-objective multi optimization problems
-PKACP [30] 1 Any Any Single - Planar kinematic arm control problem as many-task optimization problem
-SCP [33] 1 Any 60-120 Single Competitive Sensor coverage problem as competitive multitask optimization problem
-PEPVM [61] 1 3 7 Single - Parameter extraction of photovoltaic models as multitask optimization problem
-OPF [59] 5 2 24/33 Single Constrained Optimal power flow as constrained multitask optimization problem
-SOPM [69] 2 3 30 Multi Constrained Synchronous optimal pulse-width modulation as constrained multiobjective multitask problem
-MO-OPF [69] 4 2 24/33 Multi Constrained Multiobjective optimal power flow as constrained multiobjective multitask problem
-MO-SCP [61] 1 Any 60-120 Multi Competitive Multiobjective sensor coverage problem as competitive multiobjective multitask problem
 II-C3 Performance Metric
@@
 An illustrative example of metric implementation is presented in Listing LABEL:list:metric.
-The input parameter of the function is MTOData, which is generated during experimental execution.
-The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData.
-The Metric property can take values of either Min’ or Max’, indicating whether a smaller or larger metric value is preferable.
-RowName, ColumnName, and TableData are utilized to present metric results in the GUI table.
-On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively.
+The input parameter of the function is <MTOData>, which is generated during experimental execution.
+The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>.
+The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable.
+<RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table.
+On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively.
 Note that specific metric calculation functions are not elaborated here for the sake of simplicity.
-The metrics currently integrated into MToP are enumerated in Table VIII.
-TABLE VIII: Metrics included in MToP (current version).
+The metrics currently integrated into MToP are enumerated in Table V.
+TABLE V: Metrics included in MToP.
+The table lists the metric name, optimization objective, task type, and a brief description.
+SO and MO denote single-objective and multi-objective optimization, respectively.
+ST and MT represent single-task and multi-task, respectively
 Metric Objective Task Description
-Obj Single All Objective value for each task
-Obj (MTS) Single Multi Multitask score of Obj for all tasks
-Obj (AV) Single Multi Average Obj for all tasks
-Obj (CMT) Single Multi Competitive multitask Obj for all tasks
-Obj (NBR) Single Multi Number of best Obj result for all tasks
-CV Single All Constraint violation for each task
-FR Single All Feasible rate for each task
-HV Multi All Hypervolume for each task
-HV (MTS) Multi Multi Multitask score of HV for all tasks
-HV (CMT) Multi Multi Competitive multitask HV for all tasks
-IGD Multi All Inverted generational distance for each task
-IGD+ Multi All Improved IGD plus for each task
-IGD+ (MTS) Multi Multi Multitask score of IGD+ for all tasks
-IGD+ (CMT) Multi Multi Competitive multitask IGD+ for all tasks
-Run Time All All Algorithm running time for all tasks
+Obj SO ST/MT Objective value for each task
+Obj (AV) SO MT Average Obj for all tasks
+Obj (UV) SO MT Unified average Obj for all tasks
+Obj (MTS) SO MT Multitask score of Obj for all tasks
+Obj (CMT) SO MT Competitive multitask Obj for all tasks
+Obj (NBR) SO MT Number of best Obj result for all tasks
+CV SO ST/MT Constraint violation for each task
+FR SO ST/MT Feasible rate for each task
+HV MO ST/MT Hypervolume for each task
+HV (MTS) MO MT Multitask score of HV for all tasks
+HV (CMT) MO MT Competitive multitask HV for all tasks
+IGD MO ST/MT Inverted generational distance for each task
+IGD (AV) MO MT Average IGD for all tasks
+IGD+ (MTS) MO MT Multitask score of IGD for all tasks
+IGD (CMT) MO MT Competitive multitask IGD for all tasks
+IGD+ MO ST/MT Improved IGD plus for each task
+IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
+IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
+Spread MO ST/MT Spread metric for each task
+Spread (CMT) MO MT Competitive multitask spread for all tasks
+Run Time SO/MO ST/MT Algorithm running time for all tasks
 III Performing Experiments with MToP
@@
 Figure 6: Graphical user interface of MToP.
-To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 7.
+The test module (a) is used for testing algorithms and problems.
+The graphical display is shown in the center section of the test module.
+The experiment module (b) is used for executing comparative experiments.
+The results of the experiment are displayed in the right section of the experiment module.
+To launch MToP, start by running the script file <mto.m> located in the root directory.
+This action will initialize the MToP GUI interface, which is illustrated in Fig.
+6.
+The GUI interface of MToP requires MATLAB R2020b or later versions to run, while command-line execution can be done with any version.
+In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
 III-A Testing Algorithms and Problems
@@
 The left section facilitates the selection and configuration of algorithms and problems.
-Upon choosing the desired labels such as Task, Objective, and Special from the drop-down buttons, the options available for Algorithm and Problem are updated accordingly.
+Upon choosing the desired labels such as <Task>, <Objective>, and <Special> from the drop-down buttons, the options available for <Algorithm> and <Problem> are updated accordingly.
 Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them.
 Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method.
-Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size N, maximum function evaluations maxFE, and other specific parameters.
+Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size <N>, maximum function evaluations <maxFE>, and other specific parameters.
 In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically.
@@
 In the first column, users can configure the basic experiment settings.
-The No.
-of Runs textbox specifies the number of independent repeat runs for each algorithm on each problem.
-The No.
-of Results textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations.
-The Parallel button toggles parallel processing, enabling multiple threads to run independent experiments simultaneously.
-The drop-down buttons Task, Objective, and Special filter algorithms and problems based on predefined labels, similar to the test module.
-Users can add selected algorithms and problems by clicking the Add button after choosing them from the respective drop-down boxes at the bottom.
+The <No.
+of Runs> textbox specifies the number of independent repeat runs for each algorithm on each problem.
+The <No.
+of Results> textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations.
+The <Parallel> button toggles parallel processing, allowing independent experiments to run simultaneously across multiple process.
+The number of concurrent experiments depends on the number of processor cores.
+If the number of parallel experiments exceeds the available cores, MATLAB automatically handles the scheduling.
+For multiple algorithms and problems, each algorithm processes a problem in parallel.
+After one algorithm finishes, the next starts, repeating until all algorithms have processed the problem, then moving to the next problem.
+Hence, the number of algorithms and problems is not inherently limited.
+The drop-down buttons <Task>, <Objective>, and <Special> filter algorithms and problems based on predefined labels, similar to the test module.
+Users can add selected algorithms and problems by clicking the <Add> button after choosing them from the respective drop-down boxes at the bottom.
 In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them.
-Once all configurations are set, clicking the green Start button initiates the experiment execution.
-Users can also pause or stop the experiment using the Pause and Stop buttons if needed.
+Once all configurations are set, clicking the green <Start> button initiates the experiment execution.
+Users can also pause or stop the experiment using the <Pause> and <Stop> buttons if needed.
 Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module.
-Users can read and save the original data of the experiment using the Load Data and Save Data buttons.
+Users can read and save the original data of the experiment using the <Load Data> and <Save Data> buttons.
 In the results area, the first function row contains buttons for operating on the experimental data.
-The first button, labeled Save Table, exports the table results to xlsx, csv, and tex files.
-The subsequent buttons labeled Converge control the convergence y-lim types and generate convergence plots based on the selected results from the table.
-The Pareto button draws the Pareto front and non-dominated solutions based on the selected results.
+The first button, labeled <Save Table>, exports the table results to xlsx, csv, and tex files.
+The subsequent buttons labeled <Converge> control the convergence y-lim types and generate convergence plots based on the selected results from the table.
+The <Pareto> button draws the Pareto front and non-dominated solutions based on the selected results.
 The edit box controls the formatting type of the table result.
@@
 In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line.
-By inputting the algorithm list, problem list, and other parameters into the script file mto.m, experiments can be executed directly from the command line.
-Upon completion, the experiment data MTOData.mat is returned.
+By inputting the algorithm list, problem list, and other parameters into the script file <mto.m>, experiments can be executed directly from the command line.
+For example, running the command mto({MFEA, EMEA}, {CMT1, CMT2}, 30) will execute the MFEA and EMEA algorithms on the CMT1 and CMT2 problems independently 30 times.
+Upon completion, the experiment data <MTOData.mat> is returned.
 Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
@@
 IV-A Adding New Algorithms
-To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the Algorithms/ folder.
-This class file should inherit from the base class Algorithm.m.
+To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder.
+This class file should inherit from the base class <Algorithm.m>.
 Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo.
@@
 IV-B Adding New Problems
-To integrate new problems into MToP, users should create a dedicated problem class file within the Problems/ folder.
-This class file should inherit from the base class Problem.m.
-It’s advisable to define the suggested maximum function evaluations (maxFE) within the constructor of the specific problem.
+To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder.
+This class file should inherit from the base class <Problem.m>.
+It’s advisable to define the suggested maximum function evaluations <maxFE> within the constructor of the specific problem.
 Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function.
@@
 IV-C Adding New Performance Metrics
-To incorporate new performance metrics into MToP, users should create a new metrics function file within the Metrics/ folder.
-Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object.
-For the metrics displayed in the GUI table, ensure that the RowName, ColumnName, and TableData fields are properly populated.
-Typically, the RowName for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task.
-The ColumnName typically represents the algorithm name.
-Implement the calculation function for the specific performance metric in the TableData field.
-For plotting metric convergence, ensure the ConvergeData field is populated.
-In the case of multi-objective optimization, include non-dominated solutions in the ParetoData field.
+To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder.
+Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object <MTOData> and return the calculated <result> object.
+For the metrics displayed in the GUI table, ensure that the <RowName>, <ColumnName>, and <TableData> fields are properly populated.
+Typically, the <RowName> for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task.
+The <ColumnName> typically represents the algorithm name.
+Implement the calculation function for the specific performance metric in the <TableData> field.
+For plotting metric convergence, ensure the <ConvergeData> field is populated.
+In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
+IV-D Extending Optimization Types
+There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking [18].
+Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types.
+Users can refer to the existing code patterns in MToP to develop new optimization types.
+Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems.
+The metric calculation can be adjusted to evaluate the performance of the target task.
+For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
 V Conclusion and Future Work
@@
 It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field.
-Specifically, MToP is considering the incorporation of multitask combinatorial optimization problems [5, 6] and more real-world applications [8, 111] in the future.
+Specifically, MToP is considering the incorporation of many-objective optimization codes [46, 4], multitask combinatorial optimization codes [5, 6] and more real-world applications [9, 47] in the future.
 Acknowledgement

### v4

@@
 MToP: A MATLAB Optimization Platform for Evolutionary Multitasking
-Abstract
-Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past years.
+Abstract.
+Evolutionary multitasking (EMT) has emerged as a popular topic of evolutionary computation over the past decade.
 It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques.
@@
 To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT.
-MToP incorporates over 40 MTEAs, more than 150 MTO problem cases with real-world applications, and over 20 performance metrics.
-Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 40 popular single-task evolutionary algorithms to address MTO problems.
+MToP incorporates over 50 MTEAs, more than 200 MTO problem cases with real-world applications, and over 20 performance metrics.
+Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 50 popular single-task evolutionary algorithms to address MTO problems.
 MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization.
@@
 The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
-Index Terms:
-Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm.
-I Introduction
+Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm
+1. Introduction
 Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency.
 Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature.
-Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization [1, 2], multi-objective optimization [3, 4], and combinatorial optimization [5, 6].
-In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking [7].
+Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021b), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021c, a).
+In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022).
 A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
-$\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,F_{K}(\bm{x}_{K})\Big{]},\\ \textit{s.t.
-}\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$ (1)
-where $\Omega_{k}$ and $F_{k}$ are the decision space and objective function of the $k$-th task.
-Note that for a multi-objective optimization task, the $F_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task.
-For task $k$ with upper and lower bounds $L_{k}$ and $U_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space [8] as follows:
-$\bm{x}^{\prime}=\frac{\bm{x}-L_{k}}{U_{k}-L_{k}}.$ (2)
+(1) $\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,\bm{F}_{k}(\bm{x}_{K})\Big{]},\\ \textit{s.t.
+}\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$
+where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task.
+Note that for a multi-objective optimization task, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task.
+For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
+(2) $\bm{x}^{\prime}=\frac{\bm{x}-\bm{L}_{k}}{\bm{U}_{k}-\bm{L}_{k}}.$
 In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$.
 During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
-EMT has found successful applications in various domains, including engineering scheduling [5, 9], nonlinear equation systems [10], feature selection in machine learning [11], anomaly detection [12], point cloud registration [13, 14], and reinforcement learning [15].
-To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques [16, 17].
-Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources [18, 19].
-The first attempt of EMT can be traced back to the multifactorial EA [8], which introduced an implicit knowledge representation via random mating among optimization tasks.
+EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021c; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024e), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
+To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018).
+Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024f).
+The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks.
 Subsequently, numerous EAs tailored for MTO have emerged.
-These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks [20, 21], or a multi-population framework allocating separate populations for each task [22, 23].
-Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space [8, 21], autoencoding [24, 25], affine transformation [26, 27, 28], and adversarial generative models [29] have been proposed in MTEAs.
-Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment [20, 30, 31], knowledge selection [23, 32], and historical feedback [33, 34, 35], have also been investigated.
+These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020; Wang et al., 2022a).
+Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), autoencoding (Feng et al., 2019b; Zhou et al., 2021a), affine transformation (Xue et al., 2022; Lin et al., 2024; Gao et al., 2022), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs.
+Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023a; Wu et al., 2024c), knowledge selection (Wang et al., 2022a; Liang et al., 2022c), and historical feedback (Li et al., 2022c; Lin et al., 2021; Tang et al., 2021b), have also been investigated.
 Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes.
@@
 In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization.
-For instance, PlatEMO [36] is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems.
-EvoX [37] is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training.
-Evo-Panel [38] created a comprehensive dynamic visualization tool to illustrate the detailed procedures of different optimization algorithms in solving numerical benchmark functions.
-EDOLAB [39] is an optimization tool specifically for dynamic optimization.
+For instance, PlatEMO (Tian et al., 2017) is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems.
+EvoX (Huang et al., 2024) is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training.
+EDOLAB (Peng et al., 2023) is an optimization tool specifically for dynamic optimization.
 However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks.
@@
 These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms.
-Moreover, EMT has been extended to the subfields of many-task optimization [40, 41], multi-objective multitask optimization [42, 43], competitive multitask optimization [33, 44], and constrained multi-task optimization [45].
+Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019), multi-objective MTO (Gupta et al., 2017; Gao et al., 2023), competitive MTO (Li et al., 2022c, 2023a), and constrained multi-task optimization (Li et al., 2022a).
 Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
@@
 The main contributions of MToP are summarized as follows:
-  1. 1.
+  1. (1)
     MToP features a user-friendly graphical user interface (GUI) comprising test, experiment, and data-process modules.
@@
 On top of these, it provides an extensive public application programming interface (API) with template functions for population initialization, function evaluation, evolutionary operators, and environmental selection.
-  2. 2.
+  2. (2)
     MToP encompasses a wide array of algorithms, problems, and metrics, all accessible via a public API.
-Over 40 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types.
-Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 40 single-task EAs of diverse types.
-In terms of synthetic test problems, MToP incorporates over 150 benchmark MTO problems alongside several real-world applications.
+Over 50 MTEAs are implemented, catering to single-objective, multi-objective, constrained, many-task, and competitive multitask types.
+Additionally, to facilitate comparative analyses between MTEAs and popular traditional EAs, MToP integrates more than 50 single-task EAs of diverse types.
+In terms of synthetic test problems, MToP incorporates over 200 benchmark MTO problems alongside several real-world applications.
 Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
-  3. 3.
+  3. (3)
     MToP is designed to be easily extended, allowing for the seamless addition of new algorithms, problems, and metrics.
@@
 Given its status as a completely open-source project, researchers have the opportunity to leverage existing algorithms, problems, and metrics as the foundation for novel ideas.
-Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research [^1] MToP has been continuously updated since Sep.
-21, 2021, when it was uploaded to GitHub at https://github.com/intLyc/MTO-Platform..
+Through the collaborative platform GitHub, MToP undergoes continuous updates and enhancements, ensuring that it remains at the forefront of EMT research.
 The rest of this paper is organized as follows.
-Section II elaborates on the architecture of MToP.
-Section III describes approaches for performing experiments with MToP.
-Section IV introduces how to extend the contents of MToP.
-Finally, the conclusion and future work are given in Section V.
-II Architecture of MToP
+Section 2 elaborates on the architecture of MToP.
+Section 3 describes approaches for performing experiments with MToP.
+Section 4 introduces how to extend the contents of MToP.
+Finally, the conclusion and future work are given in Section 5.
+2. Architecture of MToP
 MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis.
@@
 Following standardized code patterns, MToP integrates various algorithms and performance metrics.
-II-A Functional Modules
-II-A1 Test Module
+2.1. Functional Modules
+2.1.1. Test Module
 (a) Function landscapes (1D)
@@
 (e) Metric convergence
-Figure 1: Examples of graphical display in the test module of MToP.
+Figure 1.
+Examples of graphical display in the test module of MToP.
 (a) and (b) illustrate the landscapes of single-objective problems with different tasks in the one- and two-dimensional unified search space, respectively.
@@
 (l) Gen=1000
-Figure 2: Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations.
+Figure 2.
+Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations.
 The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
@@
 In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
-Figure 3: The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
+Figure 3. The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
 The workflow of API calls for executing an algorithm on a problem is depicted in Fig.
@@
 Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
-II-A2 Experiment Module
-TABLE I: Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
+2.1.2. Experiment Module
+Table 1. Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
 GA MFEA MTEA-AD
@@
 (b) Pareto front
-Figure 4: schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
+Figure 4. schematic plotting by the experiment module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
 The experiment module offers functionalities for conducting comprehensive experiments and analyzing results.
@@
 Significance test results are also included in the table alongside the metric results.
-All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table I.
+All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table 1.
 Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig.
@@
 4 (b).
-II-A3 Data Process Module
-TABLE II: Properties of experimental data.
+2.1.3. Data Process Module
+Table 2. Properties of experimental data.
 Property Description
@@
 Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>.
-The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table II.
+The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table 2.
 This saved data can subsequently be reloaded by the experiment module of MToP.
@@
 This process can also be used to merge datasets by independent runs or problems.
-II-B File Structure
-Figure 5: File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
+2.2. File Structure
+Figure 5. File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
 In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
@@
 Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
-II-C Code Pattern
-II-C1 Algorithm
-TABLE III: Properties and methods of algorithm base class.
+2.3. Code Pattern
+2.3.1. Algorithm
+Table 3. Properties and methods of algorithm base class.
 Property or method Description
@@
 All algorithms within MToP inherit from the algorithm base class <Algorithm.m>.
-The properties and methods of this base class are detailed in Table III.
+The properties and methods of this base class are detailed in Table 3.
 This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself.
-The MTEAs implemented in MToP are outlined in Table LABEL:supp:tab:algo-mt of the supplementary files.
+The MTEAs implemented in MToP are outlined in Table 4.
 Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems.
-Specific algorithms belonging to this category are listed in Table LABEL:supp:tab:algo-st.
+Specific algorithms belonging to this category are listed in Table 5.
 Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively.
@@
 The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
-II-C2 Problem
-TABLE IV: Properties and methods of problem base class.
+Table 4.
+Multitask evolutionary algorithms included in MToP.
+The table lists the algorithm name, optimization objective, task type, framework, constraint handling, and a brief description.
+SO and MO denote single-objective and multi-objective optimization, respectively.
+MT and MaT represent multi-task and many-task, respectively.
+MF and MP represent multifactorial and multi-population, respectively.
+Algorithm Objective Task Framework Constraint Description
+MFEA (Gupta et al., 2016) SO MT MF Multifactorial evolution-based GA for MTO
+LDA-MFEA (Bali et al., 2017) SO MT MF Linearized domain adaptation in MFEA
+MFDE&MFPSO (Feng et al., 2017) SO MT MF Multifactorial evolution-based DE and PSO
+G-MFEA (Ding et al., 2019) SO MT MF Generalized MFEA with decision variable translation and shuffling
+MFEA-GHS (Liang et al., 2019) SO MT MF Genetic Transform and hyper-rectangle search in MFEA
+MFEA-DV (Yin et al., 2019) SO MT MF Enhanced MFEA with cross-task search direction
+MFEA-II (Bali et al., 2020) SO MT MF MFEA with online transfer parameter estimation
+SREMTO (Zheng et al., 2020) SO MT MF Self-regulated multitask framework for varying relatedness among tasks
+TLTLA (Ma et al., 2020) SO MT MF MFEA with two-level inter- and intra-task transfer learning
+MFEA-AKT (Zhou et al., 2021b) SO MT MF MFEA with adaptive knowledge transfer via multiple crossover operator
+ASCMFDE (Tang et al., 2021a) SO MT MF Intertask transfer in aligned subspace for MFDE
+AT-MFEA (Xue et al., 2022) SO MT MF Affine transformation-enhanced domain adaptation for MFEA
+MFEA-DGD (Liu et al., 2023) SO MT MF MFEA based on diffusion gradient descent
+MFEA-VC (Wang et al., 2024) SO MT MF MFEA with contrastive variational auto-encoder
+IMEA (Hashimoto et al., 2018) SO MT MP Multi-population framework for MTO through island model
+EMEA (Feng et al., 2019b) SO MT MP Explicit knowledge transfer for MTO via autoencoding
+MFMP (Li et al., 2020) SO MT MP Multi-population-based adaptive DE for MTO
+DEORA (Li et al., 2022c) SO MT MP Adaptive task selection for competitive MTO
+MKTDE (Li et al., 2022b) SO MT MP Meta-knowledge transfer-based multitask DE
+MTES (Bai et al., 2022) SO MT MP Multitask OpenAI-ES via gradient-free evolution multitasking
+BLKT-DE (Jiang et al., 2023b) SO MT MP Multitask DE with block-level knowledge transfer
+MTSRA (Li et al., 2023a) SO MT MP Improved adaptive DE with competitive task selection
+MTDE-ADKT (Zhang et al., 2024a) SO MT MP Multitask DE with adaptive dual knowledge transfer
+CEDA (Zhang et al., 2024b) SO MT MF/MP Constrained multitasking via co-evolution and domain adaptation
+CMO-LKT (Ban et al., 2025) SO MT MP Local knowledge transfer-based EA for constrained MTO
+MTEA-AD (Wang et al., 2022a) SO MT/MaT MP MTEA with adaptive knowledge transfer via anomaly detection
+MTES-KG (Li et al., 2024c) SO MT/MaT MP Multitask ES with knowledge-guided external sampling
+MTEA-HKTS (Zhao et al., 2024) SO MT/MaT MP MTEA for solving the problem of transfer targets
+BoKT (Jiang et al., 2023a) SO MaT MP Bi-objective knowledge transfer framework for many-task optimization
+TRADE (Wu et al., 2023) SO MaT MP Transferable adaptive parameter DE for many-task optimization
+TNG-NES (Li et al., 2024b) SO MaT MP Transfer task-averaged natural gradient for efficient many-task optimization
+MTEA-SaO (Li et al., 2023b) SO/MO MT/MaT MP Adaptive solver multitask framework with implicit knowledge transfer
+KR-MTEA (Cui et al., 2023) SO/MO MT/MaT MP Adaptive multi-task EA based on knowledge reuse
+SBO (Liaw and Ting, 2019) SO/MO MaT MP Symbiosis in biocoenosis framework for many-task optimization
+MaTDE (Chen et al., 2020) SO/MO MaT MP Many-task DE with adaptive archive-based knowledge transfer
+EMaTO-MKT (Liang et al., 2022c) SO/MO MaT MP Multi-source knowledge transfer via local distribution estimation
+MO-MFEA (Gupta et al., 2017) MO MT MF Multiobjective MFEA with non-dominated sort and crowding distance
+MO-MFEA-II (Bali et al., 2021) MO MT MF Cognizant Multitasking for parameter estimation in MO-MFEA
+EMT-PD (Liang et al., 2022b) MO MT MF Two-stage adaptive knowledge transfer based on population distribution
+EMT-GS (Liang et al., 2023) MO MT MF Generative adversarial networks for knowledge transfer
+MM-DE (Chen et al., 2018) MO MT MP Fast memetic DE for multiobjective MTO
+AMT-NSGA-II (Da et al., 2019) MO MT MP Curbing negative influences online for seamless transfer
+EMT-ET (Lin et al., 2021) MO MT MP Effective knowledge transfer approach via non-dominated sort
+MOMFEA-SADE (Liang et al., 2022a) MO MT MP Subspace alignment and adaptive Differential Evolution
+MTEA-D-DN (Wang et al., 2023) MO MT MP Neighborhood as a bridge for decomposition-based knowledge transfer
+MTEA-D-TSD  (Li et al., 2024a) MO MT MP Transfer search directions among decomposed subtasks
+MTDE-MKTA (Li and Gong, 2025) MO MT MP Multitask DE with multiple knowledge types and transfer adaptation
+RVC-MTEA (Li et al., 2024g) MO MT MP Reference vector contribution based MTEA for competitive multiobjective MTO
+MTEA-DCK (Li et al., 2025) MO MT MP Multiobjective MTEA with diversity- and convergen-based knowledge transfer
+Table 5.
+Single-task evolutionary algorithms for multitask optimization included in MToP.
+The table lists the algorithm name, optimization objective, task type, category, constraint handling, and a brief description.
+SO and MO denote single-objective and multi-objective optimization, respectively.
+GA denotes genetic algorithm, DE denotes differential evolution, SI denotes swarm intelligence, ES denotes evolution strategy, CH denotes constraint handling, ND denotes non-dominated sorting, ID denotes indicator-based, DC denotes decomposition, and RV denotes reference vector.
+Algorithms Objective Task Category Constraint Description
+GA (Zbigniew, 1996) SO ST/MT GA Genetic algorithm with SBX crossover and polynomial mutation
+Global-GA (Bull and Liu, 2024) SO ST/MT GA GA with global crossover to replace cooperative coevolution
+DE (Storn, 1996) SO ST/MT DE Differential evolution algorithm with DE/rand/1/bin operator
+jDE (Brest et al., 2006) SO ST/MT DE Self-adapting control parameters in DE
+JADE (Zhang and Sanderson, 2009) SO ST/MT DE Adaptive DE with fast convergence performance
+rank-DE (Gong and Cai, 2013) SO ST/MT DE Ranking-based mutation operators for DE
+SHADE (Tanabe and Fukunaga, 2013) SO ST/MT DE Success-history based parameter adaptation for DE
+L-SHADE (Tanabe and Fukunaga, 2014) SO ST/MT DE Linear population size reduction for SHADE
+LSHADE44 (Poláková, 2017) SO ST/MT DE L-SHADE with competing strategies applied to constrained optimization
+CAL-SHADE (Zamuda, 2017) SO ST/MT DE Adaptive constraint handling technique for L-SHADE
+jSO (Brest et al., 2017) SO ST/MT DE Champion of CEC 2017 numerical optimization competition via improved DE
+NL-SHADE-RSP (Stanovov et al., 2021) SO ST/MT DE Champion of CEC 2021 numerical optimization competition via improved DE
+PSO (Kennedy and Eberhart, 1995) SO ST/MT SI Particle swarm optimization with global and particle best update
+CSO (Cheng and Jin, 2015) SO ST/MT SI Competitive swarm optimizer for large-scale optimization
+MPA (Faramarzi et al., 2020a) SO ST/MT SI Marine predators algorithm for numerical optimization
+EO (Faramarzi et al., 2020b) SO ST/MT SI Equilibrium optimizer for numerical optimization
+AO (Abualigah et al., 2021) SO ST/MT SI Aquila optimizer for numerical optimization
+KLDE&KLPSO (Jiang et al., 2023c) SO ST/MT DE/SI Knowledge learning-based DE and PSO for numerical optimization
+CMA-ES (Hansen and Ostermeier, 2001) SO ST/MT ES Evolution strategy with derandomized covariance matrix adaptation
+IPOP-CMA-ES (Auger and Hansen, 2005) SO ST/MT ES Restart CMA-ES with increasing population size
+sep-CMA-ES (Ros and Hansen, 2008) SO ST/MT ES Separable CMA-ES for large-scale optimization
+xNES-as (Schaul, 2012) SO ST/MT ES Natural ES with adaptation sampling
+R1-NES (Sun et al., 2013) SO ST/MT ES A Linear Time natural ES for non-separable functions
+xNES&SNES (Wierstra et al., 2014) SO ST/MT ES ES with adaptive natural gradients to update distribution
+OpenAI-ES (Salimans et al., 2017) SO ST/MT ES Parallelized ES with standard normal distribution gradients
+DES (Arabas and Jagodziński, 2020) SO ST/MT ES Matrix-free covariance matrix adaptation ES
+MTV-DE (Mezura-Montes et al., 2007) SO ST/MT CH Multiple trial vectors in DE for handling constraints
+ECHT (Mallipeddi and Suganthan, 2010b) SO ST/MT CH Ensemble of constraint handling techniques
+FROFI (Wang et al., 2016) SO ST/MT CH Incorporating objective function information into the feasibility rule
+C2oDE (Wang et al., 2019) SO ST/MT CH Composite DE for constrained optimization
+CORCO (Wang et al., 2020) SO ST/MT CH Utilizing the correlation between constraints and objective function
+DeCODE (Wang et al., 2021) SO ST/MT CH Decomposition-based multiobjective approach for constrained optimization
+VMCH (Wu et al., 2022) SO ST/MT CH Voting-mechanism-based ensemble of constraint handling techniques
+CEDE-DRL (Hu et al., 2023) SO ST/MT CH Deep reinforcement learning assisted co-evolutionary DE
+CCEF-ECHT (Li et al., 2024d) SO ST/MT CH Competitive and cooperative ensemble of constraint handling techniques
+SPEA2 (Zitzler et al., 2001) MO ST/MT ND Improving the strength Pareto approach for multiobjective optimization
+NSGA-II (Deb et al., 2002) MO ST/MT ND Multiobjective GA with non-dominated sort and crowding distance
+NSGA-III (Deb and Jain, 2014) MO ST/MT ND Reference-point-based nondominated sorting for multiobjective optimization
+SMS-EMOA (Beume et al., 2007) MO ST/MT ID Hypervolume-based selection for multiobjective optimization
+MO-CMA-ES (Igel et al., 2007) MO ST/MT ES Covariance matrix adaptation for multiobjective optimization
+MOEA/D (Zhang and Li, 2007) MO ST/MT DC Decomposition-based multiobjective evolutionary algorithm
+MOEA/D-DE (Li and Zhang, 2009) MO ST/MT DC MOEA/D for complex Pareto sets in multiobjective optimization
+RVEA (Cheng et al., 2016) MO ST/MT RV Reference vector-guided EA for many-objective optimization
+AR-MOEA (Tian et al., 2018) MO ST/MT ID Adaptive reference points for MOEA/D in multiobjective optimization
+MSEA (Tian et al., 2021a) MO ST/MT ND Multistage EA for better diversity preservation in multiobjective optimization
+CCMO (Tian et al., 2021b) MO ST/MT CH Coevolutionary framework for constrained multiobjective optimization
+LMOCSO (Tian et al., 2020) MO ST/MT SI Efficient large-scale multiobjective optimization based on CSO
+CMOCSO (Ming et al., 2023) MO ST/MT SI Competitive and cooperative swarm optimizer for multiobjective optimization
+2.3.2. Problem
+All problems within MToP inherit from the problem base class <Problem.m>.
+The properties and methods of <Problem.m> are detailed in Table 7.
+As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach (Gupta et al., 2016).
+It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
+A simple problem implementation example is illustrated in Listing LABEL:list:prob.
+The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor.
+The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task.
+Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task.
+Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task.
+For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type.
+Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>.
+Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
+With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP.
+The existing problems available in MToP are enumerated in Table 6.
+In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
+Table 6.
+Benchmark and real-world optimization problems included in MToP.
+The table lists the problem name, case number, task type, dimension, optimization objective, task type, special case, and a brief description.
+SO and MO denote single-objective and multi-objective optimization, respectively.
+ST, MT and MaT represent single-task, multi-task and many-task, respectively.
+Problem Case Task Dimension Objective Task Special Description
+CEC17-MTSO (Da et al., 2017) 9 2 50 SO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
+CEC19-MaTSO (Feng et al., 2019a) 6 2-50 50 SO MaT Single-objective many-task optimization problems
+WCCI20-MTSO (Feng et al., 2020) 10 2 50 SO MT Complex single-objective MTO problems
+WCCI20-MaTSO (Feng et al., 2020) 10 2-50 50 SO MaT Complex single-objective many-task optimization problems
+CMT (Li et al., 2022a) 9 2 Any SO MT Constrained Large/partial/no feasible regions interaction and different inter-task similarity
+C2TOP&C4TOP (Li et al., 2022c) 51 2/4 50 SO MT Competitive Different optimal objective values among competitive tasks
+C-CPLX (Li et al., 2023a) 20 2 50 SO MT Competitive Complex competitive MTO problems
+LSMaTSO (Li et al., 2024b) 5 2-50 300 SO MaT Large-scale many-task optimization problems
+CEC17-MTMO (Yuan et al., 2017) 9 2 50 MO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
+CEC19-MTMO (Feng et al., 2019a) 10 2 50 MO MT Complex multi-objective MTO problems
+CEC19-MaTMO (Feng et al., 2019a) 10 2-50 50 MO MaT Multi-objective many-task optimization problems
+WCCI20-MaTMO (Feng et al., 2020) 10 2-50 50 MO MaT Complex multi-objective many-task optimization problems
+CEC21-MTMO (Feng et al., 2021b) 10 2 50 MO MT More complex multi-objective multi optimization problems
+CMOMT (Li et al., 2024g) 12 2/3 30-60 MO MT Competitive multiobjective MTO problems
+Synthetic Functions 9 1 Any SO ST Ackley, Elliptic, Griewank, Rastrigin, Rosenbrock Schwefel, Sphere, Weierstrass
+CEC06-CSO (Liang et al., 2006) 24 1 2-24 SO ST Constrained CEC 2006 competition on constrained single-objective optimization
+CEC10-CSO (Mallipeddi and Suganthan, 2010a) 18 1 10/30 SO ST Constrained CEC 2010 competition on constrained single-objective optimization
+CEC17-CSO (Wu et al., 2017) 28 1 2-100 SO ST Constrained CEC 2017 competition on constrained single-objective optimization
+CEC17-SO (Awad et al., 2017) 29 1 2-100 SO ST CEC 2017 competition on bound constrained single-objective optimization
+CEC20-RWCO (Kumar et al., 2020) 57 1 2-158 SO ST Constrained CEC 2020 competition on real-world constrained single-objective optimization
+CEC22-SO (Kumar et al., 2022) 12 1 10/20 SO ST CEC 2022 competition on bound constrained single-objective optimization
+PEPVM (Li et al., 2023b) 1 3 7 SO MT Parameter extraction of photovoltaic models as MTO problem
+SCP (Li et al., 2022c) 1 Any 60-120 SO MT Competitive Sensor coverage problem as competitive MTO problem
+OPF (Li et al., 2024c) 5 2 24/33 SO MT Constrained Optimal power flow as constrained MTO problem
+PKACP (Jiang et al., 2023a) 1 Any Any SO MaT Planar kinematic arm control problem as many-task optimization problem
+MO-SCP (Li et al., 2023b) 2 Any 60-120 MO MT Competitive Multiobjective sensor coverage problem as competitive multiobjective MTO problem
+SOPM (Li and Gong, 2025) 2 3 30 MO MT Constrained Synchronous optimal pulse-width modulation as constrained multiobjective MTO problem
+MO-OPF (Li and Gong, 2025) 4 2 24/33 MO MT Constrained Multiobjective optimal power flow as constrained multiobjective MTO problem
+OPF-CMTMO (Li et al., 2024g) 4 2 11/24 MO MT Competitive Competitive multitask optimal power flow with thermal and wind-solar power
+Table 7. Properties and methods of problem base class.
 Property or method Description
@@
 35end
-All problems within MToP inherit from the problem base class <Problem.m>.
-The properties and methods of <Problem.m> are detailed in Table IV.
-As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach [8].
-It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
-A simple problem implementation example is illustrated in Listing LABEL:list:prob.
-The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor.
-The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task.
-Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task.
-Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task.
-For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type.
-Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>.
-Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
-With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP.
-The existing problems available in MToP are enumerated in Table LABEL:supp:tab:prob-mt of the supplementary files.
-In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
-II-C3 Performance Metric
+2.3.3. Performance Metric
 1 label={list:metric},
@@
 Note that specific metric calculation functions are not elaborated here for the sake of simplicity.
-The metrics currently integrated into MToP are enumerated in Table V.
-TABLE V: Metrics included in MToP.
+The metrics currently integrated into MToP are enumerated in Table 8.
+Table 8.
+Metrics included in MToP.
 The table lists the metric name, optimization objective, task type, and a brief description.
@@
 IGD (AV) MO MT Average IGD for all tasks
-IGD+ (MTS) MO MT Multitask score of IGD for all tasks
+IGD (MTS) MO MT Multitask score of IGD for all tasks
 IGD (CMT) MO MT Competitive multitask IGD for all tasks
@@
 Run Time SO/MO ST/MT Algorithm running time for all tasks
-III Performing Experiments with MToP
+3. Performing Experiments with MToP
 (a) Test Module
 (b) Experiment Module
-Figure 6: Graphical user interface of MToP.
+Figure 6.
+Graphical user interface of MToP.
 The test module (a) is used for testing algorithms and problems.
@@
 In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
-III-A Testing Algorithms and Problems
+3.1. Testing Algorithms and Problems
 As depicted in Fig.
@@
 Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
-III-B Executing Comparative Experiments
+3.2. Executing Comparative Experiments
 The GUI of the experiment module is displayed in Fig. 6 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
@@
 The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
-III-C Executing via Alternative Command Line
+3.3. Executing via Alternative Command Line
 In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line.
@@
 Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
-IV Extending MToP
+4. Extending MToP
 Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP.
-By adhering to the code pattern outlined in Section II-C, users can develop and implement their code files with ease.
-IV-A Adding New Algorithms
+By adhering to the code pattern outlined in Section 2.3, users can develop and implement their code files with ease.
+4.1. Adding New Algorithms
 To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder.
@@
 Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
-IV-B Adding New Problems
+4.2. Adding New Problems
 To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder.
@@
 Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
-IV-C Adding New Performance Metrics
+4.3. Adding New Performance Metrics
 To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder.
@@
 In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
-IV-D Extending Optimization Types
-There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking [18].
+4.4. Extending Optimization Types
+There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking (Tan et al., 2021).
 Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types.
@@
 For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
-V Conclusion and Future Work
+5. Conclusion and Future Work
 This paper introduces MToP, an open-source MATLAB platform designed for EMT.
 MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures.
-The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT.
+The current version of MToP contains over 50 MTEAs, more than 50 single-task EAs capable of handling MTO problems, over 200 benchmark MTO problems, and several real-world applications of EMT.
 This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
@@
 It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field.
-Specifically, MToP is considering the incorporation of many-objective optimization codes [46, 4], multitask combinatorial optimization codes [5, 6] and more real-world applications [9, 47] in the future.
-Acknowledgement
+Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021c, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
+Acknowledgements.
+This work was partly supported by the National Natural Science Foundation of China under Grant No.
+62076225 and 62403477, the Open Research Fund of the State Key Lab of Digital Manufacturing Equipment & Technology under Grant No.
+IMETKF2024020, and the Major Project of Hubei Province Science and Technology under Grant No.
+2023BCA006.
 All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles.
@@
 The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes.
-Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO [36], for which the authors are also thankful.
+Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.

### v5

@@
-MToP: A MATLAB Optimization Platform for Evolutionary Multitasking
+MToP: A MATLAB Benchmarking Platform for Evolutionary Multitasking
 Abstract.
@@
 It aims to concurrently address multiple optimization tasks within limited computing resources, leveraging inter-task knowledge transfer techniques.
-Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications.
-To bridge this gap, we introduce the first open-source optimization platform, named MTO-Platform (MToP), for EMT.
+Despite the abundance of multitask evolutionary algorithms (MTEAs) proposed for multitask optimization (MTO), there remains a need for a comprehensive software platform to help researchers evaluate MTEA performance on benchmark MTO problems as well as explore real-world applications.
+To bridge this gap, we introduce the first open-source benchmarking platform, named MToP, for EMT.
 MToP incorporates over 50 MTEAs, more than 200 MTO problem cases with real-world applications, and over 20 performance metrics.
+Based on these, we provide benchmarking recommendations tailored for different MTO scenarios.
 Moreover, to facilitate comparative analyses between MTEAs and traditional evolutionary algorithms, we adapted over 50 popular single-task evolutionary algorithms to address MTO problems.
-MToP boasts a user-friendly graphical interface, facilitating results analysis, data export, and schematics visualization.
+Notably, we release extensive pre-run experimental data on benchmark suites to enhance reproducibility and reduce computational overhead for researchers.
+MToP features a user-friendly graphical interface, facilitating results analysis, data export, and schematic visualization.
 More importantly, MToP is designed with extensibility in mind, allowing users to develop new algorithms and tackle emerging problem domains.
-The source code of MToP is available at https://github.com/intLyc/MTO-Platform.
-Evolutionary multitasking, MATLAB optimization platform, multitask optimization problem, evolutionary algorithm
+The source code of MToP is available at: https://github.com/intLyc/MTO-Platform
+Evolutionary multitasking, benchmarking platform, multitask optimization problem, evolutionary algorithm
 1. Introduction
 Evolutionary computation (EC), inspired by natural evolution, has experienced rapid growth owing to its effectiveness and efficiency.
-Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and user-friendly nature.
-Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021b), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021c, a).
-In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022).
-A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ optimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
-(1) $\begin{gathered}(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})=\arg\min\Big{[}F_{1}(\bm{x}_{1}),F_{2}(\bm{x}_{2}),...,\bm{F}_{k}(\bm{x}_{K})\Big{]},\\ \textit{s.t.
-}\bm{x}_{k}\in\Omega_{k},\ k=1,2,...,K,\\ \end{gathered}$
+Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and ease of implementation.
+Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang2020CORCO; Tian2021CCMO), multi-objective optimization (Zhang2007MOEA-D; Deb2014NSGA3), and combinatorial optimization (Feng2021VRPHTO; Feng2021EEMTA).
+In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei2021Review-MTO).
+A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ minimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
+[…]
 where $\Omega_{k}$ and $\bm{F}_{k}$ are the decision space and objective function of the $k$-th task.
-Note that for a multi-objective optimization task, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of non-dominated solutions for the $k$-th task.
-For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
+Note that for a multi-objective multi-task optimization problem, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of Pareto sets for the $k$-th task.
+For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta2016MFEA) as follows:
 (2) $\bm{x}^{\prime}=\frac{\bm{x}-\bm{L}_{k}}{\bm{U}_{k}-\bm{L}_{k}}.$
 In addition, the dimensionality of task $k$ is expanded to $\max(D_{1},...,D_{K})$.
-During real function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
-EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021c; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024e), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
-To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018).
-Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024f).
-The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks.
+During objective function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
+EMT has found successful applications in various domains, including engineering scheduling (Feng2021VRPHTO; Gupta2021Half), nonlinear equation systems (Li2024Evolutionary), feature selection in machine learning (Chen2021Evolutionary), anomaly detection (Wang2022AUC), point cloud registration (Wu2024Evolutionary; Wu2023Evolutionary), and reinforcement learning (Zhang2023Multitask).
+To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan2023Knowledge; Gupta2018Insights).
+Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan2021ETO; Li2024MFEA-RL).
+The first attempt of EMT can be traced back to the multifactorial EA (Gupta2016MFEA), which introduced an implicit knowledge representation via random mating among optimization tasks.
 Subsequently, numerous EAs tailored for MTO have emerged.
-These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020; Wang et al., 2022a).
-Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), autoencoding (Feng et al., 2019b; Zhou et al., 2021a), affine transformation (Xue et al., 2022; Lin et al., 2024; Gao et al., 2022), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs.
-Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023a; Wu et al., 2024c), knowledge selection (Wang et al., 2022a; Liang et al., 2022c), and historical feedback (Li et al., 2022c; Lin et al., 2021; Tang et al., 2021b), have also been investigated.
+These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali2020MFEA2; Zhou2021MFEA-AKT), or a multi-population framework allocating separate populations for each task (Li2020MFMP; Li2023MTEA-SaO; Li2024MTES-KG).
+Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta2016MFEA; Zhou2021MFEA-AKT), auto-encoding (Feng2019EMEA; Zhou2021Learnable; Gu2025MTEA-PAE), affine transformation (Xue2022AT-MFEA; Lin2024Ensemble), and adversarial generative models (Liang2023EMT-GS) have been proposed in MTEAs.
+Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali2020MFEA2; Jiang2023BoKT), knowledge selection (Wang2022MTEA-AD; Liang2022EMaTO-MKT; Li2025MTDE-MKTA), and historical feedback (Li2022CompetitiveMTO; Lin2021EMT-ET; Li2023MTSRA), have also been investigated.
 Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes.
 This presents challenges for newcomers entering the field of EMT and for researchers seeking to conduct convenient experimental comparisons of algorithms.
-As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs.
-Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field.
-Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field.
-In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization.
-For instance, PlatEMO (Tian et al., 2017) is tailored for evolutionary multi-objective optimization and benchmarking multi-objective EAs, aiding researchers in designing EAs to address various types of multi-objective optimization problems.
-EvoX (Huang et al., 2024) is a distributed GPU-accelerated library that assists researchers in designing parallel EAs to expedite the resolution of complex optimization problems and reinforcement learning training.
-EDOLAB (Peng et al., 2023) is an optimization tool specifically for dynamic optimization.
-However, unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks.
-Especially important, the implementation of MTEAs requires incorporating inter-task solution space mapping and diverse knowledge transfer techniques.
-Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics.
-Specifically, single-task metrics evaluate the performance of each task in the MTO problem, while multitask metrics assess the overall performance of the MTO problem.
-These bring uncertainty to the pattern of results display and analysis.
-These distinctive requirements present significant challenges for the implementation of MTEAs in existing tools and platforms.
-Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019), multi-objective MTO (Gupta et al., 2017; Gao et al., 2023), competitive MTO (Li et al., 2022c, 2023a), and constrained multi-task optimization (Li et al., 2022a).
-Consequently, researchers encounter difficulties in benchmarking MTEAs across various types of MTO problems.
-To mitigate these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field.
+To overcome these challenges, we present MToP, an open-source MATLAB platform tailored for advancing the EMT field.
 MToP is designed to provide a comprehensive software platform for researchers to evaluate MTEAs on benchmark MTO problems and explore real-world applications.
-The reason for using MATLAB is that it is a widely used programming language in the field of EC.
-Moreover, MATLAB provides a user-friendly graphical interface and a powerful numerical computing environment, making it an ideal choice for developing MToP.
-More importantly, the existing MTEAs and MTO problems are mainly implemented in MATLAB, which can be directly integrated into MToP.
+The selection of MATLAB is based on several strategic factors.
+First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta2016MFEA) and MFEA-II (Bali2020MFEA2), were developed and open-sourced in MATLAB.
+This established a strong precedent and a rich ecosystem of existing code for the community.
+Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian2017PlatEMO), provides a high-quality open-source reference.
+This synergy allows for the adaptaion and reuse of well-vetted modules, such as those for multi-objective optimization, which is crucial for the multi-objective MTO subfield.
+Finally, MATLAB’s powerful numerical computing environment and its capabilities for building user-friendly graphical interfaces make it an ideal choice for a comprehensive platform.
 The main contributions of MToP are summarized as follows:
@@
 Lastly, MToP features a variety of multitask and single-task performance metrics, such as multitask score, objective value, hypervolume, and running time, providing comprehensive evaluation capabilities.
+We also release extensive pre-run experimental data to enhance reproducibility and reduce computational overhead for researchers.
   3. (3)
@@
 The rest of this paper is organized as follows.
+Section 2 reviews related software platforms in EC and justifies the need for MToP.
 Section 3 elaborates on the architecture of MToP.
-Section 3 describes approaches for performing experiments with MToP.
-Section 4 introduces how to extend the contents of MToP.
-Finally, the conclusion and future work are given in Section 5.
+Section 4 provides guidelines for using MToP.
+Section 5 presents the experimental validation of MToP.
+Finally, Section 6 gives discussions and outlook of MToP.
+2. Related Work
+As the field of EMT continues to gain momentum, there is an urgent need for a convenient and user-friendly software platform to facilitate the benchmarking of MTEAs.
+Furthermore, accessible source code and platforms are indispensable for exploring the real-world applications of the EMT field.
+Open-source and user-friendly software platforms play a crucial role in fostering the advancement of a research field.
+In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization:
+  • PlatEMO (Tian2017PlatEMO): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
+  • EDOLAB (Peng2023EDOLAB): A MATLAB-based platform focusing on evolutionary dynamic optimization.
+  • IOHprofiler (Doerr2018IOHprofiler): A modular framework, including IOHexperimenter (deNobel2024IOHexperimenter) and IOHanalyzer (Wang2022IOHanalyzer), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
+  • EvoX (Huang2024EvoX): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
+  • PyPop7 (Duan2024PyPop7): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
+  • MetaBox (Ma2023MetaBox): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
+Table 1. Comparison of core focus areas among popular evolutionary computation platforms.
+Platform Language Core Focus MTO Support
+PlatEMO (Tian2017PlatEMO) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
+EDOLAB (Peng2023EDOLAB) MATLAB Evolutionary Dynamic Optimization No
+IOHprofiler (Doerr2018IOHprofiler) C++/Python/R Benchmarking Iterative Optimization Heuristics No
+EvoX (Huang2024EvoX) Python Distributed GPU-accelerated Optimization No
+PyPop7 (Duan2024PyPop7) Python Large-scale Black-box Optimization No
+MetaBox (Ma2023MetaBox) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
+MToP (This work) MATLAB Evolutionary Multitask Optimization Yes (Native)
+A review of these tools, summarized in Table 1, justifies the development of a new, dedicated framework.
+At the time MToP development commenced [^1] MToP has been continuously updated since Sep.
+21, 2021 at: https://github.com/intLyc/MTO-Platform, a platform with native, comprehensive support for MTO benchmarking was absent.
+The most prominent existing platform, PlatEMO (Tian2017PlatEMO), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization.
+This presents a fundamental architectural mismatch for EMT.
+Unlike the traditional single-task EC field, solving MTO problems with MTEAs necessitates the simultaneous evolution of multiple optimization tasks, the implementation of inter-task solution space mapping, and diverse knowledge transfer techniques.
+Additionally, the performance metrics in EMT are diverse, encompassing both single-task and multitask metrics, which brings uncertainty to the pattern of results display and analysis.
+These distinctive requirements present significant challenges for implementation in platforms not designed for them.
+Consequently, extending current framework like PlatEMO to natively handle these MTO-specific concepts is non-trivial and leads to significant extensibility limitations.
+This is particularly evident in the context of many-task optimization, where PlatEMO’s support for specialized multitask metrics is restricted.
+Although preliminary MTO support was added to PlatEMO (v4.0) [^2] PlatEMO v4.0 (released on Oct.
+13, 2022) introduced preliminary MTO support., these additions are supplementary rather than architecturally native.
+The other platforms listed in Table 1 are similarly specialized for different, non-MTO-centric goals.
+Moreover, EMT has been extended to the subfields of many-task optimization (Chen2020MaTDE; Liaw2019SBO; Li2024TNG-NES), multi-objective MTO (Gupta2017MO-MFEA; Li2025MTEA-DCK), competitive MTO (Li2022CompetitiveMTO; Li2023MTSRA; Li2024CMO-MTO), and constrained multi-task optimization (Li2022CMTO-Benchmark; Zhang2024CEDA).
+This growing complexity further necessitates a dedicated, MTO-native framework.
+While a new framework was deemed necessary, MToP actively integrates and appropriately cites code modules from these established platforms where feasible, thereby reducing redundant development and focusing on its core MTO-specific contributions.
 3. Architecture of MToP
-MToP consists of key modules including the test, experiment, and data processing modules, which facilitate algorithm execution, visualization, and data analysis.
-These components allow for flexible experimentation through both GUI and command line interfaces.
-Following standardized code patterns, MToP integrates various algorithms and performance metrics.
+In this section, we present the architecture of MToP, encompassing its functional modules, project structure, and code patterns.
 3.1. Functional Modules
+MToP is organized into three interconnected functional modules to support a full research workflow.
+Researchers can first use the Test Module for preliminary analysis and visualization of specific algorithms and problems.
+The Experiment Module facilitates large-scale, multi-run comparative experiments.
+Finally, the Data Process Module provides tools to manage the datasets generated by the Experiment Module, such as merging or splitting results for flexible post-processing.
+The following subsections detail the architectural role and capabilities of these modules.
+A practical guide on their specific operation and usage is provided in Section 4.
 3.1.1. Test Module
@@
 (e) displays the convergence behavior of metrics after executing algorithms on problems.
-The test module is designed to assist researchers in analyzing the characteristics of MTO problems and algorithms.
-It facilitates running the selected algorithm on the chosen problem within MToP.
-The module includes visualization functions for both single-objective and multi-objective MTO problems.
-Users can select different options to display depending on the problem type, resulting in different visual outputs, as demonstrated below.
-In the case of single-objective MTO problems, the landscapes of the objective function with different tasks in the one- and two-dimensional unified search space can be depicted, as shown in Fig.
-1 (a)-(b).
-For tasks with constraints, the feasible regions with different tasks in the two-dimensional unified search space can be visualized, as illustrated in Fig.
-1 (c).
-Regarding multi-objective MTO problems, the Pareto front figures can be generated for multiple tasks, as depicted in Fig.
-1 (d).
-Furthermore, the convergence behavior of metrics can be plotted after executing algorithms on problems, as exemplified in Fig.
-1 (e).
-In addition to static schematic presentations, MToP offers dynamic display utilities such as DrawPopObj() and DrawPopDec() to enable researchers to explore algorithm behavior.
-Figure 2 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, showcasing the evolution in both objective space and decision space.
-CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives.
-The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$.
-The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.4$.
-Through dynamic changes, researchers can intuitively grasp the algorithm’s behavior and delve deeper into its dynamics.
-(a) Gen=1
-(b) Gen=100
-(c) Gen=200
-(d) Gen=500
-(e) Gen=700
-(f) Gen=1000
-(g) Gen=1
-(h) Gen=100
-(i) Gen=200
-(j) Gen=500
-(k) Gen=700
-(l) Gen=1000
-Figure 2.
-Schematic of MO-MFEA’s population on CEC17-MTMO4 in the objective space and decision space with the number of generations.
-The top row (a)-(f) shows the population in the objective space, while the bottom row (g)-(l) displays the population in the decision space.
-In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space.
-As evolution progresses through the first and middle phases (Gen=100 to 200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity.
-Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity.
-By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task.
-The schematics indicate that the population for the second task occupies a similar position in the decision space as the first task.
-However, these two tasks do not precisely overlap in the decision space.
-This observation suggests the presence of negative knowledge transfer, which could be contributing to this discrepancy.
-In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
-Figure 3. The application programming interface calling workflow of running an algorithm MFEA on a problem CEC17-MTP1 with metric Obj by MToP.
-The workflow of API calls for executing an algorithm on a problem is depicted in Fig.
-3.
-The process initiates with the <Start_Button> event in <MTO_GUI.m>, triggering the parameter setting API SetParameter() within the selected algorithm <MFEA.m>.
-Subsequently, it proceeds to the Run() function of the algorithm, which initially invokes the Initialization() API of the algorithm base class <Algorithm.m>, and then iterates through the function evaluation API Evaluation().
-During the initialization phase, the parameter settings of the problem <CEC17_MTP1.m> to be solved are obtained via the problem base class <Problem.m>.
-Similarly, the Evaluation() function retrieves the specific optimization task object through the problem base class.
-Once the algorithm completes its execution, the GUI displays the resulting performance metrics by calling the metrics function <Obj.m> via the CallMetric() API.
-All functions within MToP that execute algorithms, including subsequent experiment modules, adhere to this procedure.
-Additionally, both algorithm and problem parameters can be directly modified through the GUI of MToP.
+The Test Module is designed to assist researchers in the qualitative analysis of MTO problems and algorithms.
+It facilitates executing a single run of a selected algorithm on a chosen problem to generate a suite of diagnostic visualizations.
+These tools help inspect both problem characteristics and algorithmic performance, as illustrated in Fig.
+1.
+For single-objective MTO problems, users can depict function landscapes in one or two dimensions (Fig.
+1 (a)-(b)) or visualize the feasible regions of constrained tasks (Fig.
+1 (c)).
+For multi-objective MTO problems, the module can plot the true Pareto fronts with populations (Fig.
+1 (d)).
+In addition to these static plots, the module provides dynamic visualization utilities.
+Once a run is initiated, researchers can enable options such as Draw Dec (decision space) and Draw Obj (objective space) to observe the population’s evolution in real-time.
+Upon completion of the run, the final performance metrics are calculated, and their convergence behavior (as shown in Fig.
+1 (e)) is displayed in a dedicated results panel on the right side of the interface.
+The Test Module serves as an essential preliminary step for researchers to understand the nuances of specific MTO problems and the behavior of algorithms before embarking on large-scale experiments.
 3.1.2. Experiment Module
-Table 1. Results of Obj metric and standard deviations with Wilcoxon rank-sum test of 30 independent runs.
-GA MFEA MTEA-AD
-CI-HS-T1 9.0426e-01 (6.16e-02) 8.4445e-01 (5.51e-02) + 9.3438e-01 (5.03e-02) =
-CI-HS-T2 4.2400e+02 (4.90e+01) 2.8119e+02 (5.09e+01) + 3.1099e+02 (4.11e+01) +
-CI-MS-T1 5.4562e+00 (1.52e+00) 5.8307e+00 (8.23e-01) - 4.5085e+00 (2.69e-01) +
-CI-MS-T2 4.2474e+02 (6.45e+01) 3.1448e+02 (4.41e+01) + 3.1660e+02 (4.39e+01) +
-CI-LS-T1 2.1167e+01 (1.82e-01) 2.0234e+01 (8.35e-02) + 2.1086e+01 (2.82e-01) =
-CI-LS-T2 4.1476e+03 (6.34e+02) 4.3638e+03 (4.88e+02) = 4.5958e+03 (6.64e+02) -
-PI-HS-T1 4.2186e+02 (5.27e+01) 6.6221e+02 (1.11e+02) - 4.4662e+02 (4.13e+01) -
-PI-HS-T2 8.6994e+01 (1.93e+01) 8.9970e+01 (1.67e+01) = 1.2177e+02 (2.88e+01) -
-PI-MS-T1 5.4736e+00 (1.42e+00) 4.9276e+00 (4.76e-01) = 4.7264e+00 (3.50e-01) +
-PI-MS-T2 2.4152e+04 (1.00e+04) 1.0488e+04 (4.77e+03) + 9.1688e+03 (4.69e+03) +
-PI-LS-T1 5.6540e+00 (1.88e+00) 1.9583e+01 (2.71e+00) - 5.3895e+00 (9.11e-01) =
-PI-LS-T2 1.4062e+01 (2.48e+00) 1.9371e+01 (3.26e+00) - 5.1076e+00 (1.44e+00) +
-NI-HS-T1 2.9452e+04 (1.49e+04) 9.7165e+03 (4.93e+03) + 1.3459e+04 (4.77e+03) +
-NI-HS-T2 4.5183e+02 (6.26e+01) 3.7757e+02 (7.84e+01) + 3.8559e+02 (6.26e+01) +
-NI-MS-T1 9.0462e-01 (6.19e-02) 9.3064e-01 (5.16e-02) = 9.8876e-01 (4.48e-02) -
-NI-MS-T2 3.8662e+01 (3.97e+00) 2.7501e+01 (3.26e+00) + 2.4831e+01 (3.76e+00) +
-NI-LS-T1 4.2895e+02 (5.41e+01) 7.4411e+02 (1.30e+02) - 4.6340e+02 (4.97e+01) -
-NI-LS-T2 4.3308e+03 (5.19e+02) 4.5452e+03 (4.90e+02) = 4.7449e+03 (6.45e+02) -
-+ / - / = Base 8 / 5 / 5 9 / 6 / 3
 (a) Convergence
@@
 Figure 2. schematic plotting by the Experiment Module of MToP. (a) Convergence plot of algorithms on problems. (b) Pareto front plot of multi-objective problems.
-The experiment module offers functionalities for conducting comprehensive experiments and analyzing results.
-Users can execute multiple independent runs of various algorithms to tackle multiple MTO problems.
-Upon completion of the runs, data including experimental settings, objective values, and decision variables are recorded and stored.
-These data serve as the basis for generating metric results, conducting significance tests, and creating schematic plots.
-Metrics functions operating on the data are categorized into single-task and multitask metrics.
-The selection of single-task metrics presents the results of each task of the MTO problem in tabular format within MToP while choosing to multitask metrics displays the results of each MTO problem.
-Significance test results are also included in the table alongside the metric results.
-All table data can be exported in formats such as tex, csv, and xlsx, as demonstrated in Table 1.
-Convergence plots are generated based on the selected table data to facilitate the comparison of convergence behaviors among different algorithms, as depicted in Fig.
-4 (a).
-The convergence plot function can be applied to both single- and multi-objective problems.
-Furthermore, for multi-objective MTO algorithms and problems, the Pareto front can be plotted, as illustrated in Fig.
-4 (b).
+The Experiment Module provides the core functionalities for conducting comprehensive, large-scale experiments and analyzing the resulting data.
+It is designed to manage and execute batch runs, allowing users to test multiple algorithms across multiple MTO problems for a specified number of independent repetitions.
+Upon completion, the module systematically records all experimental settings, objective values, and decision variables.
+A key architectural feature is its support for MATLAB’s Parallel Computing Toolbox, enabling the parallel execution of independent runs to significantly reduce total experimental time.
+The module supports saving the complete, raw experimental dataset (MTOData) to a .mat file, which can be fully reloaded into the platform at any time.
+Once the data is generated, the module offers a suite of tools for post-processing and analysis.
+This includes functions to automatically calculate performance metrics, which are categorized to support both single-task and multi-task indicators, and to perform statistical significance tests to compare algorithmic performance.
+All generated tables and test results can be exported to .xlsx, .csv, or .tex formats.
+Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang2022IOHanalyzer) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
+Finally, the module provides essential visualization capabilities for interpreting the experimental outcomes.
+As shown in Fig.
+2, this includes generating convergence plots to simultaneously compare the performance trajectories of multiple algorithms across various problems (Fig.
+2 (a)), a function applicable to both single- and multi-objective problems.
+Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto fronts (Fig.
+2 (b)) to visually assess the quality of the final solution sets.
+Since all visualizations are generated as standard MATLAB figures, users can leverage MATLAB’s built-in plotting tools for further customization, annotation, and export.
 3.1.3. Data Process Module
+Each unique execution within the Experiment Module is saved as a standardized data object to a MTOData.mat file, which is described in detail in Section 3.2.4, this file stores all experimental settings, results, and metadata.
+This saved data can be fully reloaded by the Experiment Module for post-processing or managed by the Data Process Module.
+The core function of the Data Process Module is to manage and manipulate these data objects, enabling data reuse and customization.
+It provides flexible Merge and Split capabilities.
+These operations can be performed along three distinct dimensions of the dataset: by independent runs (e.g., combining two 10-run sets into one 20-run set), by algorithms (e.g., adding a new algorithm’s results to an existing dataset), or by problems (e.g., splitting a large benchmark suite into smaller subsets).
+This functionality is crucial for maintaining organized datasets and efficiently conducting comparative studies, such as when a new algorithm needs to be benchmarked against a set of previously executed experiments.
+3.2. Project Structure
+The project structure of MToP is organized into three main perspectives: class diagram, sequence workflow, file structure, and experiment data structure.
+These perspectives collectively illustrate the static architecture, dynamic interactions, and organizational layout of MToP.
+3.2.1. Class Diagram
+Figure 3. Class diagram of MToP. The main classes include Algorithm, Problem, Individual.
+Fig. 3 illustrates the class diagram of MToP, which outlines the static architecture of the platform. The design is centered around three core base classes and functions:
+  • Individualclass is a data structure encapsulating a single solution, storing its decision variables (Dec), objective values (Obj), and constraint information (Con, CV).
+  • Algorithm class is the abstract base for all optimization algorithms with parameters ($FE$ for current function evaluations, $Gen$ for generations), managing the population of Individual objects (via an aggregation “has a” relationship) and the main evolutionary loop.
+A concrete example, MFEA, is shown inheriting from it.
+The Algorithm class also holds a “use a” dependency on the Problem class, as it needs to evaluate individuals.
+  • Problem class is an abstract base for optimization tasks, defining the essential evaluate() method and problem parameters ($T$ for task, $N$ for population size, $M$ for objective, $D$ for dimension).
+  • Function module groups static utility functions such as metrics (e.g., IGD()) and evolutionary operators (e.g., GA_Crossover()).
+3.2.2. Sequence Workflow
+Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm, Problem, and Metric.
+Fig. 4 depicts the dynamic interaction workflow between the main components when running an experiment:
+  • The process is initiated by a User through either the MTO_GUI or MTO_CMD interface.
+  • Then, the interface handles the setup phase, configuring the Algorithm and Problem instances by calling setParameter() and setTasks().
+Next, the interface invokes the run() method of the Algorithm.
+The Algorithm performs its Initialization() and then enters the main optimization loop, conditioned on while notTerminated().
+  • Inside the loop, the Algorithm executes its Evaluation() method, which subsequently calls the evaluate(individual) method on the Problem object.
+The Problem returns the calculated Obj, Con values.
+Following evaluation, the Algorithm calls EvolvePopulation() to generate the next population.
+  • Once the loop terminates, the Algorithm returns the bestResult to the GUI/CMD interface.
+The interface then passes this result to the Metric module for Calculation().
+Finally, the computed metricValue is returned and displayed to the user.
+3.2.3. File Structure
+Figure 5. File structure of MToP. The root directory contains the main script file mto.m and four subfolders: Algorithms/, Problems/, Metrics/, and GUI/.
+The file structure of MToP is shown in Fig. 5. The root directory contains the main script file mto.m and four primary subfolders:
+  • Algorithms/: it contains various categories of algorithms, utility functions, and the algorithm base class Algorithm.m.
+Specific algorithm files such as MFEA.m, MO_MFEA.m, and GA.m are stored within their respective classification folders.
+The Algorithms/Utils/ folder contains multiple subfolders, including Operator/ and Selection/, which organize specific functional modules.
+The Operator/ subfolder stores variation operators such as crossover and mutation, while the Selection/ subfolder includes mating and environmental selection approaches.
+Other subfolders provide additional utility functions such as constraint handling and multi-objective optimization techniques.
+  • Problems/: it contains various categories of problems, the base function folder, and the problem base class Problem.m.
+Specific problem files like WCCI20_MTSO1.m, CMT3.m, and CEC19_MaTMO6.m are stored under their corresponding classification folders.
+Real-world application problem files are located in the Real-World Application/ folder, while the Problems/Base/ folder contains public base functions for problems.
+  • Metrics/: it contains all result evaluation metrics, organized into subfolders for single-objective Metrics/Single-objective/ and multi-objective Metrics/Multi-objective optimization.
+Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
+  • GUI/: it contains all files used by the GUI of MToP.
+Among them, MTO_GUI.m serves as the main file of the GUI, while MTO_CMD.m provides functionality for executing experiments via the command line.
+3.2.4. Experiment Data Structure
+The experiment data structure is the standardized format used by MToP to save the complete output of an experimental batch, regardless of whether it is generated by the MTO_CMD.m command-line runner or the MTO_GUI.m graphical interface.
+All experimental data is encapsulated within a single MATLAB struct variable named MTOData, which is then saved to a .mat file (e.g., MTOData.mat).
+This self-contained design ensures portability and simplifies downstream data management and post-processing.
 Table 2. Properties of experimental data.
@@
 Metrics Calculated metric results data
-Each unique execution within the experiment module is considered a data object and is saved to a data file named <MTOData.mat>.
-The settings and intermediate results of the experiment are stored in data properties, as illustrated in Table 2.
-This saved data can subsequently be reloaded by the experiment module of MToP.
-With the data process module, the data can be merged or split, enabling data reuse and customization.
-The processing of data is categorized into three types: independent runs, algorithms, and problems.
-If several sets of data with fewer independent runs are available, the data can be merged to create new data with a larger number of independent runs.
-Similarly, in cases where existing data needs augmentation with new algorithms for comparison, it is unnecessary to re-execute all comparative experiments.
-Instead, only the new algorithms need to be executed, and the resulting data can be merged with the original data.
-Likewise, if not all algorithms within the data are required, the data can be split and re-merged accordingly.
-The data processing for problems follows a similar approach, wherein new problems can be added or deleted as needed.
-In practical use, the specific operations for splitting and merging data are as follows:
-  • Data Splitting: If an experiment involves running both the MFEA and EMEA algorithms on same problems, the data can be split by selecting it in the <Data Tree> and clicking the <Algorithms Split> button to generate separate datasets for each algorithm.
-Similarly, data can also be split by independent runs or problems using the respective options.
-  • Data Merging: If the MFEA and EMEA algorithms were run on same problems in separate experiments, the data can be merged by selecting both datasets and clicking the <Algorithms Merge> button, provided the settings are identical except for the algorithm.
-This process can also be used to merge datasets by independent runs or problems.
-2.2. File Structure
-Figure 5. File structure of MToP. The root directory contains the main script file <mto.m> and four subfolders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
-In the root directory <MTO/> of MToP, you’ll find a script file named <mto.m> and four folders: <Algorithms/>, <Problems/>, <Metrics/>, and <GUI/>.
-The structure of these folders is depicted in Fig.
-5.
-The <Algorithms/> folder contains various categories of algorithms, utility functions, and the algorithm base class <Algorithm.m>.
-Specific algorithm files such as <MFEA.m>, <MO_MFEA.m>, and <GA.m> are stored within their respective classification folders.
-The <Algorithms/Utils/> folder provides public utility functions for generation operators, environmental selection approaches, constraint handling techniques, etc.
-The <Problems/> folder houses various categories of problems, the base function folder, and the problem base class <Problem.m>.
-Specific problem files like <WCCI20_MTSO1.m>, <CMT3.m>, and <CEC19_MaTMO6.m> are stored under their corresponding classification folders.
-Real-world application problem files are located in the <Real-World Application/> folder, while the <Problems/Base/> folder contains public base functions for problems.
-The <Metrics/> folder contains all result evaluation metrics, organized into subfolders for single-objective <Metrics/Single-objective/> and multi-objective <Metrics/Multi-objective> optimization.
-Specific metric files include <Obj.m>, <Obj_MTS.m>, and <IGD.m>, among others.
-Lastly, the <GUI/> folder houses all files used by the GUI of MToP.
-Among them, <MTO_GUI.m> serves as the main file of the GUI, while <MTO_CMD.m> provides functionality for executing experiments via the command line.
+The high-level properties stored within the MTOData struct are summarized in Table 2.
+The core fields are organized as multi-dimensional arrays to map clearly onto the experimental design.
+Let $P$ be the number of problems, $A$ be the number of algorithms, and $R$ be the number of Reps.
+  • Problems: This is a $1\times P$ struct array.
+Each element MTOData.Problems(i) stores the static metadata for the $i$-th problem, including its Name, number of tasks (T), objective dimensions (M), decision variable dimensions (D), and maximum function evaluations (maxFE).
+  • Algorithms: This is a $1\times A$ struct array. Each element MTOData.Algorithms(j) stores the metadata for the $j$-th algorithm, such as its Name and a struct of its specific parameters (Para).
+  • RunTimes: This is a $P\times A\times R$ numerical matrix, where the element (i, j, k) stores the wall-clock execution time for running algorithm $j$ on problem $i$ during repetition $k$.
+  • Results: This is the $P\times A\times R$ struct array containing the core optimization output.
+Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved convergence points (where $G$ is the Results Num).
+Each struct contains:
+    • Obj: Stores the objective values.
+For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each convergence point.
+For multi-objective problems, this is a $T\times 1$ cell array, where each cell Obj{t} contains a $G\times N\times M_{t}$ array representing the full objective values of $N$ solutions.
+    • CV: Stores the constraint violation values.
+This is typically a $T\times G$ array (for single-objective problems) or a $T\times G\times N$ array (for multi-objective problems) storing the summed constraint violation for the corresponding solutions in Obj.
+    • Dec: Stores the decision variables, which are populated only if the Save_Dec flag is set to true.
+For single-objective problems, this is a $T\times G\times D_{t}$ array storing the decision vector of the best solution.
+For multi-objective problems, it is a $T\times G\times N\times D_{t}$ array storing the decision vectors for all $N$ solutions.
+  • Metrics: This field is not generated by the experiment runners (CMD or GUI) themselves but is designed to be populated later by the Experiment Module when metrics are calculated and saved back to the data file.
+This consistent, multi-dimensional struct-based layout is fundamental to the platform, as it allows the Data Process Module to reliably merge and split datasets by manipulating the Problems, Algorithms, or Reps dimensions of this structure.
 3.3. Code Pattern
+This subsection details the code patterns employed in MToP for implementing algorithms, problems, and metrics.
+By adhering to these standardized patterns, developers can easily extend the platform with new methods while ensuring compatibility with the existing architecture and GUI.
 3.3.1. Algorithm
@@
 notTerminated() Determine whether to terminate and update result data
-Evaluation() Fitness function evaluation
+Evaluation() Function evaluation with algorithm state update
 run() Executing algorithm
 1 label={list:algo},
-2 caption={Algorithm implementation example for multi-population EA.},
+2 caption={Algorithm implementation example for multi-population MTEA.},
 3 captionpos=b,
@@
 14function run(Algo, Prob)
-15% initialize multiple populations for all tasks
-16pops = Initialization(Algo, Prob, Individual);
-17
-18while Algo.notTerminated(Prob, pops) % main loop
+15 % initialize multiple populations for all tasks
+16 pops = Initialization(Algo, Prob, Individual);
+17 % main loop
+18 while Algo.notTerminated(Prob, pops)
 19 for t = 1:Prob.T % for each task
 20 % generate new offspring
-21 off(index1) = Generation(pops{t});
+21 offspring(index1) = Generation(pops{t});
 22 % knowledge transfer
-23 off(index2) = KnowledgeTransfer(pops)
+23 offspring(index2) = KnowledgeTransfer(pops);
 24 % evaluate the fitness of offspring
-25 off = Algo.Evaluation(off, Prob, t);
+25 offspring = Algo.Evaluation(offspring, Prob, t);
 26 % environmental selection
-27 pops{t} = Selection(pops{t}, off);
+27 pops{t} = Selection(pops{t}, offspring);
 28 end
-29end
+29 end
 30end
 1 label={list:algo2},
-2 caption={Algorithm implementation example for multifactorial EA.},
+2 caption={Algorithm implementation example for multifactorial MTEA.},
 3 captionpos=b,
@@
 14function run(Algo, Prob)
-15% initialize populations for multifactorial evolution
-16population=Initialization_MF(Algo, Prob, Individual);
-17
-18while Algo.notTerminated(Prob, pops) % main loop
+15 % initialize populations for multifactorial evolution
+16 population = Initialization_MF(Algo, Prob, Individual);
+17 % main loop
+18 while Algo.notTerminated(Prob, population)
 19 % generate new offspring with random mating
-20 off = Generation(population);
+20 offspring = Generation(population);
 21 % evaluate based on offspring skill factor
 22 for t = 1:Prob.T
-23 idx = [off.MFFactor] == t;
-24 off(idx) = Algo.Evaluation(off(idx), Prob, t);
+23 idx = [offspring.MFFactor] == t;
+24 offspring(idx) = Algo.Evaluation(offspring(idx), Prob, t);
 25 end
 26 % environmental selection
-27 population = Selection(population, off);
-28end
+27 population = Selection(population, offspring);
+28 end
 29end
-All algorithms within MToP inherit from the algorithm base class <Algorithm.m>.
+All algorithms within MToP inherit from the algorithm base class Algorithm.m.
 The properties and methods of this base class are detailed in Table 3.
 This base class encapsulates all functions that interface with the GUI, simplifying the implementation process for algorithms, which only need to focus on the evolutionary workflow itself.
-The MTEAs implemented in MToP are outlined in Table 4.
+The MTEAs implemented in MToP are outlined in Table LABEL:tab:algo-mt of the supplementary files.
 Additionally, the single-task EAs are exclusively implemented using multi-population methods tailored for MTO problems.
-Specific algorithms belonging to this category are listed in Table 5.
+Specific algorithms belonging to this category are listed in Table LABEL:tab:algo-st.
 Listings LABEL:list:algo and LABEL:list:algo2 exemplify the implementation of multi-population and multifactorial algorithms respectively.
 The algorithm labels provided in the second line serve as identifiers for classification within the GUI.
-The run() function within each class orchestrates algorithm execution, with <Algo> representing the object itself and <Prob> denoting the problem object to be solved.
+The run() function within each class orchestrates algorithm execution, with Algo representing the object itself and Prob denoting the problem object to be solved.
 Population initialization occurs at the outset of the run() function, utilizing either the multi-population or multifactorial method provided by MToP.
-In Listing LABEL:list:algo, the variable <pops> is of <cell> data type, housing multiple populations as <population1, population2, population3...>.
-Each individual within <pops> is an <Individual> object, containing properties such as objective value <Obj>, constraint violation <CV>, and decision variables <Dec>.
-Contrastingly, Listing LABEL:list:algo2 utilizes a list data type for the <population> variable, directly containing individuals as <individual1:factor1, individual2:factor1, individual3:factor2...>, along with an additional property <MFFactor>.
+Here, most standard algorithms read the N property from the Prob object (see Table 4) to use the problem’s default population size, which ensures consistency for comparisons.
+However, this is not a rigid requirement: algorithms known for dynamic population sizing (e.g., L-SHADE, IPOP-CMA-ES) are implemented to override this default and manage their own population sizes as per their original designs.
+In Listing LABEL:list:algo, the variable pops is a cell array.
+It stores multiple populations grouped by task, in the form of {pop1, pop2, pop3, ...}.
+Each element pop is itself a separate object array containing all individuals [ind1, ind2, ind3, ...] for a specific task.
+Each Individual object within these lists contains properties such as Obj (objective value), Con (constraint value), and Dec (decision variables).
+In contrast, Listing LABEL:list:algo2 uses a object array data type for the whole population variable.
+This adopts a mixed structure, storing individuals from all tasks together in a single list.
+To differentiate which task each individual belongs to, the Individual objects in this listing include an additional property, MFFactor.
+Therefore, the structure of this single list can be conceptualized as [ind1, ind2, ind3, ...], where each individual’s MFFactor property is set to its corresponding task ID (e.g., ind1.MFFactor = 1, ind2.MFFactor = 2, ind3.MFFactor = 1, ...).
 Subsequently, the primary loop commences with the invocation of the notTerminated() function, a component of the algorithm base class.
@@
 Conversely, the multifactorial approach achieves knowledge transfer through random mating within the Generation() function.
-The Generation() function operates on the decision variables <Dec> of offspring individuals and is tailored to specific algorithms.
-Meanwhile, the Evaluation() function, embedded within the algorithm base class, automatically engages the problem object <Prob> for offspring fitness evaluation.
-This entails calculating the objective value <Obj> and constraint violation <CV> for all individuals.
-Following each loop iteration, the environmental selection function Selection() is invoked to update the new population.
+The Generation() function operates on the decision variables Dec of offspring individuals and is tailored to specific algorithms.
+The Evaluation() method serves as a state-updating wrapper in Algo.
+It performs two sequential operations: first, it invokes the evaluate() method from the Prob object to calculate the actual Obj and Con values.
+Second, it updates the algorithm’s internal state properties, such as incrementing the FE counter.
+Following, the environmental selection function Selection() is invoked to update the new population.
 While MToP offers universal Selection() functions, specific algorithms also have the option to reimplement this function.
@@
 The algorithmic structure in MToP is designed to accommodate all EAs for solving MTO problems.
-Table 4.
-Multitask evolutionary algorithms included in MToP.
-The table lists the algorithm name, optimization objective, task type, framework, constraint handling, and a brief description.
-SO and MO denote single-objective and multi-objective optimization, respectively.
-MT and MaT represent multi-task and many-task, respectively.
-MF and MP represent multifactorial and multi-population, respectively.
-Algorithm Objective Task Framework Constraint Description
-MFEA (Gupta et al., 2016) SO MT MF Multifactorial evolution-based GA for MTO
-LDA-MFEA (Bali et al., 2017) SO MT MF Linearized domain adaptation in MFEA
-MFDE&MFPSO (Feng et al., 2017) SO MT MF Multifactorial evolution-based DE and PSO
-G-MFEA (Ding et al., 2019) SO MT MF Generalized MFEA with decision variable translation and shuffling
-MFEA-GHS (Liang et al., 2019) SO MT MF Genetic Transform and hyper-rectangle search in MFEA
-MFEA-DV (Yin et al., 2019) SO MT MF Enhanced MFEA with cross-task search direction
-MFEA-II (Bali et al., 2020) SO MT MF MFEA with online transfer parameter estimation
-SREMTO (Zheng et al., 2020) SO MT MF Self-regulated multitask framework for varying relatedness among tasks
-TLTLA (Ma et al., 2020) SO MT MF MFEA with two-level inter- and intra-task transfer learning
-MFEA-AKT (Zhou et al., 2021b) SO MT MF MFEA with adaptive knowledge transfer via multiple crossover operator
-ASCMFDE (Tang et al., 2021a) SO MT MF Intertask transfer in aligned subspace for MFDE
-AT-MFEA (Xue et al., 2022) SO MT MF Affine transformation-enhanced domain adaptation for MFEA
-MFEA-DGD (Liu et al., 2023) SO MT MF MFEA based on diffusion gradient descent
-MFEA-VC (Wang et al., 2024) SO MT MF MFEA with contrastive variational auto-encoder
-IMEA (Hashimoto et al., 2018) SO MT MP Multi-population framework for MTO through island model
-EMEA (Feng et al., 2019b) SO MT MP Explicit knowledge transfer for MTO via autoencoding
-MFMP (Li et al., 2020) SO MT MP Multi-population-based adaptive DE for MTO
-DEORA (Li et al., 2022c) SO MT MP Adaptive task selection for competitive MTO
-MKTDE (Li et al., 2022b) SO MT MP Meta-knowledge transfer-based multitask DE
-MTES (Bai et al., 2022) SO MT MP Multitask OpenAI-ES via gradient-free evolution multitasking
-BLKT-DE (Jiang et al., 2023b) SO MT MP Multitask DE with block-level knowledge transfer
-MTSRA (Li et al., 2023a) SO MT MP Improved adaptive DE with competitive task selection
-MTDE-ADKT (Zhang et al., 2024a) SO MT MP Multitask DE with adaptive dual knowledge transfer
-CEDA (Zhang et al., 2024b) SO MT MF/MP Constrained multitasking via co-evolution and domain adaptation
-CMO-LKT (Ban et al., 2025) SO MT MP Local knowledge transfer-based EA for constrained MTO
-MTEA-AD (Wang et al., 2022a) SO MT/MaT MP MTEA with adaptive knowledge transfer via anomaly detection
-MTES-KG (Li et al., 2024c) SO MT/MaT MP Multitask ES with knowledge-guided external sampling
-MTEA-HKTS (Zhao et al., 2024) SO MT/MaT MP MTEA for solving the problem of transfer targets
-BoKT (Jiang et al., 2023a) SO MaT MP Bi-objective knowledge transfer framework for many-task optimization
-TRADE (Wu et al., 2023) SO MaT MP Transferable adaptive parameter DE for many-task optimization
-TNG-NES (Li et al., 2024b) SO MaT MP Transfer task-averaged natural gradient for efficient many-task optimization
-MTEA-SaO (Li et al., 2023b) SO/MO MT/MaT MP Adaptive solver multitask framework with implicit knowledge transfer
-KR-MTEA (Cui et al., 2023) SO/MO MT/MaT MP Adaptive multi-task EA based on knowledge reuse
-SBO (Liaw and Ting, 2019) SO/MO MaT MP Symbiosis in biocoenosis framework for many-task optimization
-MaTDE (Chen et al., 2020) SO/MO MaT MP Many-task DE with adaptive archive-based knowledge transfer
-EMaTO-MKT (Liang et al., 2022c) SO/MO MaT MP Multi-source knowledge transfer via local distribution estimation
-MO-MFEA (Gupta et al., 2017) MO MT MF Multiobjective MFEA with non-dominated sort and crowding distance
-MO-MFEA-II (Bali et al., 2021) MO MT MF Cognizant Multitasking for parameter estimation in MO-MFEA
-EMT-PD (Liang et al., 2022b) MO MT MF Two-stage adaptive knowledge transfer based on population distribution
-EMT-GS (Liang et al., 2023) MO MT MF Generative adversarial networks for knowledge transfer
-MM-DE (Chen et al., 2018) MO MT MP Fast memetic DE for multiobjective MTO
-AMT-NSGA-II (Da et al., 2019) MO MT MP Curbing negative influences online for seamless transfer
-EMT-ET (Lin et al., 2021) MO MT MP Effective knowledge transfer approach via non-dominated sort
-MOMFEA-SADE (Liang et al., 2022a) MO MT MP Subspace alignment and adaptive Differential Evolution
-MTEA-D-DN (Wang et al., 2023) MO MT MP Neighborhood as a bridge for decomposition-based knowledge transfer
-MTEA-D-TSD  (Li et al., 2024a) MO MT MP Transfer search directions among decomposed subtasks
-MTDE-MKTA (Li and Gong, 2025) MO MT MP Multitask DE with multiple knowledge types and transfer adaptation
-RVC-MTEA (Li et al., 2024g) MO MT MP Reference vector contribution based MTEA for competitive multiobjective MTO
-MTEA-DCK (Li et al., 2025) MO MT MP Multiobjective MTEA with diversity- and convergen-based knowledge transfer
-Table 5.
-Single-task evolutionary algorithms for multitask optimization included in MToP.
-The table lists the algorithm name, optimization objective, task type, category, constraint handling, and a brief description.
-SO and MO denote single-objective and multi-objective optimization, respectively.
-GA denotes genetic algorithm, DE denotes differential evolution, SI denotes swarm intelligence, ES denotes evolution strategy, CH denotes constraint handling, ND denotes non-dominated sorting, ID denotes indicator-based, DC denotes decomposition, and RV denotes reference vector.
-Algorithms Objective Task Category Constraint Description
-GA (Zbigniew, 1996) SO ST/MT GA Genetic algorithm with SBX crossover and polynomial mutation
-Global-GA (Bull and Liu, 2024) SO ST/MT GA GA with global crossover to replace cooperative coevolution
-DE (Storn, 1996) SO ST/MT DE Differential evolution algorithm with DE/rand/1/bin operator
-jDE (Brest et al., 2006) SO ST/MT DE Self-adapting control parameters in DE
-JADE (Zhang and Sanderson, 2009) SO ST/MT DE Adaptive DE with fast convergence performance
-rank-DE (Gong and Cai, 2013) SO ST/MT DE Ranking-based mutation operators for DE
-SHADE (Tanabe and Fukunaga, 2013) SO ST/MT DE Success-history based parameter adaptation for DE
-L-SHADE (Tanabe and Fukunaga, 2014) SO ST/MT DE Linear population size reduction for SHADE
-LSHADE44 (Poláková, 2017) SO ST/MT DE L-SHADE with competing strategies applied to constrained optimization
-CAL-SHADE (Zamuda, 2017) SO ST/MT DE Adaptive constraint handling technique for L-SHADE
-jSO (Brest et al., 2017) SO ST/MT DE Champion of CEC 2017 numerical optimization competition via improved DE
-NL-SHADE-RSP (Stanovov et al., 2021) SO ST/MT DE Champion of CEC 2021 numerical optimization competition via improved DE
-PSO (Kennedy and Eberhart, 1995) SO ST/MT SI Particle swarm optimization with global and particle best update
-CSO (Cheng and Jin, 2015) SO ST/MT SI Competitive swarm optimizer for large-scale optimization
-MPA (Faramarzi et al., 2020a) SO ST/MT SI Marine predators algorithm for numerical optimization
-EO (Faramarzi et al., 2020b) SO ST/MT SI Equilibrium optimizer for numerical optimization
-AO (Abualigah et al., 2021) SO ST/MT SI Aquila optimizer for numerical optimization
-KLDE&KLPSO (Jiang et al., 2023c) SO ST/MT DE/SI Knowledge learning-based DE and PSO for numerical optimization
-CMA-ES (Hansen and Ostermeier, 2001) SO ST/MT ES Evolution strategy with derandomized covariance matrix adaptation
-IPOP-CMA-ES (Auger and Hansen, 2005) SO ST/MT ES Restart CMA-ES with increasing population size
-sep-CMA-ES (Ros and Hansen, 2008) SO ST/MT ES Separable CMA-ES for large-scale optimization
-xNES-as (Schaul, 2012) SO ST/MT ES Natural ES with adaptation sampling
-R1-NES (Sun et al., 2013) SO ST/MT ES A Linear Time natural ES for non-separable functions
-xNES&SNES (Wierstra et al., 2014) SO ST/MT ES ES with adaptive natural gradients to update distribution
-OpenAI-ES (Salimans et al., 2017) SO ST/MT ES Parallelized ES with standard normal distribution gradients
-DES (Arabas and Jagodziński, 2020) SO ST/MT ES Matrix-free covariance matrix adaptation ES
-MTV-DE (Mezura-Montes et al., 2007) SO ST/MT CH Multiple trial vectors in DE for handling constraints
-ECHT (Mallipeddi and Suganthan, 2010b) SO ST/MT CH Ensemble of constraint handling techniques
-FROFI (Wang et al., 2016) SO ST/MT CH Incorporating objective function information into the feasibility rule
-C2oDE (Wang et al., 2019) SO ST/MT CH Composite DE for constrained optimization
-CORCO (Wang et al., 2020) SO ST/MT CH Utilizing the correlation between constraints and objective function
-DeCODE (Wang et al., 2021) SO ST/MT CH Decomposition-based multiobjective approach for constrained optimization
-VMCH (Wu et al., 2022) SO ST/MT CH Voting-mechanism-based ensemble of constraint handling techniques
-CEDE-DRL (Hu et al., 2023) SO ST/MT CH Deep reinforcement learning assisted co-evolutionary DE
-CCEF-ECHT (Li et al., 2024d) SO ST/MT CH Competitive and cooperative ensemble of constraint handling techniques
-SPEA2 (Zitzler et al., 2001) MO ST/MT ND Improving the strength Pareto approach for multiobjective optimization
-NSGA-II (Deb et al., 2002) MO ST/MT ND Multiobjective GA with non-dominated sort and crowding distance
-NSGA-III (Deb and Jain, 2014) MO ST/MT ND Reference-point-based nondominated sorting for multiobjective optimization
-SMS-EMOA (Beume et al., 2007) MO ST/MT ID Hypervolume-based selection for multiobjective optimization
-MO-CMA-ES (Igel et al., 2007) MO ST/MT ES Covariance matrix adaptation for multiobjective optimization
-MOEA/D (Zhang and Li, 2007) MO ST/MT DC Decomposition-based multiobjective evolutionary algorithm
-MOEA/D-DE (Li and Zhang, 2009) MO ST/MT DC MOEA/D for complex Pareto sets in multiobjective optimization
-RVEA (Cheng et al., 2016) MO ST/MT RV Reference vector-guided EA for many-objective optimization
-AR-MOEA (Tian et al., 2018) MO ST/MT ID Adaptive reference points for MOEA/D in multiobjective optimization
-MSEA (Tian et al., 2021a) MO ST/MT ND Multistage EA for better diversity preservation in multiobjective optimization
-CCMO (Tian et al., 2021b) MO ST/MT CH Coevolutionary framework for constrained multiobjective optimization
-LMOCSO (Tian et al., 2020) MO ST/MT SI Efficient large-scale multiobjective optimization based on CSO
-CMOCSO (Ming et al., 2023) MO ST/MT SI Competitive and cooperative swarm optimizer for multiobjective optimization
 3.3.2. Problem
-All problems within MToP inherit from the problem base class <Problem.m>.
-The properties and methods of <Problem.m> are detailed in Table 7.
-As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach (Gupta et al., 2016).
+All problems within MToP inherit from the problem base class Problem.m.
+The properties and methods of Problem.m are detailed in Table 4.
+As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach as Eq.
+(2).
 It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
 A simple problem implementation example is illustrated in Listing LABEL:list:prob.
-The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor.
+The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor.
 The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task.
-Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task.
-Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task.
-For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type.
-Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>.
-Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
+Within this function, the number of tasks T is first defined, followed by a detailed setup for each task.
+Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task.
+For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type.
+Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint value Con.
+The problem’s base class evaluate() method manages this process by using the specific function handle defined in Fnc to perform the calculation for the requested task.
+Moreover, for multi-objective optimization, the external function getOptimum() is provided.
+It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD.
+For problems where the true front is unknown (e.g., real-world applications), this function is instead used to provide the reference point for calculating metrics such as hypervolume (HV).
 With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP.
-The existing problems available in MToP are enumerated in Table 6.
-In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
-Table 6.
-Benchmark and real-world optimization problems included in MToP.
-The table lists the problem name, case number, task type, dimension, optimization objective, task type, special case, and a brief description.
-SO and MO denote single-objective and multi-objective optimization, respectively.
-ST, MT and MaT represent single-task, multi-task and many-task, respectively.
-Problem Case Task Dimension Objective Task Special Description
-CEC17-MTSO (Da et al., 2017) 9 2 50 SO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MaTSO (Feng et al., 2019a) 6 2-50 50 SO MaT Single-objective many-task optimization problems
-WCCI20-MTSO (Feng et al., 2020) 10 2 50 SO MT Complex single-objective MTO problems
-WCCI20-MaTSO (Feng et al., 2020) 10 2-50 50 SO MaT Complex single-objective many-task optimization problems
-CMT (Li et al., 2022a) 9 2 Any SO MT Constrained Large/partial/no feasible regions interaction and different inter-task similarity
-C2TOP&C4TOP (Li et al., 2022c) 51 2/4 50 SO MT Competitive Different optimal objective values among competitive tasks
-C-CPLX (Li et al., 2023a) 20 2 50 SO MT Competitive Complex competitive MTO problems
-LSMaTSO (Li et al., 2024b) 5 2-50 300 SO MaT Large-scale many-task optimization problems
-CEC17-MTMO (Yuan et al., 2017) 9 2 50 MO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
-CEC19-MTMO (Feng et al., 2019a) 10 2 50 MO MT Complex multi-objective MTO problems
-CEC19-MaTMO (Feng et al., 2019a) 10 2-50 50 MO MaT Multi-objective many-task optimization problems
-WCCI20-MaTMO (Feng et al., 2020) 10 2-50 50 MO MaT Complex multi-objective many-task optimization problems
-CEC21-MTMO (Feng et al., 2021b) 10 2 50 MO MT More complex multi-objective multi optimization problems
-CMOMT (Li et al., 2024g) 12 2/3 30-60 MO MT Competitive multiobjective MTO problems
-Synthetic Functions 9 1 Any SO ST Ackley, Elliptic, Griewank, Rastrigin, Rosenbrock Schwefel, Sphere, Weierstrass
-CEC06-CSO (Liang et al., 2006) 24 1 2-24 SO ST Constrained CEC 2006 competition on constrained single-objective optimization
-CEC10-CSO (Mallipeddi and Suganthan, 2010a) 18 1 10/30 SO ST Constrained CEC 2010 competition on constrained single-objective optimization
-CEC17-CSO (Wu et al., 2017) 28 1 2-100 SO ST Constrained CEC 2017 competition on constrained single-objective optimization
-CEC17-SO (Awad et al., 2017) 29 1 2-100 SO ST CEC 2017 competition on bound constrained single-objective optimization
-CEC20-RWCO (Kumar et al., 2020) 57 1 2-158 SO ST Constrained CEC 2020 competition on real-world constrained single-objective optimization
-CEC22-SO (Kumar et al., 2022) 12 1 10/20 SO ST CEC 2022 competition on bound constrained single-objective optimization
-PEPVM (Li et al., 2023b) 1 3 7 SO MT Parameter extraction of photovoltaic models as MTO problem
-SCP (Li et al., 2022c) 1 Any 60-120 SO MT Competitive Sensor coverage problem as competitive MTO problem
-OPF (Li et al., 2024c) 5 2 24/33 SO MT Constrained Optimal power flow as constrained MTO problem
-PKACP (Jiang et al., 2023a) 1 Any Any SO MaT Planar kinematic arm control problem as many-task optimization problem
-MO-SCP (Li et al., 2023b) 2 Any 60-120 MO MT Competitive Multiobjective sensor coverage problem as competitive multiobjective MTO problem
-SOPM (Li and Gong, 2025) 2 3 30 MO MT Constrained Synchronous optimal pulse-width modulation as constrained multiobjective MTO problem
-MO-OPF (Li and Gong, 2025) 4 2 24/33 MO MT Constrained Multiobjective optimal power flow as constrained multiobjective MTO problem
-OPF-CMTMO (Li et al., 2024g) 4 2 11/24 MO MT Competitive Competitive multitask optimal power flow with thermal and wind-solar power
+The existing problems available in MToP are enumerated in Table LABEL:tab:prob-mt of the supplementary files.
 Table 4. Properties and methods of problem base class.
@@
 T Number of optimization tasks
-N Population size for each task
+N Default population size for each task
 M Number of objectives for all tasks
@@
 getOptimum() Get optimal solutions for multi-objective optimization
+evaluate() Fitness function evaluation
 1 label={list:prob},
@@
 29end
-30function optimum = getOptimum(Prob)
+30function optimum = getOptimum(Prob) % optional
 31 % return optimum points for each task
@@
 8
-9result.Metric = ’Min’;
+9 result.Metric = ’Min’;
 10
-11% Data for shown in the GUI table
-12result.RowName = {MTOData.Problems.Name};
-13result.ColumnName = {MTOData.Algorithms.Name};
-14result.TableData = CalculateTableData(MTOData);
+11 % Data for shown in the GUI table
+12 result.RowName = {MTOData.Problems.Name};
+13 result.ColumnName = {MTOData.Algorithms.Name};
+14 result.TableData = CalculateTableData(MTOData);
 15
-16% Data for shown in the GUI convergence plot
-17result.ConvergeData = CalculateConvergeData(MTOData);
+16 % Data for shown in the GUI convergence plot (optional)
+17 result.ConvergeData = CalculateConvergeData(MTOData);
 18
-19% Data for shown in the GUI Pareto plot
-20result.ParetoData = CalculateParetoData(MTOData);
+19 % Data for shown in the GUI Pareto plot (optional)
+20 result.ParetoData = CalculateParetoData(MTOData);
 21end
@@
 An illustrative example of metric implementation is presented in Listing LABEL:list:metric.
-The input parameter of the function is <MTOData>, which is generated during experimental execution.
-The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>.
-The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable.
-<RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table.
-On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively.
+The input parameter of the function is MTOData, which is generated during experimental execution.
+The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData.
+The Metric property can take values of either Min or Max, indicating whether a smaller or larger metric value is preferable.
+RowName, ColumnName, and TableData are utilized to present metric results in the GUI table.
+On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively.
 Note that specific metric calculation functions are not elaborated here for the sake of simplicity.
-The metrics currently integrated into MToP are enumerated in Table 8.
-Table 8.
-Metrics included in MToP.
-The table lists the metric name, optimization objective, task type, and a brief description.
-SO and MO denote single-objective and multi-objective optimization, respectively.
-ST and MT represent single-task and multi-task, respectively
-Metric Objective Task Description
-Obj SO ST/MT Objective value for each task
-Obj (AV) SO MT Average Obj for all tasks
-Obj (UV) SO MT Unified average Obj for all tasks
-Obj (MTS) SO MT Multitask score of Obj for all tasks
-Obj (CMT) SO MT Competitive multitask Obj for all tasks
-Obj (NBR) SO MT Number of best Obj result for all tasks
-CV SO ST/MT Constraint violation for each task
-FR SO ST/MT Feasible rate for each task
-HV MO ST/MT Hypervolume for each task
-HV (MTS) MO MT Multitask score of HV for all tasks
-HV (CMT) MO MT Competitive multitask HV for all tasks
-IGD MO ST/MT Inverted generational distance for each task
-IGD (AV) MO MT Average IGD for all tasks
-IGD (MTS) MO MT Multitask score of IGD for all tasks
-IGD (CMT) MO MT Competitive multitask IGD for all tasks
-IGD+ MO ST/MT Improved IGD plus for each task
-IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
-IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
-Spread MO ST/MT Spread metric for each task
-Spread (CMT) MO MT Competitive multitask spread for all tasks
-Run Time SO/MO ST/MT Algorithm running time for all tasks
-3. Performing Experiments with MToP
+The metrics currently integrated into MToP are enumerated in Table LABEL:tab:metric of the supplementary files.
+4. Guidelines for Using MToP
 (a) Test Module
@@
 The results of the experiment are displayed in the right section of the Experiment Module.
-To launch MToP, start by running the script file <mto.m> located in the root directory.
+To launch MToP, start by running the script file mto.m located in the root directory.
 This action will initialize the MToP GUI interface, which is illustrated in Fig.
 6.
-The GUI interface of MToP requires MATLAB R2020b or later versions to run, while command-line execution can be done with any version.
-In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
+The GUI interface of MToP requires MATLAB R2022b or later versions to run, while command-line execution can be done with any version.
 4.1. Testing Algorithms and Problems
@@
 6 (a), the Test Module interface is structured into left, center, and right sections.
-The left section facilitates the selection and configuration of algorithms and problems.
-Upon choosing the desired labels such as <Task>, <Objective>, and <Special> from the drop-down buttons, the options available for <Algorithm> and <Problem> are updated accordingly.
-Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them.
-Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method.
-Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size <N>, maximum function evaluations <maxFE>, and other specific parameters.
-In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically.
-Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
-3.2. Executing Comparative Experiments
-The GUI of the experiment module is displayed in Fig. 6 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
-In the first column, users can configure the basic experiment settings.
-The <No.
-of Runs> textbox specifies the number of independent repeat runs for each algorithm on each problem.
-The <No.
-of Results> textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations.
-The <Parallel> button toggles parallel processing, allowing independent experiments to run simultaneously across multiple process.
-The number of concurrent experiments depends on the number of processor cores.
-If the number of parallel experiments exceeds the available cores, MATLAB automatically handles the scheduling.
-For multiple algorithms and problems, each algorithm processes a problem in parallel.
-After one algorithm finishes, the next starts, repeating until all algorithms have processed the problem, then moving to the next problem.
-Hence, the number of algorithms and problems is not inherently limited.
-The drop-down buttons <Task>, <Objective>, and <Special> filter algorithms and problems based on predefined labels, similar to the test module.
-Users can add selected algorithms and problems by clicking the <Add> button after choosing them from the respective drop-down boxes at the bottom.
-In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them.
-Once all configurations are set, clicking the green <Start> button initiates the experiment execution.
-Users can also pause or stop the experiment using the <Pause> and <Stop> buttons if needed.
-Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module.
-Users can read and save the original data of the experiment using the <Load Data> and <Save Data> buttons.
-In the results area, the first function row contains buttons for operating on the experimental data.
-The first button, labeled <Save Table>, exports the table results to xlsx, csv, and tex files.
-The subsequent buttons labeled <Converge> control the convergence y-lim types and generate convergence plots based on the selected results from the table.
-The <Pareto> button draws the Pareto front and non-dominated solutions based on the selected results.
-The edit box controls the formatting type of the table result.
-Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively.
-After selecting the statistical test method, calculated statistical results denoted by “+ / - / =” followed by each metric result can also be displayed in the table.
-MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests.
-The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.
+The left section is dedicated to configuration, the center section serves as the main display for visualizations, and the right section shows the final metric results after a run.
+This module is designed for preliminary analysis, debugging, and qualitative visualization of a single algorithm’s performance on a single problem.
+4.1.1. Viewing Problem Characteristics
+Figure 7. Problem characteristics visualization in the Test Module of MToP GUI.
+Before executing an algorithm, researchers can use the Test Module to visualize the characteristics of the selected problem.
+As shown in Fig.
+7, after selecting a problem (e.g., CEC17-MTMO1-CI-HS), the central display area provides a drop-down menu (initially labeled Tasks Figure (1D Unified) in the figure).
+Clicking this menu reveals various visualization options.
+These options vary by problem type but include plotting 1D or 2D function landscapes to understand the search space, visualizing feasible regions for constrained problems, and displaying the true Pareto front for multi-objective tasks after running the algorithm.
+This allows for an initial assessment of the problem’s difficulty and features prior to optimization.
+4.1.2. Setting Algorithm and Problem Parameters
+Figure 8. Algorithm and problem parameter setting in the Test Module of MToP GUI.
+The left section of the module handles the selection and configuration of algorithms and problems.
+As illustrated in Fig.
+8, users first select the desired algorithm (e.g., MO-MFEA) and problem (e.g., CEC17-MTMO1-CI-HS) from the respective lists.
+Filtering options (e.g., Task, Objective, Special) help narrow down the choices.
+Once selected, their default parameters appear in the text boxes below.
+To modify a parameter, the user can double-click the parameter’s name in the list, which opens an editable field (as shown in the figure for Change Name and Change Parameter).
+For example, the algorithm’s name can be changed, or a specific numeric parameter like the problem’s population size N can be adjusted.
+These changes are applied to the objects when the experiment is started by pressing the Start button in the central panel.
+4.1.3. Visualizing Algorithm Behavior
+After configuring the parameters and starting the run, MToP offers dynamic display utilities such as Draw Obj and Draw Dec to enable researchers to explore algorithm behavior in real-time.
+Figure 9 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA, showcasing the evolution in both objective space and decision space.
+CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives.
+The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$.
+The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.5005$.
+Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior and delve deeper into its search dynamics.
+(a) Gen=1
+(b) Gen=200
+(c) Gen=500
+(d) Gen=700
+(e) Gen=1000
+(f) MTDE-MKTA
+(g) Gen=1
+(h) Gen=200
+(i) Gen=500
+(j) Gen=700
+(k) Gen=1000
+(l) MTDE-MKTA
+Figure 9.
+Schematic of population snapshots on CEC17-MTMO4 across generations.
+The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Frontier (PF).
+The bottom row (g)-(l) shows the decision space, where each black polyline represents the 50-dimensional decision vector of one individual in the population.
+Panels (a)-(e) and (g)-(k) depict MO-MFEA at Gen=1, 200, 500, 700, and 1000.
+Panels (f) and (l) show the example of MTDE-MKTA at Gen=1000.
+In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space.
+As evolution progresses through the first and middle phases (Gen=200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity.
+Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity.
+By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task.
+The schematics provide the key insight: the population for the second task is clustered in the similar decision space region as the population for the first task.
+However, the true optimal regions of these two tasks do not precisely overlap in the decision space.
+This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region, causing it to converge on a local optimum for Task 2.
+This visualization functionality in MToP is crucial for diagnosing such algorithmic failures.
+Furthermore, MToP can be used to validate the effectiveness of remedial approaches.
+For instance, Figure 9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue.
+In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto frontier for both tasks.
+This is further confirmed in the decision space, where the inset plot clearly shows the convergence to the correct optimal values ($0.5$ for dimensions $[2,40]$ and $0.5005$ for $[41,50]$), thus overcoming the negative transfer problem.
+In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
+4.2. Performing Comparative Experiments
+The GUI of the Experiment Module, shown in Fig.
+6 (b), features a three-column layout: the left and middle columns are for experimental setup, and the right column is for results analysis.
+The detailed steps for conducting an experiment are illustrated in Fig.
+10.
+Figure 10. Experiment execution steps in the Experiment Module of MToP GUI.
+4.2.1. Setting Experiment Parameters
+Experiment parameters are set in the Set Exp Parameter panel (Step 1).
+This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of convergence points to save per run).
+Toggles for Parallel execution and SaveDec (to save decision variables) are also located here.
+Enabling Parallel allows independent runs to execute simultaneously.
+The number of concurrent processes depends on the system’s processor cores, and MATLAB automatically handles the scheduling.
+The typical execution flow processes one problem at a time, running all added algorithms on it in parallel before moving to the next problem.
+4.2.2. Selecting Algorithms and Problems
+As shown in Fig.
+10 (Steps 2-4), the user first selects the algorithms and problems to be included in the experiment.
+To simplify this process, users can use the tag filters (Step 2) such as Task, Objective, and Constrained to narrow down the lists based on predefined labels.
+From the Algorithm List (Step 3), the user selects one or more algorithms and clicks the Add button to move them to the Added Algorithm Tree.
+An identical process is used to select problems from the Problem List and add them to the Added Problem Tree (Step 4).
+As shown in Steps 3 and 4, users can set parameters for specific algorithms and problems.
+By clicking on an item in the Added Algorithm Tree or Added Problem Tree, its individual parameters (e.g., population size N, operator parameters RMP) are displayed in the text box below, where they can be directly edited.
+To assist researchers in this selection, especially when starting a new study, Table 5 provides a curated guide to benchmark suites and high-performing algorithms.
+The listed benchmarks are widely used and recognized within the EMT community, serving as standard testbeds.
+The recommended algorithms are not arbitrary; they represent the top 5 algorithms based on the Friedman ranking of their average metric performance over 30 independent runs on each benchmark set.
+The full experimental validation for this ranking is detailed later in Section 5.1.
+This table serves as an excellent starting point for researchers and can also offer valuable insights when selecting algorithms for new real-world problems.
+Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table LABEL:tab:metric, respectively, in the supplementary files.
+Table 5.
+Recommended benchmarks, algorithms and metrics for different MTO problem categories in MToP.
+The selected algorithms are based on their Friedman rankings on each benchmark.
+Detailed references and descriptions of these algorithms and benchmarks can be found in supplementary files.
+Benchmark Recommended Algorithms (selected by Friedman rankings) Recommended Metric(s)
+Single-objective Multi-task
+CEC17-MTSO MFEA-GHS, MTES-KG, MFMP, MTDE-ADKT, MTEA-HKTS Obj
+WCCI20-MTSO MTES-KG, MFMP, MTDE-ADKT, MTEA-PAE, MFEA-DGD Obj
+Constrained Single-objective Multi-task
+CMT MTEA-PAE, MTES-KG, CEDA-MP, MFEA-GHS, MTEA-AD Obj, CV
+Competitive Single-objective Multi-task
+C2TOP&C4TOP MTSRA, MFMP, MTES-KG, MTEA-HKTS, DEORA Obj (CMT)
+C-CPLX MTES-KG, MFMP, MTSRA, MTEA-PAE, MTEA-HKTS Obj (CMT)
+Single-objective Many-task
+CEC19-MaTSO DTSKT, TNG-SNES, MTES-KG, MTEA-HKTS, SBCMAES Obj (MTS), Obj (UV)
+WCCI20-MaTSO DTSKT, TNG-SNES, MTES-KG, SBCMAES, EMaTO-MKT Obj (MTS), Obj (UV)
+Multi-objective Multi-task
+CEC17-MTMO MO-MTEA-PAE, MTDE-MKTA, MTEA-DCK, MTEA-D-DN, KR-MTEA IGD+, HV
+CEC19-MTMO MTDE-MKTA, MTEA-DCK, MTEA-D-TSD, MO-MTEA-PAE, MTEA-D-DN IGD+, HV
+CEC21-MTMO MTEA-D-TSD, MO-MTEA-PAE, MTEA-DCK, MO-MTEA-SaO, MTDE-MKTA IGD+, HV
+Competitive Multi-objective Multi-task
+CMOMT RVC-MTEA, MO-MTEA-PAE, MO-MTEA-SaO, MO-EMEA, MO-MFEA-II IGD+ (CMT), HV (CMT)
+4.2.3. Running Experiment and Calculating Metrics
+Once all configurations are set, the experiment is initiated by clicking the Start button in the Run / Load Experiment panel (Step 5).
+The Pause and Stop buttons can be used to control the execution.
+After the experiment completes, the user can move to the results panel (Step 6) to analyze the performance.
+The first step is to select a Metric (e.g., Obj) and a Show Type (e.g., Mean&Std).
+MToP then calculates the results and populates the main table.
+Table 5 also provides guidance on which metrics are most appropriate, as the choice is critical for correct analysis:
+  • For standard multi-task problems (e.g., CEC17-MTSO, CEC19-MTMO), the number of tasks is small, making it practical and recommended to analyze per-task metrics directly, such as Obj, IGD+, or HV.
+For multi-objective metrics, IGD+ is generally recommended over the traditional IGD as it provides a more comprehensive and stable measure of both convergence and distribution.
+Furthermore, the HV metric is also provided, as it is a valuable indicator that does not require a known true Pareto front, making it suitable for real-world problems where the optimum is unknown.
+  • For constrained multi-task problems (e.g., CMT), the evaluation must account for both feasibility and objective quality.
+The standard approach is to report the Obj (objective value) for runs that find at least one feasible solution, but to report the CV (constraint violation) for runs that fail to find any feasible solutions.
+  • For competitive multi-task problems (e.g., C-CPLX, CMOMT), per-task analysis is not correct, and aggregated metrics like Obj (CMT) or IGD+ (CMT) are necessary to capture the overall trade-off across tasks.
+  • For many-task problems (e.g., CEC19-MaTSO), the large number of tasks makes reporting per-task metrics impractical.
+Here, we recommend using widely-adopted summary metrics such as the multi-task score Obj (MTS) (relative performance across tasks) or the unified value Obj (UV) (adjusted absolute performance across tasks).
+  • For all problem types, the Run Time metric can be used to compare the computational efficiency and execution time of different algorithms.
+4.2.4. Performing Statistical Tests
+The results panel (Step 6) also includes a suite for statistical analysis.
+MToP provides two distinct non-parametric approaches for performance comparison: direct pairwise tests and a global omnibus test with post-hoc analysis.
+First, for direct pairwise comparisons, MToP offers the Wilcoxon rank-sum test and the Wilcoxon signed-rank test.
+When a user selects one of these methods, MToP performs a direct statistical comparison between each algorithm and a user-selected Base Algorithm.
+MToP’s logic then annotates the results table with “+” (significantly better), “-” (significantly worse), or “=” (no significant difference) based on a $p$-value threshold of $0.05$.
+It is important to note that these are uncorrected pairwise tests intended for exploratory analysis.
+Second, for a more robust global comparison, MToP provides the Friedman test.
+This is an omnibus test that first checks for any statistically significant differences among the entire set of selected algorithms.
+Users can choose to run this test on the mean results across all runs (mean) or on the complete dataset combining all runs (all reps).
+Upon selection, MToP automatically performs the global Friedman test and then immediately conducts a post-hoc test.
+This post-hoc analysis calculates the $z$-value and corresponding $p$-value to compare the mean rank of each algorithm against the mean rank of the Base Algorithm.
+The results table is then populated with the mean rank for each algorithm and the $p$-value for its comparison against the base, allowing for a statistically grounded, rank-based assessment.
+It is a known characteristic of null-hypothesis significance testing that with a sufficiently large number of replications, even trivial performance differences can be deemed statistically significant.
+This does not necessarily imply practical significance.
+MToP is therefore designed to facilitate a multi-faceted analysis.
+MToP provides both relative performance comparisons (via the Wilcoxon and Friedman statistical test results) and crucial absolute performance measures (the Mean&Std values).
+Researchers are thus equipped to interpret the statistical results in conjunction with the absolute performance data and the MToP’s visualizations (e.g., Converge plots) to make a more informed judgment about the practical significance of the observed differences.
+4.2.5. Exporting Experiment Data and Results
+MToP provides two primary methods for saving data, located in the Run / Load Experiment panel (Fig.
+10, Step 5).
+The Save Data and Load Data buttons are used to save or reload the entire experimental session (the raw MTOData object), which allows for experiments to be reloaded and analyzed later.
+For saving processed results, the Export button provides more granular control. Clicking this button opens a new dialog window offering three distinct export options:
+  • Current Table (tex, xlsx, csv): This option saves the data exactly as it is currently displayed in the main results table, including the calculated metrics and any statistical annotations (e.g., “+ / - / =” and “Ranking”).
+It supports tex, xlsx, and csv formats for easy integration into publications and spreadsheets.
+  • IOHanalyzer Data (csv): This option exports the complete convergence history (not just the final values) for all selected runs into a csv format.
+This format is specifically structured for use with the IOHanalyzer (Wang2022IOHanalyzer), facilitating advanced external analysis.
+  • Best Decision Variable (mat): This option extracts the decision variables of the best-found solutions of all algorithms and repetitions and saves them to a .mat file for further inspection or use in other applications.
+For single-objective MTO problems, the best solution is determined by the lowest objective value.
+For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu2024Finding; Choong2022Jack).
+4.2.6. Visualizing Metric Convergence and Pareto Front
+Finally, the module offers integrated plotting tools at the bottom of the results panel (Step 6).
+By selecting one or more rows and columns in the results table, the user can generate plots.
+Clicking the Converge button generates a convergence plot for the selected metric results, with a Log toggle available to switch the y-axis to a logarithmic scale, and a Range toggle to display the 95% confidence interval around the mean curve.
+For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto fronts).
+Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang2022IOHanalyzer) for further external analysis and visualization.
+4.3. Processing Experiment Data
+MToP provides a dedicated Data Process Module within the GUI for merging and splitting experimental data files.
+This module allows users to combine multiple MTOData.mat files into a single comprehensive dataset or to partition a large dataset into smaller, more manageable segments based on specific criteria such as algorithms, problems, or repetitions.
+The merging function is particularly useful for aggregating results from separate experimental runs, while the splitting function aids in isolating subsets of data for focused analysis.
+Users can access this module directly from the main GUI, where they can select the desired operation (merge or split), specify input files, and define output parameters.
+The processed data is saved in the standard MTOData.mat format, ensuring compatibility with other modules within MToP.
+Figure 11. Data processing steps in the Data Process Module of MToP GUI.
+4.3.1. Loading and Viewing Data Files
+As shown in the Load / Save Data panel of Fig.
+11, the user begins by clicking the Load Data button.
+This opens a file dialog to select one or more MTOData.mat files, which are then added to the Data Tree.
+By clicking on a loaded data object, its high-level structure (e.g., Reps, Algorithms, Problems) is displayed.
+Users can further click on these fields, such as Problems, to view the detailed contents in a table, allowing for verification of the data before processing.
+The Delete Data button can also be used to remove any unwanted or intermediate data objects from the Data Tree.
+4.3.2. Splitting Experiment Data
+The Split Data panel in Fig.
+11 illustrates the partitioning functionality.
+To divide a large dataset, the user selects a data object from the Data Tree.
+MToP provides three splitting options: Reps Split, Algorithm Split, and Problem Split.
+For instance, selecting Algorithm Split will parse the dataset and create new, separate data objects for each algorithm it contains (e.g., MTOData (Split Algorithm: MFEA)).
+These new objects then appear in the Data Tree, ready for isolated analysis or saving.
+4.3.3. Merging Experiment Data
+Conversely, the Merge Data panel demonstrates how to combine datasets.
+The user selects two or more data objects from the Data Tree that share compatible settings.
+For example, to aggregate independent runs, the user can select multiple files and click Reps Merge.
+This creates a single new dataset (e.g., Data (Merge Reps)) with a larger total number of repetitions.
+The Algorithm Merge and Problem Merge buttons function similarly, allowing for the consolidation of data across algorithms or problems.
+Note that when merging, MToP checks for consistency in experiment settings to ensure valid combinations (e.g., same problems and reps when merging algorithms).
+4.3.4. Adjusting Data Precision
+The module also provides tools for changing data precision.
+Users can select a data object and specify the desired number of decimal places for metric values.
+This is particularly useful for reducing file size or standardizing data formats before saving.
+4.3.5. Saving Processed Data
+After all processing operations (splitting, merging, or reordering) are complete, the new or modified data objects can be exported.
+As shown in the Load / Save Data panel, the user selects the desired data object from the Data Tree (e.g., Data (Merge Reps)) and clicks the Save Data button to save it as a new MTOData.mat file.
 4.4. Executing via Alternative Command Line
-In addition to GUI execution, MToP offers alternative options for conducting experiments via the command line.
-By inputting the algorithm list, problem list, and other parameters into the script file <mto.m>, experiments can be executed directly from the command line.
-For example, running the command mto({MFEA, EMEA}, {CMT1, CMT2}, 30) will execute the MFEA and EMEA algorithms on the CMT1 and CMT2 problems independently 30 times.
-Upon completion, the experiment data <MTOData.mat> is returned.
-Similarly, by inputting the data into the metrics function, users can obtain the calculated metric results.
-4. Extending MToP
-Users can seamlessly integrate new algorithms, problems, and performance metrics into MToP.
-By adhering to the code pattern outlined in Section 2.3, users can develop and implement their code files with ease.
-4.1. Adding New Algorithms
-To incorporate a new algorithm into MToP, users should first generate a MATLAB class file within the <Algorithms/> folder.
-This class file should inherit from the base class <Algorithm.m>.
-Subsequently, users can implement the algorithmic framework following the provided code pattern, as demonstrated in Listing LABEL:list:algo.
-Utilizing functions such as Initialization(), Generation(), and Selection() provided by MToP is recommended.
-However, users can also choose to reimplement these functions to suit their specific requirements.
-Additionally, users seeking to create customizable parameters that can be modified via the MToP GUI should implement the getParameter() and setParameter() APIs within the new algorithm class file.
-4.2. Adding New Problems
-To integrate new problems into MToP, users should create a dedicated problem class file within the <Problems/> folder.
-This class file should inherit from the base class <Problem.m>.
-It’s advisable to define the suggested maximum function evaluations <maxFE> within the constructor of the specific problem.
-Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function.
-In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points.
-Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
-4.3. Adding New Performance Metrics
-To incorporate new performance metrics into MToP, users should create a new metrics function file within the <Metrics/> folder.
-Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object <MTOData> and return the calculated <result> object.
-For the metrics displayed in the GUI table, ensure that the <RowName>, <ColumnName>, and <TableData> fields are properly populated.
-Typically, the <RowName> for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task.
-The <ColumnName> typically represents the algorithm name.
-Implement the calculation function for the specific performance metric in the <TableData> field.
-For plotting metric convergence, ensure the <ConvergeData> field is populated.
-In the case of multi-objective optimization, include non-dominated solutions in the <ParetoData> field.
-4.4. Extending Optimization Types
-There are also some areas, such as sequential transfer optimization and multiform optimization, in the evolutionary transfer optimization field similar to evolutionary multitasking (Tan et al., 2021).
-Although MToP is currently focused on evolutionary multitasking, it is possible to extend the platform to include these new optimization types.
-Users can refer to the existing code patterns in MToP to develop new optimization types.
-Specifically, for the sequential transfer optimization, it is possible to set the target task as the first task and the remaining tasks as source tasks for algorithms and problems.
-The metric calculation can be adjusted to evaluate the performance of the target task.
-For multiform optimization, it can dynamically construct multiform tasks within the implemented algorithms depending on the problem to be solved.
-5. Conclusion and Future Work
+For users who require batch processing, integration with external scripts, or operation in non-GUI environments, MToP provides a comprehensive and flexible command-line interface.
+The main mto.m script has been created to serve as a dual-purpose entry point: calling mto with no arguments launches the GUI, while calling it with arguments executes a command-line experiment.
+The function MTOData = mto(varargin) is designed for maximum flexibility, accepting both positional arguments and Name-Value pairs.
+For instance, users can run basic experiments using commands like result = mto(MFEA, CMT1) for quick runs.
+More complex configurations such as mto({MFEA, EMEA}, {CMT1, CMT2}, ’Reps’, 30, ’Par_Flag’, true, ’Save_Name’, ’A.mat’) allow for detailed customization, including parallel execution and specifying output filenames.
+Upon completion, the function returns the standard MTOData struct and simultaneously saves it to the specified .mat file.
+This command-line interface workflow extends beyond just running experiments; the returned MTOData object can be directly passed to any metric function to calculate results, just as in the GUI.
+This enables a complete, script-based workflow: setup, run, and analyze.
+A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD, and programmatically plot the final convergence curves and Pareto fronts is provided in the cmd_examples.m script located in MToP’s root directory.
+5. Validation and Comparison
+This section first performs reproducible validation tests to ensure that the algorithms and problems implemented in MToP are reliable and accurate.
+Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian2017PlatEMO), is conducted to demonstrate MToP’s comparable efficiency.
+5.1. Reproducible Validation
+A core requirement for a scientific platform is the reliability and reproducibility of its components.
+To validate the implementations of the extensive algorithm and problem libraries within MToP, we conducted a comprehensive experimental study, running all algorithms (including both MTEAs and single-task EAs) on all compatible benchmark suites for 30 independent runs.
+To ensure fair comparisons and high reproducibility, a global random seed controller is employed.
+For each independent repetition rep, the random number generator is explicitly seeded using rng(Global_Seed + rep - 1).
+This ensures that all algorithms within the same repetition begin with identical random conditions, which is crucial for reliable comparisons.
+In order to enhance reproducibility for the wider community and to provide a baseline for future research, we have made the complete raw experimental data from this large-scale validation publicly available.
+This dataset validates our implementations and can be used directly by other researchers to avoid redundant, computationally expensive experiments.
+All data can be accessed and downloaded.
+[^3] Pre-run experiment data can be found at:
+  Google Drive: https://drive.google.com/drive/folders/1IpwXNuOYcnpC3IXbAx3VnLGVV899k3bG?usp=share_link
+  Hugging Face: https://huggingface.co/datasets/intLyc/MToP-MTOData/tree/main
+  Baidu Netdisk: https://pan.baidu.com/s/1Pk06fBj_4gidkiZe4f1Oww?pwd=mtop
+5.2. Comparison with Other Platforms
+Figure 12. Running time comparison between MToP and PlatEMO on DTLZ problems using algorithms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA.
+Beyond internal validation, it is crucial to benchmark MToP’s performance and efficiency against established, state-of-the-art platforms.
+We selected PlatEMO (Tian2017PlatEMO), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison.
+To ensure a fair and direct comparison, we selected four popular and well-understood single-task multi-objective algorithms that are implemented in both platforms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA.
+We executed these algorithms on the standard DTLZ1-7 benchmark suite.
+Both platforms were run under identical experimental settings, including 30 independent runs and the use of parallel execution.
+First, we compared the computational efficiency.
+Figure 12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite.
+The results show that the execution times for MToP and PlatEMO are highly comparable.
+For NSGA-II, SPEA2, and MOEA-D, the run times are nearly identical.
+For SMS-EMOA, which involves computationally intensive hypervolume calculations, both platforms show similar overhead.
+This demonstrates that MToP’s architecture is lightweight and efficient, with no significant computational overhead compared to PlatEMO.
+Second, we validated the correctness of our algorithmic implementations by comparing the final solution quality.
+Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD metric for 30 runs.
+A Wilcoxon rank-sum test (at a 0.05 significance level) was conducted between the MToP and PlatEMO results for each case.
+The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (23 out of 28).
+The few minor differences are likely due to small implementation-level variations, but the overall performance is statistically indistinguishable.
+This similarity in performance is further confirmed visually in Figure LABEL:fig:pareto-comparison, which plots the final Pareto fronts from a sample run for each algorithm.
+The acquired fronts from both platforms are visually coincident.
+6. Discussion and Outlook
 This paper introduces MToP, an open-source MATLAB platform designed for EMT.
@@
 The current version of MToP contains over 50 MTEAs, more than 50 single-task EAs capable of handling MTO problems, over 200 benchmark MTO problems, and several real-world applications of EMT.
-This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
-While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs.
+This paper provides recommendations for selecting algorithms and benchmarks on different types of MTO problems.
+It also offers guidelines for using MToP, including testing algorithms and problems, executing comparative experiments, processing experiment data, and running via the command line.
+Finally, this paper validates the reproducibility of implemented algorithms and problems in MToP, provides open-sourced experimental data, and compares the running time and results of MToP with other EC platforms.
+While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some implementation errors and bugs.
 Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub.
-The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
+MToP encourages feedback and contributions from users and researchers, with many codes already received and integrated from the community.
 Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community.
 It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field.
-Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021c, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
+Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng2016RVEA; Deb2014NSGA3), multitask combinatorial optimization codes (Feng2021VRPHTO; Feng2021EEMTA) and more real-world applications (Gupta2021Half; Huang2023Evolutionary) in the future.
 Acknowledgements.
-This work was partly supported by the National Natural Science Foundation of China under Grant No.
-62076225 and 62403477, the Open Research Fund of the State Key Lab of Digital Manufacturing Equipment & Technology under Grant No.
-IMETKF2024020, and the Major Project of Hubei Province Science and Technology under Grant No.
-2023BCA006.
+This work was partly supported by the National Natural Science Foundation of China (Grant No.
+62076225), the Major Research & Development Program of Hubei Province (Grant No.
+2025BEB002 & 2023BCA006), the Regional Innovation System Program of Hubei Province (Grant No.
+2025EIA022), and the Hubei Superior and Distinctive Discipline Group of “New-Energy Vehicles and Smart Transportation”.
 All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles.
@@
 The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes.
-Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.
+Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian2017PlatEMO), for which the authors are also thankful.

### v6

@@
 Evolutionary algorithms (EAs), the cornerstone of EC, have demonstrated remarkable success in addressing black-box optimization problems due to their robustness and ease of implementation.
-Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang2020CORCO; Tian2021CCMO), multi-objective optimization (Zhang2007MOEA-D; Deb2014NSGA3), and combinatorial optimization (Feng2021VRPHTO; Feng2021EEMTA).
-In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei2021Review-MTO).
+Researchers have dedicated significant efforts to designing tailored EAs for various complex black-box optimization problems, including constrained optimization (Wang et al., 2020; Tian et al., 2021), multi-objective optimization (Zhang and Li, 2007; Deb and Jain, 2014), and combinatorial optimization (Feng et al., 2021b, a).
+In recent years, driven by escalating computational demands and the emergence of cloud computing, there has been a growing emphasis on utilizing EAs to tackle multiple optimization tasks concurrently, known as evolutionary multitasking (Wei et al., 2022).
 A multitask optimization (MTO) problem within the realm of EMT, comprising $K$ minimization tasks, aims to find solutions $(\bm{x}_{1}^{\ast},\bm{x}_{2}^{\ast},...,\bm{x}_{K}^{\ast})$ for all tasks, which can be formulated as follows:
@@
 Note that for a multi-objective multi-task optimization problem, the $\bm{F}_{k}$ contains multiple objective functions $(f_{1},f_{2},...,f_{M_{k}})$ and $\bm{x}_{k}^{\ast}$ becomes a set of Pareto sets for the $k$-th task.
-For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta2016MFEA) as follows:
+For task $k$ with upper and lower bounds $\bm{L}_{k}$ and $\bm{U}_{k}$ and dimension $D_{k}$, the solution $\bm{x}$ of its decision space is typically mapped to the unified search space (Gupta et al., 2016) as follows:
 (2) $\bm{x}^{\prime}=\frac{\bm{x}-\bm{L}_{k}}{\bm{U}_{k}-\bm{L}_{k}}.$
@@
 During objective function evaluation, dimensions and upper and lower bounds are linearly reproduced without loss of precision.
-EMT has found successful applications in various domains, including engineering scheduling (Feng2021VRPHTO; Gupta2021Half), nonlinear equation systems (Li2024Evolutionary), feature selection in machine learning (Chen2021Evolutionary), anomaly detection (Wang2022AUC), point cloud registration (Wu2024Evolutionary; Wu2023Evolutionary), and reinforcement learning (Zhang2023Multitask).
-To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan2023Knowledge; Gupta2018Insights).
-Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan2021ETO; Li2024MFEA-RL).
-The first attempt of EMT can be traced back to the multifactorial EA (Gupta2016MFEA), which introduced an implicit knowledge representation via random mating among optimization tasks.
+EMT has found successful applications in various domains, including engineering scheduling (Feng et al., 2021b; Gupta et al., 2022), nonlinear equation systems (Li et al., 2024a), feature selection in machine learning (Chen et al., 2022), anomaly detection (Wang et al., 2022b), point cloud registration (Wu et al., 2024a, b), and reinforcement learning (Zhang et al., 2023).
+To expedite and enhance the concurrent resolution of multiple optimization tasks, researchers have endeavored to leverage task similarity to augment EAs with knowledge extraction and transfer techniques (Tan et al., 2023; Gupta et al., 2018).
+Through knowledge transfer, EAs can effectively exploit implicit parallelism to achieve superior solutions across multiple tasks while conserving computational resources (Tan et al., 2021; Li et al., 2024b).
+The first attempt of EMT can be traced back to the multifactorial EA (Gupta et al., 2016), which introduced an implicit knowledge representation via random mating among optimization tasks.
 Subsequently, numerous EAs tailored for MTO have emerged.
-These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali2020MFEA2; Zhou2021MFEA-AKT), or a multi-population framework allocating separate populations for each task (Li2020MFMP; Li2023MTEA-SaO; Li2024MTES-KG).
-Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta2016MFEA; Zhou2021MFEA-AKT), auto-encoding (Feng2019EMEA; Zhou2021Learnable; Gu2025MTEA-PAE), affine transformation (Xue2022AT-MFEA; Lin2024Ensemble), and adversarial generative models (Liang2023EMT-GS) have been proposed in MTEAs.
-Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali2020MFEA2; Jiang2023BoKT), knowledge selection (Wang2022MTEA-AD; Liang2022EMaTO-MKT; Li2025MTDE-MKTA), and historical feedback (Li2022CompetitiveMTO; Lin2021EMT-ET; Li2023MTSRA), have also been investigated.
+These multitask evolutionary algorithms (MTEAs) adopt either a multifactorial framework utilizing a single population for multiple tasks (Bali et al., 2020; Zhou et al., 2021b), or a multi-population framework allocating separate populations for each task (Li et al., 2020, 2023b, 2024d).
+Moreover, to facilitate decision space mapping across different tasks, various techniques such as unified search space (Gupta et al., 2016; Zhou et al., 2021b), auto-encoding (Feng et al., 2019; Zhou et al., 2021a; Gu et al., 2025), affine transformation (Xue et al., 2022; Lin et al., 2024), and adversarial generative models (Liang et al., 2023) have been proposed in MTEAs.
+Given the significant impact of knowledge transfer on solving MTO problems, MTEAs with adaptive control strategies for knowledge transfer, such as similarity judgment (Bali et al., 2020; Jiang et al., 2023), knowledge selection (Wang et al., 2022a; Liang et al., 2022; Li and Gong, 2025), and historical feedback (Li et al., 2022a; Lin et al., 2021; Li et al., 2023a), have also been investigated.
 Despite the proliferation of MTEAs proposed by researchers, there is currently no standardized programming language, code pattern, or software platform for EMT source codes.
@@
 The selection of MATLAB is based on several strategic factors.
-First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta2016MFEA) and MFEA-II (Bali2020MFEA2), were developed and open-sourced in MATLAB.
+First, many foundational and seminal works in the EMT field, including the original MFEA (Gupta et al., 2016) and MFEA-II (Bali et al., 2020), were developed and open-sourced in MATLAB.
 This established a strong precedent and a rich ecosystem of existing code for the community.
-Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian2017PlatEMO), provides a high-quality open-source reference.
+Second, the prominence of other successful EC platforms in MATLAB, most notably PlatEMO (Tian et al., 2017), provides a high-quality open-source reference.
 This synergy allows for the adaptaion and reuse of well-vetted modules, such as those for multi-objective optimization, which is crucial for the multi-objective MTO subfield.
@@
 In the field of EC, several popular and successful software platforms have significantly contributed to the development of EAs and evolutionary optimization:
-  • PlatEMO (Tian2017PlatEMO): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
-  • EDOLAB (Peng2023EDOLAB): A MATLAB-based platform focusing on evolutionary dynamic optimization.
-  • IOHprofiler (Doerr2018IOHprofiler): A modular framework, including IOHexperimenter (deNobel2024IOHexperimenter) and IOHanalyzer (Wang2022IOHanalyzer), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
-  • EvoX (Huang2024EvoX): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
-  • PyPop7 (Duan2024PyPop7): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
-  • MetaBox (Ma2023MetaBox): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
+  • PlatEMO (Tian et al., 2017): A comprehensive MATLAB platform tailored for evolutionary multi-objective optimization.
+  • EDOLAB (Peng et al., 2023): A MATLAB-based platform focusing on evolutionary dynamic optimization.
+  • IOHprofiler (Doerr et al., 2018): A modular framework, including IOHexperimenter (de Nobel et al., 2024) and IOHanalyzer (Wang et al., 2022c), for the detailed benchmarking and analysis of iterative optimization heuristics, with a primary focus on single-objective and multi-objective problems.
+  • EvoX (Huang et al., 2024): A distributed GPU-accelerated library focused on scalability and expediting complex optimization and reinforcement learning tasks.
+  • PyPop7 (Duan et al., 2024): A pure-Python library for population-based single-objective black-box optimization, particularly for large-scale optimization.
+  • MetaBox (Ma et al., 2023): A Python-based platform for meta-black-box optimization, which uses meta-learning to design optimizers, with its latest version using MTO as a training scenario.
 Table 1. Comparison of core focus areas among popular evolutionary computation platforms.
 Platform Language Core Focus MTO Support
-PlatEMO (Tian2017PlatEMO) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
-EDOLAB (Peng2023EDOLAB) MATLAB Evolutionary Dynamic Optimization No
-IOHprofiler (Doerr2018IOHprofiler) C++/Python/R Benchmarking Iterative Optimization Heuristics No
-EvoX (Huang2024EvoX) Python Distributed GPU-accelerated Optimization No
-PyPop7 (Duan2024PyPop7) Python Large-scale Black-box Optimization No
-MetaBox (Ma2023MetaBox) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
+PlatEMO (Tian et al., 2017) MATLAB Evolutionary Multi-objective Optimization Partial (v4)
+EDOLAB (Peng et al., 2023) MATLAB Evolutionary Dynamic Optimization No
+IOHprofiler (Doerr et al., 2018) C++/Python/R Benchmarking Iterative Optimization Heuristics No
+EvoX (Huang et al., 2024) Python Distributed GPU-accelerated Optimization No
+PyPop7 (Duan et al., 2024) Python Large-scale Black-box Optimization No
+MetaBox (Ma et al., 2023) Python Meta-black-box Optimization via Meta-learning As a test scenario (v2)
 MToP (This work) MATLAB Evolutionary Multitask Optimization Yes (Native)
@@
 21, 2021 at: https://github.com/intLyc/MTO-Platform, a platform with native, comprehensive support for MTO benchmarking was absent.
-The most prominent existing platform, PlatEMO (Tian2017PlatEMO), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization.
+The most prominent existing platform, PlatEMO (Tian et al., 2017), has a core architecture that is multi-objective optimization-centric, meticulously designed for solving single-task multi-objective optimization.
 This presents a fundamental architectural mismatch for EMT.
@@
 The other platforms listed in Table 1 are similarly specialized for different, non-MTO-centric goals.
-Moreover, EMT has been extended to the subfields of many-task optimization (Chen2020MaTDE; Liaw2019SBO; Li2024TNG-NES), multi-objective MTO (Gupta2017MO-MFEA; Li2025MTEA-DCK), competitive MTO (Li2022CompetitiveMTO; Li2023MTSRA; Li2024CMO-MTO), and constrained multi-task optimization (Li2022CMTO-Benchmark; Zhang2024CEDA).
+Moreover, EMT has been extended to the subfields of many-task optimization (Chen et al., 2020; Liaw and Ting, 2019; Li et al., 2024c), multi-objective MTO (Gupta et al., 2017; Li et al., 2025), competitive MTO (Li et al., 2022a, 2023a, 2024e), and constrained multi-task optimization (Li et al., 2022b; Zhang et al., 2024).
 This growing complexity further necessitates a dedicated, MTO-native framework.
@@
 A practical guide on their specific operation and usage is provided in Section 4.
+Notably, MToP is architecturally designed without any inherent software limitation on the number of tasks.
+Users can programmatically define benchmarks with an arbitrary number of tasks without hard-coded limits.
 3.1.1. Test Module
@@
 All generated tables and test results can be exported to .xlsx, .csv, or .tex formats.
-Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang2022IOHanalyzer) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
+Furthermore, it provides granular export options for specific use cases, such as exporting convergence data in .csv format compatible with IOHanalyzer (Wang et al., 2022c) or saving the optimal solution (single-objective MTO) or set of Pareto sets (multi-objective MTO) to .mat files for external study.
 Finally, the module provides essential visualization capabilities for interpreting the experimental outcomes.
@@
 2 (a)), a function applicable to both single- and multi-objective problems.
-Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto fronts (Fig.
+Furthermore, for multi-objective MTO tasks, it can plot the acquired Pareto front approximations (Fig.
 2 (b)) to visually assess the quality of the final solution sets.
@@
   • Problem class is an abstract base for optimization tasks, defining the essential evaluate() method and problem parameters ($T$ for task, $N$ for population size, $M$ for objective, $D$ for dimension).
-  • Function module groups static utility functions such as metrics (e.g., IGD()) and evolutionary operators (e.g., GA_Crossover()).
+  • Function module groups static utility functions such as metrics (e.g., IGDp(), i.e., IGD+) and evolutionary operators (e.g., GA_Crossover()).
 3.2.2. Sequence Workflow
-Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm, Problem, and Metric.
-Fig. 4 depicts the dynamic interaction workflow between the main components when running an experiment:
+Figure 4. Sequence diagram of MToP. The main workflow involves MTO_GUI / MTO_CMD, Algorithm (MFEA), Problem (CMT1), and Metric (Obj).
+Fig. 4 illustrates the dynamic interaction workflow using a concrete example of the MFEA algorithm solving the CMT1 problem:
   • The process is initiated by a User through either the MTO_GUI or MTO_CMD interface.
-  • Then, the interface handles the setup phase, configuring the Algorithm and Problem instances by calling setParameter() and setTasks().
-Next, the interface invokes the run() method of the Algorithm.
-The Algorithm performs its Initialization() and then enters the main optimization loop, conditioned on while notTerminated().
-  • Inside the loop, the Algorithm executes its Evaluation() method, which subsequently calls the evaluate(individual) method on the Problem object.
-The Problem returns the calculated Obj, Con values.
-Following evaluation, the Algorithm calls EvolvePopulation() to generate the next population.
-  • Once the loop terminates, the Algorithm returns the bestResult to the GUI/CMD interface.
-The interface then passes this result to the Metric module for Calculation().
-Finally, the computed metricValue is returned and displayed to the user.
+  • The interface configures the specific algorithm (e.g., MFEA) and problem (e.g., CMT1) instances by calling setParameter() and setTasks().
+Subsequently, the interface invokes the run() method of MFEA, which performs Initialization() and enters the main optimization loop conditioned on while notTerminated().
+  • Inside the loop, the workflow first executes the Generation() phase. Here, the algorithm calls specific operators, such as GA_Crossover() and GA_Mutation(), to produce offspring.
+  • Following generation, the Evaluation() method is triggered.
+This invokes the evaluate() method on the CMT1 object, which returns the objective and constraint values (Obj, Con).
+Finally, the Selection() strategy is applied to determine the survivors for the next generation.
+  • Once the loop terminates, MFEA returns the bestResult to the interface.
+The interface then passes this result to the metric module (e.g., Obj) for Calculation().
+Finally, the computed metricValue is displayed to the user.
 3.2.3. File Structure
@@
   • Metrics/: it contains all result evaluation metrics, organized into subfolders for single-objective Metrics/Single-objective/ and multi-objective Metrics/Multi-objective optimization.
-Specific metric files include Obj.m, Obj_MTS.m, and IGD.m, among others.
+Specific metric files include Obj.m, Obj_MTS.m, and IGDp.m, among others.
   • GUI/: it contains all files used by the GUI of MToP.
@@
   • Results: This is the $P\times A\times R$ struct array containing the core optimization output.
-Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved convergence points (where $G$ is the Results Num).
+Each struct MTOData.Results(i, j, k) contains the detailed results for a single run, aggregated across all $T$ tasks and all $G$ saved checkpoints (where $G$ is the Results Num).
 Each struct contains:
     • Obj: Stores the objective values.
-For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each convergence point.
+For single-objective problems, this is a $T\times G$ numerical array storing the single best value for each task at each checkpoint.
 For multi-objective problems, this is a $T\times 1$ cell array, where each cell Obj{t} contains a $G\times N\times M_{t}$ array representing the full objective values of $N$ solutions.
@@
 Moreover, for multi-objective optimization, the external function getOptimum() is provided.
-It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD.
+It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD+.
 For problems where the true front is unknown (e.g., real-world applications), this function is instead used to provide the reference point for calculating metrics such as hypervolume (HV).
@@
 Note that specific metric calculation functions are not elaborated here for the sake of simplicity.
-The metrics currently integrated into MToP are enumerated in Table LABEL:tab:metric of the supplementary files.
+The metrics currently integrated into MToP are enumerated in Table 5.
+Table 5.
+Metrics in MToP.
+Abbreviations: SO/MO = single-/multi-objective; ST/MT = single-/multi-task; CV = constraint violation; FR = feasible rate; HV = hypervolume; IGD = inverted generational distance; IGD+ = improved IGD plus; AV/UV = average/unified average; MTS = multitask score; CMT = competitive multitask; NBR = number of best results.
+Metric Objective Task Description
+Obj SO ST/MT Objective value for each task
+Obj (AV) SO MT Average Obj for all tasks
+Obj (UV) SO MT Unified average Obj for all tasks
+Obj (MTS) SO MT Multitask score of Obj for all tasks
+Obj (CMT) SO MT Competitive multitask Obj for all tasks
+Obj (NBR) SO MT Number of best Obj result for all tasks
+CV SO ST/MT Constraint violation for each task
+FR SO ST/MT Feasible rate for each task
+HV MO ST/MT Hypervolume for each task
+HV (MTS) MO MT Multitask score of HV for all tasks
+HV (CMT) MO MT Competitive multitask HV for all tasks
+IGD MO ST/MT Inverted generational distance for each task
+IGD (AV) MO MT Average IGD for all tasks
+IGD (MTS) MO MT Multitask score of IGD for all tasks
+IGD (CMT) MO MT Competitive multitask IGD for all tasks
+IGD+ MO ST/MT Improved IGD plus for each task
+IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
+IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
+Spread MO ST/MT Spread metric for each task
+Spread (CMT) MO MT Competitive multitask spread for all tasks
+Run Time SO/MO ST/MT Algorithm running time for all tasks
 4. Guidelines for Using MToP
@@
 After configuring the parameters and starting the run, MToP offers dynamic display utilities such as Draw Obj and Draw Dec to enable researchers to explore algorithm behavior in real-time.
-Figure 9 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA, showcasing the evolution in both objective space and decision space.
+Fig.
+9 illustrates an example of population variation of MO-MFEA (Gupta et al., 2017) on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA (Li and Gong, 2025), showcasing the evolution in both objective space and decision space.
 CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives.
@@
 The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.5005$.
-Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior and delve deeper into its search dynamics.
+Before analyzing the visual patterns, it is essential to formally distinguish the knowledge transfer mechanisms of the compared algorithms to understand the underlying causes of their behaviors.
+MO-MFEA relies on implicit knowledge transfer via assortative mating.
+Let $\bm{x}_{i}$ and $\bm{x}_{j}$ be parents from different tasks.
+The offspring $\bm{u}_{i}$ is generated using the simulated binary crossover, which directly mixes the decision variables:
+(3) $\bm{u}_{i,d}=\begin{cases}0.5[(1+\gamma)\bm{x}_{i,d}+(1-\gamma)\bm{x}_{j,d}]&\text{if }r\leq 0.5\\ 0.5[(1-\gamma)\bm{x}_{i,d}+(1+\gamma)\bm{x}_{j,d}]&\text{otherwise}\end{cases}$
+where $\gamma$ is the spread factor and $r$ is a random number. This mechanism lacks explicit domain adaptation, which may lead to negative transfer if the optimal regions of the tasks are misaligned.
+In contrast, MTDE-MKTA employs an explicit knowledge transfer strategy based on evolution path maintenance.
+It models the population distributions of the source and target tasks as Gaussian distributions $\mathcal{N}(\bm{\mu}_{s},\bm{\sigma}_{s})$ and $\mathcal{N}(\bm{\mu}_{t},\bm{\sigma}_{t})$, respectively.
+A solution $\bm{x}$ from the source task is explicitly transformed into a candidate solution $\bm{y}$ for the target task by aligning their statistical distributions:
+(4) $\bm{y}=\bm{\mu}_{t}+\frac{\bm{\sigma}_{t}\circ(\bm{x}-\bm{\mu}_{s})}{\bm{\sigma}_{s}}$
+where $\circ$ denotes element-wise multiplication and division.
+This transformation adapts the transferred solution to the target task’s search space, thereby correcting the decision variable bias and mitigating negative transfer.
 (a) Gen=1
@@
 Schematic of population snapshots on CEC17-MTMO4 across generations.
-The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Frontier (PF).
-The bottom row (g)-(l) shows the decision space, where each black polyline represents the 50-dimensional decision vector of one individual in the population.
+The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Front.
+The bottom row (g)-(l) shows the 50-dimensional decision space.
+The horizontal axis represents the dimension index (1 to 50), and the vertical axis represents the unified decision variable value.
+Each individual in the population is visualized as a polyline connecting its values across all dimensions.
+The dense overlap of these polylines (appearing as black bands or lines) visually reflects the convergence consistency and the value distribution of the entire population.
 Panels (a)-(e) and (g)-(k) depict MO-MFEA at Gen=1, 200, 500, 700, and 1000.
 Panels (f) and (l) show the example of MTDE-MKTA at Gen=1000.
+Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior.
 In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space.
@@
 Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity.
-By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task.
+By Gen=1000, MO-MFEA achieves the optimal Pareto front on the first task but becomes trapped in local optima on the second task.
 The schematics provide the key insight: the population for the second task is clustered in the similar decision space region as the population for the first task.
 However, the true optimal regions of these two tasks do not precisely overlap in the decision space.
-This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region, causing it to converge on a local optimum for Task 2.
-This visualization functionality in MToP is crucial for diagnosing such algorithmic failures.
+This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region.
+This confirms the limitation of the implicit transfer mechanism (Eq.
+(3)) discussed above: without domain adaptation, the direct mixing of variables drags the population towards the wrong attractor.
 Furthermore, MToP can be used to validate the effectiveness of remedial approaches.
-For instance, Figure 9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue.
-In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto frontier for both tasks.
-This is further confirmed in the decision space, where the inset plot clearly shows the convergence to the correct optimal values ($0.5$ for dimensions $[2,40]$ and $0.5005$ for $[41,50]$), thus overcoming the negative transfer problem.
-In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
+For instance, Fig.
+9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue.
+In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto front for both tasks.
+This validates the efficacy of the explicit transformation (Eq.
+(4)), which successfully re-mapped the guiding solutions to the correct decision space region ($0.5005$), thus overcoming the negative transfer problem.
 4.2. Performing Comparative Experiments
@@
 Experiment parameters are set in the Set Exp Parameter panel (Step 1).
-This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of convergence points to save per run).
+This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of checkpoints to save per run).
 Toggles for Parallel execution and SaveDec (to save decision variables) are also located here.
@@
 This table serves as an excellent starting point for researchers and can also offer valuable insights when selecting algorithms for new real-world problems.
-Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table LABEL:tab:metric, respectively, in the supplementary files.
+Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table 5, respectively.
 Table 6.
@@
   • For standard multi-task problems (e.g., CEC17-MTSO, CEC19-MTMO), the number of tasks is small, making it practical and recommended to analyze per-task metrics directly, such as Obj, IGD+, or HV.
-For multi-objective metrics, IGD+ is generally recommended over the traditional IGD as it provides a more comprehensive and stable measure of both convergence and distribution.
+For multi-objective metrics, IGD+ is generally recommended over the traditional IGD because, unlike IGD, it is weakly Pareto-compliant and therefore provides a more comprehensive and stable measure of both convergence and distribution.
 Furthermore, the HV metric is also provided, as it is a valuable indicator that does not require a known true Pareto front, making it suitable for real-world problems where the optimum is unknown.
@@
   • IOHanalyzer Data (csv): This option exports the complete convergence history (not just the final values) for all selected runs into a csv format.
-This format is specifically structured for use with the IOHanalyzer (Wang2022IOHanalyzer), facilitating advanced external analysis.
+This format is specifically structured for use with the IOHanalyzer (Wang et al., 2022c), facilitating advanced external analysis.
   • Best Decision Variable (mat): This option extracts the decision variables of the best-found solutions of all algorithms and repetitions and saves them to a .mat file for further inspection or use in other applications.
 For single-objective MTO problems, the best solution is determined by the lowest objective value.
-For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu2024Finding; Choong2022Jack).
+For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu et al., 2024; Choong et al., 2023).
 4.2.6. Visualizing Metric Convergence and Pareto Front
@@
 Clicking the Converge button generates a convergence plot for the selected metric results, with a Log toggle available to switch the y-axis to a logarithmic scale, and a Range toggle to display the 95% confidence interval around the mean curve.
-For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto fronts).
-Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang2022IOHanalyzer) for further external analysis and visualization.
+For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto front approximations).
+Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang et al., 2022c) for further external analysis and visualization.
 4.3. Processing Experiment Data
@@
 This enables a complete, script-based workflow: setup, run, and analyze.
-A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD, and programmatically plot the final convergence curves and Pareto fronts is provided in the cmd_examples.m script located in MToP’s root directory.
+A comprehensive set of examples demonstrating this full process including how to instantiate and modify algorithm/problem objects, run experiments, calculate metrics like Obj and IGD+, and programmatically plot the final convergence curves and Pareto front approximations is provided in the cmd_examples.m script located in MToP’s root directory.
 5. Validation and Comparison
 This section first performs reproducible validation tests to ensure that the algorithms and problems implemented in MToP are reliable and accurate.
-Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian2017PlatEMO), is conducted to demonstrate MToP’s comparable efficiency.
+Following this, a performance comparison between MToP and a popular EC platform, PlatEMO (Tian et al., 2017), is conducted to demonstrate MToP’s comparable efficiency.
 5.1. Reproducible Validation
@@
 Beyond internal validation, it is crucial to benchmark MToP’s performance and efficiency against established, state-of-the-art platforms.
-We selected PlatEMO (Tian2017PlatEMO), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison.
+We selected PlatEMO (Tian et al., 2017), a widely-used and highly-regarded platform for evolutionary computation, as the basis for our comparison.
 To ensure a fair and direct comparison, we selected four popular and well-understood single-task multi-objective algorithms that are implemented in both platforms: NSGA-II, SPEA2, MOEA-D, and SMS-EMOA.
@@
 First, we compared the computational efficiency.
-Figure 12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite.
+Fig.
+12 displays the average wall-clock run time for each algorithm across the entire DTLZ suite.
 The results show that the execution times for MToP and PlatEMO are highly comparable.
@@
 Second, we validated the correctness of our algorithmic implementations by comparing the final solution quality.
-Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD metric for 30 runs.
+Table LABEL:tab:results-dtlz presents the mean and standard deviation of the IGD+ metric for 30 runs.
 A Wilcoxon rank-sum test (at a 0.05 significance level) was conducted between the MToP and PlatEMO results for each case.
-The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (23 out of 28).
+The + / - / = summary shows that there is no statistically significant difference in the vast majority of cases (24 out of 28).
 The few minor differences are likely due to small implementation-level variations, but the overall performance is statistically indistinguishable.
-This similarity in performance is further confirmed visually in Figure LABEL:fig:pareto-comparison, which plots the final Pareto fronts from a sample run for each algorithm.
+This similarity in performance is further confirmed visually in Fig.
+LABEL:fig:pareto-comparison, which plots the final Pareto front approximations from a sample run for each algorithm.
 The acquired fronts from both platforms are visually coincident.
@@
 It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field.
-Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng2016RVEA; Deb2014NSGA3), multitask combinatorial optimization codes (Feng2021VRPHTO; Feng2021EEMTA) and more real-world applications (Gupta2021Half; Huang2023Evolutionary) in the future.
+Specifically, MToP is considering the incorporation of many-objective optimization codes (Cheng et al., 2016; Deb and Jain, 2014), multitask combinatorial optimization codes (Feng et al., 2021b, a) and more real-world applications (Gupta et al., 2022; Huang et al., 2023) in the future.
 Acknowledgements.
@@
 The authors would like to thank the researchers who assisted in modifying the code into the MToP pattern and the researchers who open-sourced their codes.
-Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian2017PlatEMO), for which the authors are also thankful.
+Some of the function implementations in MToP about multi-objective optimization are referred to the code implementation in PlatEMO (Tian et al., 2017), for which the authors are also thankful.